### v1

# Neural models of multiscale systems: conceptual limitations, stochastic parametrizations, and a climate application

Abstract
This work explores key conceptual limitations in data-driven modeling of multiscale dynamical systems, focusing on neural emulators and stochastic climate modeling. A skillful climate model should capture both stationary statistics and responses to external perturbations. While current autoregressive neural models often reproduce the former, they typically struggle with the latter. We begin by analyzing a low-dimensional dynamical system to expose, by analogy, fundamental limitations that persist in high-dimensional settings. Specifically, we construct neural stochastic models under two scenarios: one where the full state vector is observed, and another with only partial observations (i.e. a subset of variables). In the first case, the models accurately capture both equilibrium statistics and forced responses in ensemble mean and variance. In the more realistic case of partial observations, two key challenges emerge: (i) identifying the proper variables to model, and (ii) parameterizing the influence of unobserved degrees of freedom. These issues are not specific to neural networks but reflect fundamental limitations of data-driven modeling and the need to target the slow dynamics of the system. We argue that physically grounded strategies – such as coarse-graining and stochastic parameterizations – are critical, both conceptually and practically, for the skillful emulation of complex systems like the coupled climate system. Building on these insights, we turn to a more realistic application: a stochastic reduced neural model of the sea surface temperature field and the net radiative flux at the top of the atmosphere, assessing its stationary statistics, response to temperature forcing, and interpretability.
preprint: APS/123-QED
I Introduction
A key challenge in current applied mathematics is to build data-driven models from partial observations of complex physical systems. Among the most challenging examples is the climate system: a forced, dissipative, and chaotic system whose dynamics span at least 15 orders of magnitude across spatial and temporal scales [1]. This problem has been extensively studied over the past three decades, mainly in the realm of reduced-order models [2]. Seminal contributions include the work of Penland and collaborators [3, 4, 5], the multilevel modeling strategies of Kravtsov and colleagues [6, 7, 8, 9], and the contributions of Majda et al. [10, 11]. More recent proposal also include dynamical Markov-chain models to simulate the evolution of probability distributions as in [12, 13, 14] as well as novel ideas from generative models [15].
Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of the weather [16, 17, 18] and climate [19, 20, 21, 22, 23, 24, 25] systems. These tools build on earlier efforts in data-driven reduced-order modeling by considering new powerful architectures (i.e. neural networks) and by attempting to simulate the system in its full complexity and dimensionality. For example, [24] and [25] respectively seek to emulate the full ocean and atmosphere components of two state-of-the-art climate models. These models often excel at reproducing stationary statistics and generating forecasts across a range of time horizons, marking a promising direction for computationally efficient climate modeling. However, neural network emulators also exhibit several limitations. Long trajectory rollouts often suffer from error accumulation, which can drive the system toward instability. While promising approaches have been proposed to address this issue, e.g. [26, 27], these are currently research directions still to be explored across different models and scenarios. Instead current implementations maintain stability by prescribing part of the climate system as boundary conditions (e.g. [24] is an ocean emulator driven by pre-computed atmospheric fields as boundary conditions). We refer the reader to Pedersen et al. [26] for a detailed discussion of the stability challenges in autoregressive modeling. An additional issue, relevant for this study, is that skill in capturing unforced variability may not always guarantee accurate responses in perturbation experiments, limiting their use for causal mechanistic studies of climate dynamics. This leads to two important consequences: neural emulators (i) usually do not generalize in out-of-distribution climate change experiments and (ii) they may often not capture the right causal relations across variables. Note that the two issues (i) and (ii) above are likely related. The work of Van Loon et al. [28] provides us with a clear and insightful example of this limitation in the context of the “pattern effect” problem [29, 30], i.e. the influence of sea surface temperature (SST) warming patterns on the Earth’s energy imbalance. Van Loon et al. [28] considered the ACE2 emulator [21] and built a Green’s function by iteratively applying warming perturbations in different locations in the sea surface temperature (SST) field and computing the responses in the radiative flux at the top of the atmosphere (TOA). In doing so, [28] showed how such Green’s function does not capture in a quantitative way the causal relations between the SST and the radiative fluxes. This is just one example of the limitations of neural emulators in capturing responses to perturbations. Even if such issues are addressed in specific cases, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent. Other examples related to capturing forced responses to changes in the boundary conditions can be found in [24, 25]. Additionally, we note that current emulators usually focus on one climate component at a time, ocean or atmosphere, and the challenges described above, and also later in this paper, are expected to get worse when a fully coupled system will be considered. In general, limitations of neural emulators in capturing responses to external perturbations imply that such tools may not be yet ready to explore causal questions in climate dynamics.
While the field of neural emulation is advancing rapidly, it is useful to step back and examine its fundamental limitations in capturing forced responses, both from a theoretical standpoint and in controlled settings. In the first part of this paper, we build on a series of previous works by A. Vulpiani and collaborators on the limitations of data-driven modeling [31, 32, 33]. We analyze a simplified stochastic dynamical system to illustrate fundamental limitations that are expected to persist in high-dimensional settings. Specifically, we ask the following question: Can neural autoregressive models, trained on (practically) infinitely long time series, accurately capture the response to general external perturbations? In addressing this question, we consider two scenarios:
  • Fully observed state vector. Data-driven modeling when the entire state vector is available. Here, we assume no knowledge of the underlying equations, but we have access to very long time series of the complete system state vector.
  • Partially observed state vector. Data-driven modeling in the presence of partial observations. This reflects common constraints in modeling high-dimensional systems from data, where the full set of relevant variables is typically unknown. In this more realistic setting, we again assume no knowledge of the underlying equations and only partial access to the system’s state vector.
To tackle such questions we proceed pedagocially by focusing on a “triad model” first proposed by Majda and collaborators in [34] and used as a testbed for Fluctuation-Dissipation theorem strategies in [35, 36]. The model mimics the abstract structure of dynamical cores in large-scale climate models with dynamics characterized by a linear operator and an energy conserving nonlinear, quadratic operator. We address the questions above by training neural stochastic emulators on a long integration of the triad model and test the models’ performance on stationary and perturbed cases. In doing so, we face the following issue: quantifying how an emulator respond to general forcing is an ill-posed question, as there are infinite possible forcings. Nonetheless, the linear response to a general forcing can be computed by convolving the forcing with an impulse response operator [37, 38]. Such operator can be retrieved from the model, and is here considered as the building block to get right for an emulator to, at least, capture the linear part of the response to general external perturbations. In the case of a fully observed state vector we show that the inferred neural model is able to reproduce stationary statistics of the system, i.e. the probability distributions and autocorrelation functions, as well as the response to small external perturbations. Therefore, there is no fundamental limitation for a neural model to respond to external perturbation, i.e. shifts in the input distribution, at least for perturbations of small amplitudes. However, in the case of partial observations, as in more realistic settings, we argue how the performance of a data-driven model necessarily depends on (i) the choice of variables to model and (ii) how to parametrize the unobserved degrees of freedom. The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [5, 39, 40, 41]. Therefore, we will argue that a promising perspective is to build on such established approaches, emphasizing coarse-grained dynamics and stochastic parameterizations, while using neural networks as powerful tools within this framework, rather than as wholesale replacements.
We emphasize that the focus on a simplified dynamical system is not a limitation in this context. In fact, the goal is not to demonstrate the success of a method, but to illustrate, by analogy, fundamental issues that are already evident in low dimensions and are likely to become more pronounced in higher-dimensional settings. As Smith (2002) [42] argues, while it is unreasonable to expect solutions to low-dimensional problems to generalize to high dimensions, so too it is unlikely that limitations observed in low-dimensional systems will vanish in high-dimensional, operational models.
Building on these ideas, we conclude this study by considering a real-world example. Throughout this paper, we emphasize, where possible, the distinction between “real-world examples/systems” and merely high-dimensional dynamics. By real-world examples we refer to situations where (i) we observe only the temporal evolution of a subset of variables without access to the full state vector (i.e. a projection of the full dynamics) or (ii) where even in the presence of all variables the system encompasses a huge number of spatial and temporal scales (as for the climate and weather system). Both cases (i) and (ii) lead to the need to formulate effective rather than primitive equations for both practical and conceptual reasons [40]. For an early instance of case (ii), see the work of Charney (1948) [43], Charney et al. (1950) [44] and the discussion in Dalmedico (2001) [41]. The example we consider simulates the joint evolution of the sea surface temperature (SST) field and the net radiative flux at the top of the atmosphere (TOA). Our motivation lies in understanding the spatial structure of climate feedbacks by addressing the so-called pattern effect, i.e. the dependence of Earth’s energy imbalance on the spatial patterns of surface warming [30]. To this end, we develop a simplified stochastic neural model based on 20 SST modes and a global mean, scalar time series for the radiative flux. Unobserved degrees of freedom are parametrized through multiplicative (state-dependent) noise. We demonstrate that the model captures both the stationary statistics and the radiative fluxes response to imposed temperature perturbations in two climate change experiments. Such model can then be used to explore the linkages between surface temperature warming patterns and radiative fluxes, as well as a “reconstruction” tool, by predicting the change in one variable given prescribed changes in the other. Finally, we apply ideas from response theory to infer direct causal relationships between surface temperature and TOA radiative flux. These results show qualitative agreement with perturbation experiments using atmospheric-only climate models and align with current physical understanding of radiative feedbacks in the climate system.
The paper is organized as follows: in Section II we introduce the theoretical ideas underpinning this work. In Section III we discuss key conceptual difficulties in data-driven modeling of real-world dynamical systems. In Section IV we present a real-world application by focusing on radiative feedbacks to temperature perturbations. Conclusions then follow in Section V.
II Theoretical Framework: Data-driven stochastic climate modeling and response theory
We begin by outlining the main theoretical concepts that underpin this work. First, we briefly introduce general ideas for constructing Langevin-like equations from partial observations of multiscale systems. We then review key aspects of response theory, which offers a general framework for understanding and assessing the ability of models to respond to perturbations.
II.1 Reduced stochastic models: theoretical and practical considerations
Why reduced order modeling and effective equations?
Consider a nonlinear dynamical system $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$. The main interest in data-driven modeling, is to estimate the operator $\mathbf{f}$ directly from data. In doing so we face a serious limitation: in many high-dimensional complex systems we have access only to partial observations of the state vector $\mathbf{x}$. A temptation could be to use Taken’s embedding theorem [45]. While theoretically appealing, this strategy has significant practical limitations: it is not valid for stochastic systems and it becomes infeasible as the dimensionality of the system increases [46, 32, 40, 33, 47, 48]. A more promising direction leverages the multiscale nature of real-world dynamical systems by focusing on coarse-grained, effective dynamics. In some cases, it is possible to identify slow variables exhibiting low-dimensional effective dynamics and treat the cumulative effect of unresolved fast variables on the slow ones as stochastic forcing terms [49, 50, 51, 3, 4, 37, 36, 2, 52]. In these cases, the system state vector $\mathbf{x}$ can be decomposed into slow and fast components, $\mathbf{x}=(\mathbf{x}_{s},\mathbf{x}_{f})$, and motivates the derivation of an effective Langevin equation for the slow variables of the form:
$d\mathbf{x}_{s}=\mathbf{f}_{eff}(\mathbf{x}_{s})dt+\mathbf{\Sigma}(\mathbf{x}_{s})d\mathbf{W},$ (1)
where $\mathbf{f}_{eff}$ represents the deterministic “drift” dynamics, $\mathbf{x}_{s}$ are the slow variables and $\mathbf{W}$ is the standard Wiener process. The noise term parametrize the cumulative effect of unresolved fast processes $\mathbf{x}_{f}$ on the slow ones. In the general case the noise is “multiplicative”, i.e. state dependent: $\mathbf{\Sigma}(\mathbf{x}_{s})$. However, in many practical applications, additive noise can be preferred, with $\mathbf{\Sigma}$ independent of $\mathbf{x}_{s}$. In what follows, we will simplify notation by referring to $\mathbf{x}$ instead of $\mathbf{x}_{s}$, and we will clarify the precise meaning as needed in the text.
The reduced-order modeling approach emerges naturally in the study of large-scale climate dynamics, where recurrent spatiotemporal patterns, i.e. modes of variability, dominate the system’s behavior [53, 54]. Unlike traditional general circulation models (GCMs) that aim to simulate as many spatiotemporal scales as possible, reduced models target specific processes/problems, capturing a few key degrees of freedom while treating the rest as stochastic forcing. This not only reduces computational cost but also enhances mechanistic understanding, by isolating core processes and studying their qualitative behavior (see e.g. [55], as well as the introduction of [44]). Consequently, there is a growing need for systematic methods to reduce the dimensionality of high-dimensional systems and to quantify causal interactions across their degrees of freedom. This has fueled the development of low-dimensional stochastic models learned either from data [51, 3, 4, 56, 10, 7, 2] or from known equations [49, 50, 57, 58], as well as inference-based techniques to infer interactions across modes of variability [59, 60, 61, 62]. In this paper, we will primarily focus on data-driven stochastic modeling from partial observations. Nevertheless, we emphasize that even when all the relevant variables of the system are known, in multiscale physical systems there is often a need to formulate new effective (even deterministic) equations of the slow variables rather than focusing on the primitive equations. A classic example in the weather/climate literature is given by Charney (1948,1949) [63, 43] as well as Charney et al. (1950) [44]. Charney was among the first to recognize the importance of filtering out what he termed “meteorologically insignificant noise” from the primitive equations, leading to more interpretable and computationally efficient effective equations.
Data-driven reduced-order modeling: practical considerations and prior work.
In practice, given a phenomenon of interest, we typically begin with spatiotemporal data consisting of $N$ time series of length $T$. We then often proceed as follows:
  • Coarse-grained dynamics. In spatiotemporal systems like climate, coarse-graining involves: (i) selecting a few relevant modes, (ii) isolating specific temporal scales, and (iii) choosing a limited set of variables. Regarding point (i), we reduce the dimensionality of the system in terms of some bases, a common choice is the one of Empirical Orthogonal Functions (EOFs) [39, 8]. We retain only the first $n<N$ of such bases, and project the data onto them to obtain the principal components (PCs), which serve as the reduced time series for model training. Although EOFs are convenient, they may not be always optimal, and alternatives such as those in [64, 62, 65] (among many others) may be more suitable depending on the application. The performance of the final model is sensitive to all coarse-graining choices (i–iii); see, e.g., [66, 67]. Considering the “right” coarse-grained procedures is perhaps the most difficult step in modeling from data. There is in fact no general mathematical guideline in high-dimensional dynamical systems for choosing the proper variables to model, an old problem [68] emphasized in many new contributions, such as [31, 69, 47, 48].
  • Inferring a stochastic model. Given a reduced order representation of the system, we are now left with the task of inferring a stochastic model of the form of Eq. (1). Many studies have shown that, for specific observables (e.g., sea surface temperature anomalies), a linear drift with additive noise provides forecasting skill comparable to full complexity models [4, 5, 70, 71]. This leads to a Linear Inverse Model (LIM) [3] of the form
    $d\mathbf{x}=\mathbf{L}\mathbf{x}dt+\mathbf{\Sigma}d\mathbf{W},$
    where $\mathbf{L}$ is a linear operator, $\mathbf{\Sigma}$ is a noise matrix and $\mathbf{W}$ is the standard Wiener process. An interesting recent extension to tackle colored noise can be found in [72]. The multilevel regression proposed by Kratsov et al. (2005) [6, 7] generalizes these ideas by incorporating quadratic nonlinearities and non-Markovian effects with a series of “matryoshka”-like levels for residuals. In many practical cases, however, temporal coarse-graining (e.g., averaging over a few time steps) can effectively remove memory effects, allowing a Markovian closure [73]. In the framework proposed in [6, 7], the drift consists of both linear and quadratic nonlinear components. The inclusion of quadratic nonlinearities is physically motivated rather than ad-hoc. As shown in Section 2 of Lorenz (1963) [74], projecting forced-dissipative flows (in periodic domains) onto Fourier modes, yield a set of infinite coupled ODEs with linear and quadratic interactions for the modes’ amplitudes. In particular, the quadratic nonlinearity $\mathbf{B}(\mathbf{x},\mathbf{x})$ should be energy-conserving, satisfying $\mathbf{x}\cdot\mathbf{B}(\mathbf{x},\mathbf{x})=0$. This abstract representation underlies the starting point of the Majda-Timofeyev-Vanden Eijnden (MTV) mode reduction framework [49, 50, 57, 58]. Additionally, Majda and Harlim focused on the energy conservation constraint above to propose physically constrained regression models [11]. Guided by these insights, the neural models proposed in this work will have a deterministic dynamics specified by a linear term $\mathbf{L}$ and general nonlinear terms $\mathbf{n}(\mathbf{x})$; in formulae:
    $\frac{d\mathbf{x}}{dt}=\mathbf{L}\mathbf{x}+\mathbf{n}(\mathbf{x}).$
    It is in fact a sensible choice to guide the learning by specifying a decomposition into linear and nonlinear terms, instead of letting the neural network “figure out” the functional form by itself (see also [75]). Furthermore, we will move beyond additive noise and incorporate multiplicative (state-dependent) noise, both in one and higher dimensions. As we will show, incorporating state-dependent noise can be at times critical for accurately capturing the system’s response to external perturbations in the presence of partial observations.
II.2 Response theory: a (very) brief overview
In this paper, we argue that a reduced-order model should capture not only the system’s stationary statistics but also its response to external perturbations. This naturally raises the question: How to assess skillful responses to general perturbations? Since it is impossible to test responses to the full space of possible forcings, linear response theory provides a principled framework for analyzing the system’s linear response to perturbations.
Impulse response operator.
Given a $n$-dimensional coarse-grained state vector $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$, we perturb a degree of freedom $x_{j}$ with a small, step function perturbation $x_{j}(0)\rightarrow x_{j}(0)+\delta x_{j}(t)$, with $\delta x_{j}(t)=\Delta_{j}\delta(t)$, where $\delta(t)$ is a delta function and $\Delta_{j}$ is a small perturbation amplitude [76, 47]. We then consider a general observable $A(x_{k}(t))$ (i.e. a general functional of the system) and quantify its response, in terms of ensemble average, as:
$\delta\langle A(x_{k}(t))\rangle=\langle A(x_{k}(t))\rangle_{\text{p}}-\langle A(x_{k}(t))\rangle~{},$ (2)
[…]
$R_{k,j}(t)=\lim_{\delta x_{j}(0)\to 0}\frac{\delta\langle A(x_{k}(t))\rangle}{\delta x_{j}(0)}.$ (3)
The utility of the inferred impulse response operator is two-fold:
  • Causal inference. Understanding causal relations across degrees of freedoms in physical systems require us to study the effect of interventions on the system [77]. The inferred response operator quantifies time-dependent cause-effect dependencies across degrees of freedoms $x_{j}$ and $x_{k}$; under the assumption that the reduced order models correctly capture the physics of the system. For more details to this physics-based approach to causal inference see [78, 79, 47, 80, 62].
  • Response to generic time-dependent forcing. The response operator $\mathbf{R}(t)$ can be used to predict the response of the system to small, but general, time-dependent external forcing $\mathbf{f}(t)\in\mathbb{R}^{N}$ applied directly to the system’s degrees of freedoms as:
[…]
    Importantly, while we assume linearity in the system’s response, the formulas above are valid for both linear and nonlinear systems.
A data-driven model capable of accurately capturing the impulse response operator will also capture the time-dependent causal interactions across degrees of freedom. Response theory can then be used to assess the skills of emulators by going beyond purely statistical metrics such as r-squared, by for example assessing the causal sensitivity of the system to external perturbation, see e.g. discussion in Falasca et al. (2024) [62]. Since the linear response to small perturbations is given by the convolution formula in Eq. (4), recovering the impulse response operator ensures a faithful approximation of the system’s linear response to general forcings. For this reason, we view the impulse response operator as a fundamental building block that a data-driven model must reproduce to reliably capture at least the linear component of responses to general perturbations, and is necessary (but not sufficient) to capture nonlinear responses as well.
Inferring the response operator in practice.
Given a model $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W}$ we build the impulse response operator as follows:
  1. 1.
    We simulate a very long trajectory of the model, starting from a random initial condition and remove an initial transient. We then sample $N_{e}$ random points from the simulation to define an ensemble of $N_{e}$ initial conditions on the model’s attractor. $N_{e}$ should be very large (i.e. $N_{e}\gg 1$) in order to sample the whole attractor.
  2. 2.
    For each one of the $N_{e}$ initial conditions, we impose an impulse perturbation $x_{j}(0)\rightarrow x_{j}(0)+\Delta_{j}\delta(t)$ of size $\Delta_{j}$ on the $j$-th degree of freedom at time $t=0$. $\delta(t)$ is the Dirac delta function. The amplitude $\Delta_{j}$ should be theoretically infinitesimally small. In practice we do as follows: we consider the long time series of $x_{j}(t)$ from the long control integration above and define $\Delta_{j}=10^{-2}\sigma_{j}$, where $\sigma_{j}$ is the standard deviation of time series $x_{j}(t)$.
  3. 3.
    Therefore, for a given initial condition $\mathbf{x}(0)$, we simulate two trajectories: a control trajectory without perturbation and a perturbed one, where an impulse perturbation $\Delta_{j}$ has been imposed on the $j$-th degree of freedom at time $t=0$. This procedure is repeated in parallel for all initial conditions, resulting in an ensemble of $N_{e}$ paired of control and perturbed trajectories.
  4. 4.
    At each time $t$ we then estimate the time dependent ensemble average $\langle A(x_{k}(t))\rangle$ of an observable $A(x_{k}(t))$ for both the perturbed and unperturbed run. We refer to the perturbed and control ensemble average respectively as $\langle A(x_{k}(t))\rangle_{\text{p}}$ and $\langle A(x_{k}(t))\rangle$. The observables considered in this study are going to be: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the perturbed distribution, therefore quantifying the response in ensemble mean and variance.
  5. 5.
    We define the impulse response operator for observable $A(x_{k}(t))$ as
    $R_{k,j}(t)=\frac{\delta A(x_{k}(t))}{\Delta_{j}}=\frac{\langle A(x_{k}(t))\rangle_{\text{p}}-\langle A(x_{k}(t))\rangle}{\Delta_{j}}.$
III Limitations of data-driven modeling by analogy: the case of partial observations
As argued by Cecconi et al. (2012) [31], the main challenge in modeling real-world dynamical systems from data stems from their high dimensionality. Learning a dynamical system from data requires sufficiently long time series to sample many recurrences. However, the recurrence time, i.e. the time needed for a trajectory to return to a previous region in state space, scales exponentially with the system’s dimensionality [31, 81]. The success of data-driven methods in real-world systems is therefore often attributed to the presence of an effective low-dimensional dynamics. In other words, the dynamics of the system usually lives on a low-dimensional manifold embedded in a higher-dimensional space. This so-called “manifold hypothesis” underpins research in fields ranging from spatiotemporal chaos [82, 83, 84, 85] to neuroscience [86, 87, 88, 89, 90], biophysics [91, 92], artificial neural networks [93, 94] and climate dynamics [74, 1, 95, 96, 54]. In practice, however, even when a system exhibits low effective dimensionality, we still face several difficult challenges. First, it is often unclear how to construct an optimal basis for efficient modeling in the system’s intrinsic coordinates [64]. Second, and more fundamentally, we typically observe only a subset of relevant variables of a complex, multiscale physical system. Here we illustrate how learning on only partial observations, as is common in real-world applications [10, 33], can hinder data-driven modeling, even when the underlying dynamics is low-dimensional and seemingly tractable. Importantly, while the discussion is framed in terms of neural emulators, motivated by current challenges in climate science, the issues we highlight are general and inherent to data-driven approaches for real-world dynamical systems.
III.1 A triad model as a prototype stochastic climate model
We consider a triad model originally introduced by Majda and collaborators in the context of low-frequency climate dynamics [34, 36]. The model is:
$\begin{split}dx_{1}&=(L_{11}x_{1}+L_{12}x_{2}+L_{13}x_{3}+Ix_{1}x_{2})dt\\ dx_{2}&=(-L_{12}x_{1}-Ix_{1}^{2}-\gamma_{2}\epsilon^{-1}x_{2})dt+\sigma_{2}\epsilon^{-1/2}dW_{2}\\ dx_{3}&=(-L_{13}x_{1}-\gamma_{3}\epsilon^{-1}x_{3})dt+\sigma_{3}\epsilon^{-1/2}dW_{3}\\ \end{split}$ (5)
This idealized system mimics the functional structure of a reduced stochastic climate model, featuring linear couplings $L_{11},L_{12},L_{13}$ and energy-conserving quadratic nonlinearities. The parameter $I$ controls the strength of the nonlinearity: when $I=0$, the model reduces to purely linear dynamics. The parameter $\epsilon$ sets the time-scale separation between the slow mode $x_{1}$ and the fast modes $(x_{2},x_{3})$, allowing exploration of multiscale dynamics in a controlled setting. The terms $W_{2}$ and $W_{3}$ represent independent Wiener processes, introducing stochastic forcing into the second and third equations. This stochastic forcing, together with the linear damping terms $\gamma_{2}$ and $\gamma_{3}$, serves as a linear proxy for the nonlinear self-interactions of the fast variables (see Eq. (4) in [49] or Section 3.3 in [37] for further details). Importantly, the model generates relatively small, but significant, deviations from Gaussianity which is a relevant feature in the study of low-frequency climate dynamics [34] and found in both general circulation models and observational data, see e.g. [97, 98, 99].
III.2 Data-driven modeling when we observe the full state vector
We consider the relatively simple case of a low-dimensional dynamical system with only one characteristic time scale (i.e. $\epsilon=1$). The parameters chosen for this first test are: $L_{11}=-2$, $L_{12}=1$, $L_{13}=1$, $\gamma_{2}=0.6$, $\gamma_{3}=0.4$, $\sigma_{2}=1.2$, $\sigma_{3}=0.8$, $I=1$, $\epsilon=1$. We now integrate the system in (5) with an Euler–Maruyama scheme for a period of length $T=10^{7}$ with time step $dt=0.01$ (i.e. $10^{5}$ Model Time Units). As we will show later, given the system’s decorrelation time scale, the resulting time series can be effectively treated as infinitely long. With this long trajectory of the full state vector $\mathbf{x}(t)=[x_{1},x_{2},x_{3}](t)$ in hand, we now attempt to learn data-driven models directly from the time series. We construct both linear and nonlinear stochastic models. We specify the functional forms of these models and assess their ability to reproduce both the stationary and perturbed statistics of the original system defined in Equation (5).
III.2.1 Functional form of the data-driven models considered
We consider two data-driven models and report their functional forms below. To keep the main text focused, we leave the technical details of the fitting procedures, along with details of the neural networks, in Appendix A.
  • Linear model. We consider the following linear model:
    $d\mathbf{x}=\mathbf{L}\mathbf{x}dt+\mathbf{\Sigma}d\mathbf{W},$ (6)
    $\mathbf{L}$ is a linear operator, $\mathbf{\Sigma}\mathbf{\Sigma}^{\text{T}}$ is the noise covariance matrix and $\mathbf{W}$ is the standard Wiener process.
  • Nonlinear model. The nonlinear model takes the following form:
    $d\mathbf{x}=(\mathbf{L}\mathbf{x}+\mathbf{n}(\mathbf{x}))dt+\mathbf{\Sigma}d\mathbf{W}.$ (7)
    $\mathbf{L}$ is a linear operator and $\mathbf{n}(\mathbf{x})$ is parametrized by a Multilayer Perceptron (MLP). Crucially, the term $(\mathbf{L}+\mathbf{n}(\mathbf{x}))$ is fitted jointly. As for the linear model, the $\mathbf{\Sigma}\mathbf{\Sigma}^{\text{T}}$ is the noise covariance matrix and $\mathbf{W}$ is the standard Wiener process. In what follows we are going to often refer to this model as “neural model” or “neural emulator”.
The two models above are trained on a long simulation of the triad system in Eq. (5).
Note that this is the only Section in this work where we compare linear and nonlinear autoregressive models. The linear method is considered for two reasons. First, we show that linear inverse modeling (LIM), as originally proposed in [3], serves as a strong baseline for neural autoregressive models, successfully capturing both the mean and variance of the system’s stationary distribution, the autocorrelation functions, and the ensemble-mean response to external perturbations. Second, and consistent with theoretical expectations [36], we demonstrate that linear models are fundamentally limited in capturing changes in the full probability distribution: for instance, they cannot represent any response in the variance to external forcings. This limitation restricts their usefulness for studying forced changes in distributions beyond the mean, which, dependent on the research question, can be a central objective in climate science.
III.2.2 Stationary statistics
We integrate the two data-driven models inferred from data and specified in Eq. (6) and (7) for $T=10^{7}$, yielding time series of the same length as those obtained from the original triad system. We then compare the modeled and original probability density functions (PDFs) and autocorrelation functions (ACFs), as shown in Figure 1. The neural model described in Equation (7) successfully captures the stationary distributions of all three modes $x_{1}(t),x_{2}(t)$ and $x_{3}(t)$. In particular, it outperforms the linear model by accurately reproducing the non-Gaussian features of the first mode $x_{1}$. Both models also reproduce the autocorrelation functions well, as seen in the second row of Figure 1. In summary, the neural emulator defined in Equation (7) is capable of capturing both the stationary statistics and temporal correlations of the full state vector $\mathbf{x}(t)$. At the same time, the linear model serves as a strong baseline for evaluating more sophisticated data-driven approaches.
Figure 1: First row: stationary distributions of the variables $x_{1}(t),x_{2}(t)$ and $x_{3}(t)$ as modeled by the original triad system in Eq. (5), the linear emulator in Eq. (6) and the neural emulator in Eq. (7). Stationary distributions are obtained after running the three systems for $10^{5}$ model time units. Second row: same as the first row but for the autocorrelation functions.
III.2.3 Response of probability distributions to external perturbations
Impulse response functions
We now examine the impulse response of the original triad model (Equation (5)), the linear model (Equation (6)), and the nonlinear neural model (Equation (7)). Specifically, we analyze the time-dependent response of observables $A(x_{k}(t))$, with $k=1,2,3$, to an impulse perturbation imposed on $x_{1}(0)$ at time $t=0$. The observables considered here are the mean and variance of the time-dependent perturbed distribution. The ensemble size considered here is $N_{e}=10^{5}$ trajectories. The results, shown in Figure 2, reveal that the nonlinear neural model captures the mean response of the distribution quantitatively, while the linear model captures it only qualitatively. The most significant differences arise in the response of the variance: the linear model yields exactly zero variance response, clearly highlighting the limitations of linear dynamics. This is expected, as a linear model with a Gaussian invariant measure produces zero response to any quadratic functionals, such as the variance, consistent with the theoretical results in [36]. In contrast, the neural model captures the variance response with relatively high fidelity. As argued in Section II.2, the impulse response operator plays a foundational role in determining a model’s ability to capture the linear part of the response to general perturbations. We therefore consider the neural model presented here as a skillful model both in terms of reproducing the stationary distribution as well as the response to forcing.
Figure 2: Impulse response functions as modeled by the original triad system in Eq. (5), the linear emulator in Eq. (6) and the neural emulator in Eq. (7). First row: response of the ensemble mean to impulse perturbations imposed on $x_{1}(0)$ at time $t=0$. Second row: response of the ensemble variance to impulse perturbations imposed on $x_{1}(0)$ at time $t=0$. Impulse response functions are computed using an ensemble size of $N_{e}=10^{5}$ members.
Responses to a specific logarithmic forcing.
Having evaluated the ability of the data-driven neural model to capture the impulse response operator, we now consider a more targeted test involving a specific, time-dependent forcing. We introduce a logarithmic forcing term to the right-hand side of the original triad model (Equation (5)), the linear model (Equation (6)), and the neural model (Equation (7)). The forcing is applied only to the first mode $x_{1}$ and takes the form $(0.01log(1+t),0,0)$ per time step. Each model is integrated for 2500 time steps (25 Model Time Units), and we compute the difference between the forced and unforced simulations in terms of ensemble-mean and ensemble-variance responses. Results are presented in Figure 3. The neural model accurately reproduces the forced responses across all three components. The linear model performs qualitatively well but exhibits significant deviations in the ensemble-mean response of the second mode $x_{2}$ as shown in Figure 3(b). This is consistent with the earlier Section: the largest error in the linear model’s impulse response also occurred in $x_{2}$ (see Figure 2(b)). This supports the interpretation that accurately capturing the impulse response operator is critical for a model’s reliability in predicting responses to general, time-dependent forcings. As expected from the analysis of the response operator, the linear model is not able to capture the response in ensemble variance. The neural model shows skillful response, capturing the time-dependent changes in ensemble variance relatively well, with the largest error in the second mode $x_{2}$. In general, considering that the changes in variance are very small, we conclude that the neural emulator is skillful in responding to the imposed logarithmic forcing.
Figure 3: Response to a small logarithmic forcing $F=(0.01log(1+t),0,0)$ imposed on the right-hand-side in Eq. (5), the linear emulator in Eq. (6) and the neural emulator in Eq. (7). First row: response of the ensemble mean to external forcing. Second row: response of the ensemble variance to external forcing. Responses are computed using an ensemble of $N_{e}=10^{5}$ members.
III.2.4 Main takeaways
Given a (practically) infinitely long trajectory of a stochastic dynamical system, and assuming the ideal condition where the full state vector is observed, it is possible to construct a neural stochastic model that is stable, reproduces the stationary statistics, and responds accurately to small but general external perturbations and forcings. While these results are encouraging, they remain far from any real-world scenario. In practice, even when a low-dimensional effective dynamics exists, the main challenge is that we typically have access only to partial observations of a complex, interconnected system. Such challenges lead to the important problem of having to choose the proper variables to model as well as the proposal of stochastic parametrizations to account for the unobserved degrees of freedom.
III.3 Data-driven modeling when we do not observe the full state vector
We now return to the triad model defined in Eq. (5) and consider the more realistic scenario of model inference under partial observations and multiple timescales. The parameters chosen for this test are: $L_{11}=-2$, $L_{12}=0.2$, $L_{13}=0.1$, $\gamma_{2}=0.6$, $\gamma_{3}=0.4$, $\sigma_{2}=1.2$, $\sigma_{3}=0.8$, $I=1$, $\epsilon=0.01$. In this setting the variable $x_{1}$ clearly is the slow mode, forcing the much faster modes $x_{2}$ and $x_{3}$. We again integrate the system in (5) with an Euler–Maruyama scheme for a period of length $T=10^{7}$ with time step $dt=0.01$ (i.e. $10^{5}$ Model Time Units).
Unlike the case considered in the previous section, where the entire state vector $\mathbf{x}(t)=[x_{1},x_{2},x_{3}](t)$ was available, we now consider a more challenging scenario: only one component of the system is observed (or available or chosen), either $x_{1}$, $x_{2}$, or $x_{3}$. The modeling task is thus reduced to constructing a scalar stochastic model for the observed component alone. We require the scalar stochastic model to satisfy two criteria:
  1. i.
    it must reproduce the statistics of the observed component;
  2. ii.
    it must correctly captures its response to perturbations as in the original triad system.
Criterion (ii) is particularly demanding because, in the triad system defined by Eq. (5), perturbations can propagate through the network of interacting variables, inducing cascades of feedbacks. These time-dependent feedback loops must be effectively parametrized within the scalar model, despite the absence of direct observations of some components. While clearly nontrivial, this problem is representative of real-world data-driven modeling scenarios, where models are often built from partial observations of high-dimensional systems.
Proposing a scalar data-driven model from partial observations that satisfies the two criteria above raises two central challenges:
  • Choice of variable to model. The derived reduced order model is now dependent on the variable we choose to model. Not all components can be modeled to meet the above criteria. In the simple case considered here the choice of $x_{2}$ or $x_{3}$ will be wrong as these two modes represent the fast dynamics forced by the much slower mode $x_{1}$. Therefore, it would be impossible to parametrize the effect of $x_{1}$ from information of $x_{2}$ or $x_{3}$ alone. The only possible way forward is to choose the slow mode $x_{1}$. Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model in the presence of only partial observations [37, 31, 52, 69].
  • The need for suitable parametrizations. Having identified the proper variable $x_{1}$ to model, the next challenge is to appropriately parametrize the cumulative effects of the unresolved components $x_{2}$ or $x_{3}$ onto the slow mode $x_{1}$. This leads to the classic problem of stochastic parametrization [100, 101, 2, 1, 39, 102]. As we will show, the success of the resulting model, particularly in capturing forced responses, hinges critically on the quality of this parametrization.
III.3.1 Functional form of the reduced order models considered
Given the correct choice of variable $x_{1}$ to model we can now proceed in proposing the two data-driven scalar models considered. As in the previous Section, to keep the main text focused and readable, we leave the technical details of the fitting procedures for Appendix A. We consider two nonlinear, neural models differing with the stochastic parametrization considered:
  • Neural model with additive noise. We consider the following neural model fitted using an artificial neural network, i.e. a Multilayer Perceptron (MLP) with one hidden layer. The model takes the following form:
    $dx=(Lx+n(x))dt+\sigma dW,$ (8)
    $L$ is a linear operator (here just a scalar), $n(x)$ is a nonlinear operator and $\sigma^{2}$ is the noise variance. $W$ is the standard Wiener process. Note: the whole drift $(Lx+n(x))$ is fitted by the neural network.
  • Neural model with multiplicative noise. We consider a neural model as before, but now defined with multiplicative rather than additive noise:
    $dx=(Lx+n(x))dt+\sigma(x)dW.$ (9)
    Differently from the previous model in Eq. (8), now the noise is also state dependent. The multiplicative noise $\sigma(x)$ is fitted with an additional neural network, see Appendix A. As before, we are going to often refer to the two models above as “neural model” or “neural emulator”.
To train the two models above we consider a long simulation of the triad system in Eq. (5) using the set of parameters reported in Section III.3. We then extract only the time series of $x_{1}$. The two models in Eq. (8) and (9) are then trained on this long, scalar time series of $x_{1}(t)$ alone.
III.3.2 Stationary and perturbed statistics of the scalar model
We integrate the two data-driven scalar models and specified in Eq. (8) and (9) for $T=10^{7}$ time steps, yielding time series of the same length as the time series of $x_{1}(t)$ obtained from the long integration of the triad system used for training. To evaluate the models’ ability to capture both stationary and perturbed statistics, we examine the stationary probability distribution and the impulse response functions. Specifically, we focus on responses in the ensemble mean and variance by studying the time dependent change of the perturbed probability distribution $P_{p}(x,t)$ (where the subscript “$p$” stands for “perturbed”) to impulse perturbations in the variable $x$ itself, see Section II.2 for more details.
The neural model with additive noise (Figure 4(a)) fails to accurately capture the stationary statistics in a quantitative way. Moreover, the ensemble mean response to an impulse perturbation decays too slowly, and notably, the model fails to reproduce any significant response in the ensemble variance. In contrast, incorporating multiplicative noise (Eq. (9)) leads to a substantially improved representation of the stationary distribution, as seen in Figure 4(e), which closely matches the true distribution. The mean response remains largely unchanged, decaying too slowly, but the model now captures a qualitatively correct signal in the ensemble variance, albeit with a similarly slow decay.
These results indicate that the inclusion of multiplicative noise improves the reduced-order model’s ability to represent both stationary statistics and variance responses to perturbations. In the full triad system, a perturbation to $x_{1}$ at time $t=0$ leads to changes in $x_{1}$, $x_{2}$ and $x_{3}$ at the next time step $dt$. Consequently, $x_{2}$ and $x_{3}$ feedback on $x_{1}$ at the consequent time step. This leads to a cascade of state-dependent interactions until the system relaxes back to stationarity. Such complex, state-dependent feedbacks cannot be represented by additive noise alone and require the inclusion of state dependent noise $\sigma(x)$, effectively parametrizing such interactions. However, even when multiplicative noise is considered, the responses to impulse perturbations are still only qualitatively right. The example illustrates that accurately capturing stationary statistics does not guarantee faithful reproduction of perturbed statistics.
Figure 4: First row: stationary probability distribution (panel (a)); ensemble mean response to an impulse perturbation (panel (b)); ensemble variance response to an impulse perturbation (panel (c)) for the variable $x_{1}$ in the triad system in Eq. (5) and the variable $x$ in the neural scalar emulator with additive noise in Eq. (8). Second row: same as the first row but we are now considering the neural scalar emulator with multiplicative noise in Eq. (9). Impulse response functions are computed using an ensemble of $N_{e}=10^{5}$ members.
The presence of slowly decaying responses in the ensemble mean and variance to perturbations is indicative of a system that is no longer Markovian. The reduced-order stochastic scalar model is trained solely on the time series of $x_{1}(t)$. However, the time series $x_{1}(t)$ is generated by the full triad system in Eq. (5) and implicitly reflects the influence of the unobserved variables $x_{2}$ and $x_{3}$ on it. The variability of the unobserved variables $x_{2}$ and $x_{3}$ manifest as memory effects in the reduced order scalar model for $x_{1}$. One strategy for addressing such memory effects is the multilevel regression approach first introduced by Kratsov, Kondrashov and Ghil (2005) [6] and later generalized by Kondrashov, Checkroun and Ghil (2014) [7]. The multilevel regression strategy method augments the reduced-order model with a hierarchy of linear equations solving for the residuals from the deterministic dynamics, and can be thought as a practical implementation of the Mori–Zwanzig formalism in statistical mechanics [7, 2]. Ultimately, the remaining residuals become white in time, allowing the system to be closed using a Wiener process. Similarly, it could be possible to address the issue through recurrent neural networks [103]. An alternative, practical approach for handling non-Markovian effects is to coarse-grain the data in the temporal dimension, for instance, by averaging over a time window. This suppresses fast fluctuations within that window leading to an effective “Markovianization” of the the dynamics. Due to its simplicity and ease of implementation, this method is commonly used in the literature (see for example [4]) and is also adopted in the present study.
We therefore consider again the time series $x_{1}(t)$ from the long simulation of the triad system in Eq. (5). As an additional step, we average the time series over a short temporal window of 10 time steps (recall that $dt=0.01$, so this is equivalent to $0.1$ Model Time Units). Importantly, this is not equivalent to sub-sampling the original dynamics by taking every 10th data point. Instead, averaging over a 10-step window smooths out fast variability and effectively defines a new, coarse-grained “more Markovian” system. We then re-train the neural models with additive and multiplicative noise on the new coarse-grained data. Next, we simulate the neural models over long timescales to analyze their stationary distributions, as well as their ensemble mean and variance responses to impulse perturbations. Results are shown in Figure 5. As before, the stochastic model with additive noise fails to capture the correct stationary statistics (Figure 5(a)). While the response in the ensemble mean improves relative to the non-coarse-grained case, the response in ensemble variance remains suppressed. In contrast, the model with multiplicative noise not only captures the stationary distribution and further improves the mean response, but now also closely reproduces the response in variance, capturing the peak magnitude and deviating only slightly in its temporal decay. Note that the stationary response in ensemble mean and variance to impulse perturbations of the original degree of freedom $x_{1}$ are the same in both Figures 4 and 5, the only thing changing in the latter is that the reduced-order model is defined in a coarse-grained dynamics.
Figure 5: Same as in Figure 4 however here the scalar emulators in Eq. (8) and Eq. (9) have been trained on a long trajectory of $x_{1}(t)$ after coarse-graining by averaging every 10 time-steps. No additional preprocessing was performed for the original triad system in Eq. (5), i.e. the blue curves in this Figure and in Figure (4) are the same.
III.3.3 Main takeaways
Even with access to very long time series, modeling real-world dynamical systems from data alone remains fundamentally constrained by observing only a subset of the full state vector. This leads to two major challenges: (i) the resulting data-driven model will depend heavily on the specific variable chosen for modeling, and (ii) once a suitable slow variable is identified, the cumulative effect of the unobserved degrees of freedom must be appropriately parametrized. These issues, demonstrated at the “practical level” within the context of a low-dimensional model, become even more severe in realistic high-dimensional systems, which often lack a clear time-scale separation. These challenges are intrinsic to data-driven approaches for dynamical systems, and simply extending time series length or training increasingly large neural networks is unlikely to resolve them.
The core difficulty is conceptual: there is no rigorous mathematical framework for selecting the proper variables to model in real-world, high-dimensional dynamical systems. This poses fundamental limitations for purely data-driven approaches. In practice, such choices rely on physical intuition and prior knowledge. When possible, selecting the right variables amounts to identifying an appropriate coarse-grained description of the dynamics, including relevant fields and suitable spatial and temporal scales. Consequently, effective data-driven modeling of real-world complex systems hinges on deep physical insight and strong domain expertise.
IV A real-world example: radiative responses to temperature perturbations in a climate model
Building on the ideas discussed above, we present a stochastic reduced-order model tailored to a real-world application. The emulator is designed to capture the large-scale dynamics of sea surface temperature (SST) and the global mean net radiative flux at the top of the atmosphere (TOA). To achieve this, we restrict our focus to a limited range of spatiotemporal scales by selecting a few dominant modes, using monthly-averaged data, and high-pass filtering to remove multidecadal variability. The cumulative impact of fast, unresolved processes on the slow dynamics is parametrized by multiplicative, state-dependent noise. We then evaluate the model’s response to temperature changes and assess its utility in improving our understanding of radiative feedbacks.
IV.1 Scientific problem at hand
Radiative forcing (e.g., as resulting from a change in $\text{CO}_{2}$ concentration) drives warming in the Earth system, leading to changes in sea surface temperatures (SSTs) and associated radiative feedbacks. Recent studies have shown that even given the same global mean temperature warming, different spatial patterns of SST warming can produce markedly different radiative feedbacks. For example, warming concentrated in the eastern Pacific tends to produce positive radiative feedbacks, thereby amplifying the initial temperature change. In contrast, greater warming in the western Pacific or warm-pool region tends to have a stabilizing effect [29, 30]. This sensitivity of the Earth’s energy imbalance to the spatial pattern of surface temperature change is commonly referred to as the “pattern effect”. The pattern effect captures the fast response of top-of-atmosphere (TOA) radiative fluxes to spatial variations in temperature. It is theoretically defined in the framework of energy balance models by Taylor expanding the radiative response over the steady state temperature (see for example Section 2 of Meyssignac et al. (2023) [104]). A general data-driven framework to explore radiative responses from data has recently been proposed in [105], with the pattern effect arising as a limiting case in the case of fast responses. The framework in [105] focuses on both direct causal links between SST and radiative fluxes as well as cumulative (in time) links.
A widely used approach to study the relationship between SST patterns and TOA radiative fluxes is through Green’s function experiments [106], in which the radiative response to localized SST perturbations is quantified using atmosphere-only models [107, 108, 109, 30, 110]. These experiments are extremely valuable for rigorously testing model sensitivities. However, they are computationally expensive and typically require large SST perturbations (e.g., 1–4 K) to generate a measurable response within short simulations. These large perturbations can lead to nonlinear responses, invalidating the linearity assumptions adopted in the Green’s function framework [111]. There is therefore a need for fast and interpretable data-driven models capable of performing large-ensemble experiments at low computational cost and with small perturbation amplitudes. Moreover, climate models are known to be biased, while data-driven emulators offer, in principle, a way to learn climate dynamics directly from observations.
There are two broad strategies to build such emulators. One is to construct a general-purpose emulator of the entire climate system and then use it to study specific problems. However, as shown by Van Loon et al. (2025) in the context of the ACE2 emulator [28], this strategy may not always succeed. A second, more traditional and targeted approach is to design an emulator focused on the problem at hand, e.g. [5]. While this approach sacrifices generality, it offers important advantages: the focus on a specific problem often comes together with a greater understanding of what are the proper variables to model as well as the spatial and temporal scales to consider. These choices, typically informed by physical intuition and prior literature, can directly shape the architecture and inputs of the emulator. Moreover, by isolating only core relevant processes for the specific problem considered, such an emulator can remain interpretable, thereby helping to bridge the gap between simulation and understanding [55].
IV.2 A reduced order model of surface temperature and radiative fluxes
Given the general problem outlined above, we aim to build a skillful emulator of sea surface temperature and radiative fluxes that allows us to study stationary and perturbed statistics. Building on the previous Sections we consider a functional form of the neural model defined follows:
$d\mathbf{x}=(\mathbf{L}\mathbf{x}+\mathbf{n}(\mathbf{x}))dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W},$ (10)
Where $\mathbf{W}$ is the standard Wiener process. As in the previous Sections, we use a neural network to fit the drift part $(\mathbf{L}+\mathbf{n}(\mathbf{x}))$ and another network to fit the multiplicative noise $\mathbf{\Sigma}(\mathbf{x})$. As before, to improve readability we refer to Appendix A for technical details.
We now face the challenging task of selecting appropriate preprocessing steps to define the coarse-grained dynamics to be modeled. These coarse-graining choices are crucial, as the real climate system involves not only SST and radiative fluxes, but also many other interacting variables, i.e. this is a practical example of the case of a partially observed state vector. We now describe the dataset used for this experiment, along with the preprocessing steps taken. The preprocessing procedure largely follows that described in [105] and we will refer the reader to that work for additional details and clarifications where needed.
IV.2.1 Data, preprocessing and coarse-graining
Data.
The focus in on a state-of-the-art coupled climate model GFDL-CM4 [112]. The model offers data from a long control run simulating the coupled climate system for 600 years with constant $\text{CO}_{2}$ fixed at pre-industrial level. The ocean component is the MOM6 ocean model [113] with a horizontal grid spacing of 0.25 and 75 vertical layers. The atmospheric component is the AM4 model [114, 115] with a horizontal grid spacing of roughly 100 km and 33 vertical layers. The model considered in Eq. (10) will be trained on the long control run. The variables considered are two: SST and the global mean net radiative flux at the top of the atmosphere (hereafter, often referred to as TOA). The focus on global mean net radiative flux, rather than its spatiotemporal representation is motivated by the fact that such variable does not show large scale patterns of variability. It is in fact a fast variable in respect to SST and the main signal we are interested in is its global average. The SST field is considered only in the latitudinal range $[-45,45]$, a choice informed by earlier studies highlighting the prominence of radiative feedbacks in the tropics and subtropics [30, 108].
Preprocessing and coarse-graining.
Climate dynamics is not only encompassing the SST and radiative fluxes. Therefore, we aim in building a reduced order model focusing on the slow dynamics and parametrizing unobserved degrees of freedom as noise. We consider the control run and proceed in coarse-graining both the spatial and temporal scales. First, starting from daily data, we compute monthly averages. We then remove the seasonal cycle from both the SST and radiative fluxes, allowing us to focus on seasonal anomalies, rather than periodic signals driven by the annual solar forcing. We then decompose the SST spatiotemporal field in a set of $n$ modes, i.e. empirical orthogonal functions (EOFs). Projecting the computed EOFs onto the original data give us $n$ Principal Components (PCs). In formulae, given the SST field $\mathbf{X}(t)\in\mathbb{R}^{N}$, we decompose it as:
$\begin{split}\mathbf{X}(t)=\sum_{i=1}^{n}x_{i}(t)\mathbf{e}_{i}+\sum_{j=n+1}^{N}y_{j}(t)\mathbf{e}_{j},\end{split}$ (11)
and retain only the first $n=20$ modes $x_{i}(t)$, while the remaining components $y_{j}(t)$ represent the fast modes. $\mathbf{e}$ are orthonormal bases (i.e. the EOFs). Therefore, the time series $x_{i}(t)$ will be used to train the emulator in Eq. (10). These two steps, i.e. one month averaging and the projection over EOFs, attempt to define a new coarse-grained system for the slow variability of the SST field. Note that Penland and Sardeshmukh [4] and Penland (2006) [5] already showed how a similar preprocessing would lead to a skillful linear Markov model of the tropical Pacific SST. These SST principal components are then concatenated with the scalar net radiative flux time series: to simplify the notation, we are going to refer to the state vector of the reduced order model as simply as $\mathbf{x}(t)\in\mathbb{R}^{n+1}$.
Is a 600 year-long control run a long-enough simulation to learn a model of SST and radiative fluxes? Whether this is a sufficiently long simulation depend on whether the time series considered sample enough recurrences, see Cecconi et al. (2012) [31]. In a multiscale system like climate, even in the presence of constant external forcing, we observe processes such as multidecadal oscillations which are only sampled a few times in a 600-year long experiment. Therefore, we make the following choice: we high-pass filter the data with a cut-off frequency of $10^{-1}$ years. This removes all low-frequency variability which we would have little chance to learn from data in the first place, and focus on variability inside 10 years only. This choice involves a trade-off: our model will not be able to capture multidecadal variability and we will restrict ourselves to explore subdecadal variability only. At the same time this leads to an effectively larger sample size than we would have had by keeping all time scales. Additionally, this step also allow us to remove drifts present in the control run of the CM4 run, mainly coming from an ocean component that still did not equilibrate.
In this paper we will focus on testing the ability to “generalize” to unseen climate scenarios: we will fix the SST variable with a precomputed trajectory coming from two, 150 years long forced runs, namely a $1$pctCO2 and 4×4$\times$4 ×CO2, and use this information to reconstruct the change in global mean TOA in these simulations. The $1$pctCO2 is an idealized experiment simulating the climate system under a 1% $\text{CO}_{2}$ increase per year starting from pre-industrial conditions. The 4×4$\times$4 ×CO2 is an idealized experiment simulating the climate system under an abrupt quadrupling of $\text{CO}_{2}$ concentration from its pre-industrial level at time $t=0$. Importantly, considering the $1$pctCO2 and 4×4$\times$4 ×CO2 simulations leads to an additional challenge: the radiative fluxes are directly influenced by an increase in atmospheric $\text{CO}_{2}$. However, our reduced order model only has access to SST and radiative fluxes and it cannot explicitly include $\text{CO}_{2}$ concentration. To address this, we remove the portion of the net TOA radiative flux that is attributable solely to rising $\text{CO}_{2}$. This allow us to isolate and study the component of radiative feedback that is driven by changes in SST. This correction relies on the well established fact that the global-mean radiative forcing from $\text{CO}_{2}$ alone increases approximately logarithmically with its concentration [116, 117]. As is standard, for the 4×4$\times$4 ×CO2 experiment, we remove a standard constant of $8Wm^{-2}$ [118]. For the $1$pctCO2 run we instead apply a time-dependent correction of the form $\alpha\log[C_{t}/C_{0}]$; $C_{0}$ and $C_{t}$ respectively representing the concentration of $\text{CO}_{2}$ in the control run and forced experiments.
We note that considering 1 month averages in TOA fluxes is not arbitrary: the immediate response of TOA fluxes to perturbations in SST is fast (e.g., $\sim 20$ days in tropical large-scale convective overturning motion in the GFDL-CM4 model). Therefore, reconstructing the net radiative fluxes given information of the SST in the two forced runs above is a relatively easy task for this system and it converts to a 1 time step prediction. For this point we refer the reader to the conceptual discussion in Section 7 of [105]. In other words, the preprocessing considered here maximizes the chance of having a data-driven model that correctly reconstruct the TOA variable in unseen climates.
IV.2.2 Stationary statistics and forced responses
Stationary statistics.
Given the $n+1$ state vector $\mathbf{x}(t)\in\mathbb{R}^{n+1}$ defined in the preprocessing step above, we can now train the emulator defined in Eq. (10) on the long control run of the GFDL-CM4 model. The trained emulator is then used to simulate a 600 years long climate trajectory (same length as the one used in the original model GFDL-CM4). The stationary probability distributions of the first SST mode and of the net radiative flux at the TOA are found to be qualitatively well captured, see Figure 6(a,b). The largest discrepancies between the emulator and the original model are found for the autocorrelation functions. Autocorrelation functions are shown for the first SST mode (largely corresponding to the El Niño Southern Oscillation (ENSO) [119]) and for the TOA variable.
Forced responses.
The pattern effect problem discussed in IV.1 concerns identifying a direct causal link between patterns of SST warming and changes in radiative fluxes at the TOA. Before attempting any mechanistic interpretation or physical understanding, we first assess whether the neural model we have built possesses predictive skill in capturing forced responses. Specifically, we test if the neural model can be used to reconstruct the global mean radiative flux from a pre-computed, forced sea surface temperature field in forced experiments. In doing so, we also test an important question in neural models: is the neural network model able to generalize to out-of-distribution predictions? To answer these question we consider the two, 150-years long, forced runs of the GFDL-CM4 climate model introduced in Section IV.2.1. We extract the sea surface temperature from these forced runs and project it onto the set of $n=20$ EOF bases found in the control run. We refer to such forced modes as $\mathbf{x}_{\text{SST}}^{\text{f}}$. We then simulate the neural emulator for 150 years by constraining the SST degrees of freedom to be exactly equal to $\mathbf{x}_{\text{SST}}^{\text{f}}$. Given $\mathbf{x}_{\text{SST}}^{\text{f}}(t)$ the model returns the TOA variable, $x_{\text{TOA}}(t+1)$, at the next time step, starting from initial condition $x_{\text{TOA}}(0)=0$. The results is shown in Figure 6(e,f). In both the case of $1$pctCO2 or 4×4$\times$4 ×CO2 runs, the model is able to reconstruct the change in radiative fluxes at the TOA in a quantitative way. We highlight that our preprocessing allowed us to reduce this reconstruction step into a prediction of the $21$-st entry of the state vector at time $t+1$, given information of the first 20 entries at the previous time $t$. This is only possible because of our focus on 1-month averages: the radiative fluxes at the TOA largely respond already at a sub-monthly time scales to temperature perturbations, therefore allowing us to focus on a direct statistical link $\mathbf{x}_{\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$. We also note that it is perhaps surprising that the neural network is able to respond faithfully even when for nonstationary inputs and with temperature values far out of the training distribution (in physical units, the mean temperature anomaly shifts from a $0K$ to $5K$ from the control in both forced simulations).
Figure 6: Panel (a): stationary distributions obtained for the 1st principal component of the sea surface temperature (SST) field as in the original GFDL-CM4 climate model and as emulated by the neural emulator with multiplicative noise in Eq. (10). Units: nondimensional. Panel (b): same as panel (a) but for the global mean radiative flux at the top of the atmosphere (TOA). Units: $W/m^{2}$. The stationary distributions have been obtained after simulating the neural emulator for 600 years. Second row: same as the first row but for the autocorrelation functions. Third row: reconstruction of the net radiative flux at the TOA by the neural emulator in Eq. (10) after constraining the SST degrees of freedom to follow a precomputed trajectory in two forced experiments in the GFDL-CM4 model. The forced experiments considered are the $1$pctCO2 and 4×4$\times$4 ×CO2, see Section IV.2.1 for details. Note: the net flux at the TOA variable is negative upward (i.e., leaving the planet).
IV.2.3 Interpretability and understanding
In the previous Section we showed that the neural emulator proposed in this study can in fact be used to reconstruct the instantaneous change in radiative feedback at the top of the atmosphere given pre-computed changes in the temperature field. As emphasized before, the task of the emulator in the reconstruction/prediction task reduces to the following mapping: $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\in\mathbb{R}^{n}$, $x_{\text{TOA}}(t)\in\mathbb{R}$ and with initial condition $x_{\text{TOA}}(0)=0$. We now ask the following question: What are the SST patterns that contribute the most to this skillfull prediction? To answer this question, we must go beyond the prediction task presented above and identify which components of the sea surface temperature input contributes the most to changes in radiative flux. We explore this question through the framework of linear response theory, by capturing the direct causal link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $x_{j,\text{SST}}(t)$ represents the $j$-th component of the surface temperature field at time $t$.
We briefly describe the strategy to infer this direct causal link with response theory, and refers to Section II.2 for details. We consider the neural emulator in Eq. (10) and iteratively perturb the $j$-th sea surface temperature component of the state vector at time $t=0$ with a small impulse perturbation $x_{j}(0)+\Delta_{j}\delta(t)$. $\delta(t)$ being a Dirac delta and $\Delta_{j}$ being a perturbation of very small amplitude. The amplitude $\Delta_{j}$ depends on the component $x_{j,\text{SST}}$ itself (Section II.2). We then compute the ensemble mean response of the net radiative flux at the TOA at the next time step $t=1$, referred here as $\delta\langle x_{\text{TOA}}(1)\rangle$. To estimate the ensemble mean response we consider $10^{3}$ initial conditions. We write this response as follows:
$\begin{split}R_{\text{TOA},j}(1)=\frac{\delta\langle x_{\text{TOA}}(1)\rangle}{\Delta_{j}\delta(t)},\end{split}$ (12)
ad remind the reader that the index $j$ only considers the SST degrees of freedom. Therefore, the response at 1-time-step $R_{\text{TOA},j}(1)$ defines the direct causal link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$ between the $j$-th degree of freedom in SST and the net radiative flux at the TOA.
The vector $\mathbf{R_{\text{TOA}}}(1)\in\mathbb{R}^{n}$ is projected back to the higher, original $N$-dimensional space. To do so we consider the previously computed EOFs, i.e. the orthonormal vectors $\mathbf{e}\in\mathbb{R}^{n,N}$ and project the 1-time-step responses as follows:
$\begin{split}\mathbf{S}=\mathbf{R_{\text{TOA}}}(1)\cdot\mathbf{e}=\sum_{j}^{n}R_{\text{TOA},j}(1)~{}\mathbf{e}_{j},\end{split}$ (13)
where $\mathbf{S}\in\mathbb{R}^{N}$ is a sensitivity map defined in the high-dimensional $N$ dimensional space. Such map approximates at each grid point $i$ in the ocean, the sensitivity $\frac{\partial x_{\text{TOA}}}{\partial x_{\text{SST},i}}$, where $x_{\text{TOA}}$ is the global mean radiative flux at the TOA and $x_{\text{SST},i}$ is the SST at grid point $i$.
The sensitivity map $\mathbf{S}$ obtained using the neural emulator in Eq. (10) is shown in Figure 7. The obtained sensitivity map is in qualitative agreement with previous results in response theory (Figure 2(a) in [105]), with other data-driven approaches (see for example Figure 7(b) in [120] and Figure 2 in [121]), but, most importantly, with the sensitivity map obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [29], see Figure 5(a) in [30]). Note that in the latter case, the maps exhibit fewer spatial features due to the coarse resolution of the perturbation patches, which span tens of degrees. The marked feature of this sensitivity map is the sensitivity dipole found the tropical Pacific, with negative values on the western side of the basin and positive values on the eastern side, see for example [30]. This means that an initial warming in the eastern Pacific basin will result in a direct, positive radiative feedback, further amplifying the imposed temperature change; the opposite is true for the western Pacific. The mechanistic explanations for such mechanisms can be found for example in [122, 111] and we refer the interested reader to these papers for details. In brief, the positive sensitivity in $\mathbf{S}$ in the eastern Pacific is associated with a high-climatology of low-level clouds: a warming perturbation is expected to decrease the cloud cover, leading to anomalies in incoming shortwave radiation and therefore a positive sensitivity. On the other hand, the western Pacific is a region dominated by deep convective phenomena. Because of this, a surface temperature warming in the western Pacific basin tend to travel vertically through the troposphere and then horizontally through gravity waves. This process strengthen the inversion layer in the eastern Pacific basin (and in other regions with high-climatology of low-level clouds), leading to a larger low-level cloud fraction and increasing the reflected shortwave at the top of the atmosphere.
To conclude this Section, the analysis in Figure 7 confirms that the good reconstruction observed in the forced regime (Figure 6(e,f)) arises for the physically correct reasons.
Figure 7: Sensitivity map obtained using the neural emulator in Eq. (10). At each grid point $i$ we are plotting the 1 time step response in the global mean, net radiative flux at the top of the atmosphere (TOA) driven by an impulse perturbation in the sea surface temperature (SST) at that grid point. Note: the computation is done in a 21 dimensional space spanned by a set of empirical orthogonal functions and this map is a projection of the results obtained in the low-dimensional space. The response has been obtained by averaging over an ensemble of $10^{3}$ initial conditions.
IV.2.4 Main takeaways
In this Section, we developed a stochastic neural emulator that captures the joint evolution of the sea surface temperature (SST) field and the global mean net radiative flux at the top of the atmosphere (TOA) in the GFDL-CM4 climate model. Addressing a specific, well-posed problem, proved crucial in guiding the construction of the model: it enabled us to build on previous literature, restrict attention to just two variables and a limited range of spatial and temporal scales. These simplifications led to a stable emulator that responds to external forcing and can be used to investigate causal relationships via perturbation experiments. The model reconstruct the net TOA radiative flux when forced with a pre-computed SST trajectory. Beyond reconstruction, we illustrated that the model can be used to perform perturbation experiment, allowing us to enhance mechanistic understanding of the underlying physical processes. Findings are in qualitative agreement with results from traditional perturbation experiments in numerical models, giving additional confidence in the emulator’s ability to capture key interactions across variables.
V Conclusions
In the first part of this paper, we addressed, at a pedagogical level, foundational issues in neural emulation of dynamical systems with climate applications in mind. We considered two scenarios: (i) when the full state vector is observed, and (ii) the more realistic case of partial observations. Our findings can be summarized as follows:
  1. i.
    Fully observed state vector. In this idealized case, it is possible to infer a dynamical model from data alone that reproduces the system’s stationary statistics and responds accurately to small, external perturbations.
  2. ii.
    Partially observed state vector. This more realistic case presents fundamental difficulties, even with access to high-quality data and powerful neural models architectures. The model skill now depends critically on two factors: (a) the choice of variables to model and (b) how unobserved degrees of freedom are parameterized.
These results highlight intrinsic limitations of purely data-driven modeling in high-dimensional systems, where the full set of relevant variables is typically unknown and there is no rigorous mathematical framework for selecting the proper variables to model. The identification of such variables, when possible, is inherently problem-dependent and often leads to adopting a coarse-grained representation of the dynamics. The next challenge is to parameterize the cumulative effect of unobserved degrees of freedom. This final step underscores that the issue of parameterization, traditionally associated with numerical modeling, is equally central to data-driven approaches using neural emulators. A promising perspective is to reconcile neural networks with principled modeling strategies, such as reduced-order modeling and stochastic parameterizations, by focusing on coarse-grained dynamics rather than constructing deterministic models that attempt to resolve as many spatiotemporal scales as possible.
In the second part of the paper, we built on these insights by constructing a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net top-of-atmosphere (TOA) radiative flux. The model was trained on long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies. We showed that the emulator reproduces the stationary statistics of the system and, more importantly, allows us to reconstruct TOA radiative fluxes in forced simulations when provided with forced SST trajectories. To deepen this analysis, we applied linear response theory to extract the direct causal links between temperature modes and TOA fluxes, targeting the so-called “pattern effect” [30, 29]. These links, when projected back onto the high-dimensional state space, were found to be consistent with recent studies. This suggests that the emulator’s skill in reconstructing TOA fluxes under forcing arises for the right physical reasons. In this way, focusing on a few dominant modes, on a limited range of time scales, and using stochastic parametrizations allowed us to build a skillful and interpretable emulator.
Modeling real-world dynamical systems from data alone is severely constrained by high dimensionality and the fact that we typically observe only a subset of relevant variables [46, 32, 40]. This leads to the fundamental challenge of identifying, when possible, the proper variables to model, a problem recognized since the work of Onsager and Machlup (1953) [68] and highlighted in more recent studies [47, 48]. These challenges are especially pronounced in climate dynamics, where observations are always partial, and the underlying system is both high-dimensional and multiscale. In this context, the multiscale structure of the climate system can be an advantage, allowing us to model coarse-grained dynamics while treating unresolved processes stochastically. The need to formulate models tackling the effective dynamics of slow variables, rather than the full set of primitive equations, has a long history in climate science, even in non-data-driven settings where all governing equations are known. The early work of Charney and von Neumann, for example, demonstrated the conceptual and practical benefits of filtering out “meteorologically insignificant noise” and emphasized the need for interpretable and computationally efficient effective equations in metereology [123, 63, 43, 44]. These insights remain broadly relevant when developing neural emulators, especially in future attempts of building atmosphere-ocean coupled emulators. In purely data-driven settings, focusing on stochastic models that target coarse-grained processes offers a promising way forward by reconciling rigorous approaches such as reduced-order modeling with the flexibility and power of neural networks. Possible strategies include learning stochastic parameterizations directly from data using deep learning (as in this paper), guiding variable selection through causal inference [124, 125], incorporating information theory and response theory into the training phase [56], studying the evolution of the climate system in its “latent” dynamics [65] (i.e. on the attractor) and incorporating empirical knowledge and physics constraints in the learning phase [11, 7, 126, 25]. Finally, the rather hard difficulty (at times irreducible) of data-driven modeling for the coupled atmosphere–ocean system can at times be ameliorated by building models tailored to solve specific problems, by building on prior literature and insights, and developed with a clear understanding of the relevant variables and the spatiotemporal scales at play.
Appendix A Technical and practical details on the model formulations
A.1 Multivariate data-driven stochastic models
Consider a $N$-dimensional dynamical system described by a stochastic differential equation $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W}$. This equation can be discretized by the forward Euler method as follows:
$\begin{split}\Delta\mathbf{x}=\mathbf{x}(t+\Delta t)-\mathbf{x}(t)=\mathbf{f}(\mathbf{x}(t))\Delta t+\mathbf{\Sigma}(\mathbf{x})\sqrt{\Delta t}\bm{\xi}(t)~{}\end{split}$ (14)
where $\bm{\xi}(t)$ is a Gaussian white noise with zero mean, unit variance, and $\sqrt{\Delta t}\bm{\xi}(t)$ is the increment of the Wiener process $\Delta\mathbf{W}=\mathbf{W}(t+\Delta t)-\mathbf{W}(t)$. Without loss of generality, we rescale time units by the time sampling interval such that a non-dimensional time increment is $\Delta t=1$ [6, 127, 7, 8]. The strategy is then to fit a rescaled version of Eq. 14 from observational data alone. Such rescaled equation takes the following form:
$\begin{split}\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)=\mathbf{f}(\mathbf{x}(t))+\mathbf{\Sigma}(\mathbf{x})\bm{\xi}(t).\end{split}$ (15)
We proceed in two steps:
  1. i.
    Fitting the drift. We compute $\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)$ and fit the deterministic dynamics such that $\Delta\mathbf{x}\approx\mathbf{f}(\mathbf{x}(t))$. In the main paper, we considered two cases:
    • Linear drift. In this case, given a trajectory $\mathbf{x}(t)\in\mathbb{R}^{N}$, we need to fit a matrix $\mathbf{L}\in\mathbb{R}^{N,N}$ such that $\Delta\mathbf{x}\approx\mathbf{L}\mathbf{x}(t)$ as close as possible. In practice, this is solved as a linear regression problem by minimizing the mean squared error. We use the linear regression module from the scikit-learn library [128].
    • Nonlinear drift. We consider the following form: $\Delta\mathbf{x}\approx\mathbf{f}(\mathbf{x}(t))=\mathbf{L}\mathbf{x}(t)+\mathbf{n}(\mathbf{x}(t))$, where $\mathbf{L}\in\mathbb{R}^{N,N}$ is the linear operator and $\mathbf{n}(\mathbf{x}(t))$ is the nonlinear operator parameterized by a multilayer perceptron (MLP). Both operators are fitted jointly by minimizing the mean squared error using a version of the gradient descent algorithm (Adam, [129]) with weight decay. The MLP considered here always consisted of one hidden layer. The number of neurons in the hidden layer depended on the application: we considered 30 neurons in the experiments with the triad model in Section III.2.1 and 1000 neurons in the real-world application in Section IV. We played with more hidden layers but found a comparable performance, and sometimes a larger network also led to instabilities. The activation function used is the swish function (SiLU), which we found yields smoother responses to external perturbations. In contrast, the commonly used ReLU activation produced noisier responses. Furthermore, to ensure high accuracy in perturbation-response experiments, all data were processed using double-precision floating-point format.
  2. ii.
    Fitting the stochastic term. Having fitted the deterministic dynamics $\mathbf{f}$, we compute the residuals $\mathbf{r}(t)\in\mathbb{R}^{N}$ as $\mathbf{r}(t)=\Delta\mathbf{x}-\mathbf{f}(\mathbf{x}(t))$. We then construct a suitable stochastic parametrization for this residual term. In the paper, we focus on two cases: additive and multiplicative noise.
    • Additive noise. Given the residuals $\mathbf{r}(t)$, we proceed by computing their covariance matrix $\mathbf{C}\in\mathbb{R}^{N,N}$ as $\mathbf{C}=\langle\mathbf{r}\mathbf{r}^{\text{T}}\rangle$, where brackets denote time averaging. We then consider the Cholesky decomposition of $\mathbf{C}$ to derive a lower triangular matrix $\mathbf{\Sigma}\in\mathbb{R}^{N,N}$ such that $\mathbf{C}=\mathbf{\Sigma}\mathbf{\Sigma}^{\text{T}}$. The full fitted equation solved by the emulator can then be written as:
      $\begin{split}\mathbf{x}(t+1)=\mathbf{x}(t)+\mathbf{f}(\mathbf{x}(t))+\mathbf{\Sigma}\bm{\xi}(t)~{}\end{split}$ (16)
      and we remind the reader that $\bm{\xi}(t)\in\mathbb{R}^{N}$ is a Gaussian white noise with zero mean, unit variance.
    • Multiplicative noise. Here, we consider state-dependent noise. To approximate it in a computationally efficient manner, we proceed as follows:
      • Given the residuals $\mathbf{r}(t)$ at each time $t$, we compute their outer product $\mathbf{C}(t)=\mathbf{r}(t)\mathbf{r}(t)^{\text{T}}$ which is a matrix of rank one.
      • We train a neural network predicting a lower triangular matrix $\mathbf{\Sigma}(\mathbf{x}(t))$ with positive diagonal, which can be interpreted as a Cholesky factor, such that $\mathbf{C}(t)\approx\mathbf{\Sigma}(\mathbf{x}(t))\mathbf{\Sigma}(\mathbf{x}(t))^{\text{T}}$ as close as possible.
      In practice, the loss function minimizes the mean squared error (MSE) between $\mathbf{C}(t)$ and $\mathbf{\Sigma}(\mathbf{x}(t))\mathbf{\Sigma}(\mathbf{x}(t))^{\text{T}}$ at each batch. A minimizer of the MSE loss approximates the conditional mathematical expectation $\mathbb{E}[\mathbf{C}(t)|\mathbf{x}(t)$], that is the covariance matrix averaged over all residuals corresponding to the same state vector, as a function of this state vector, and can be seen as a generalization of a method presented in [130] to multiple dimensions. The neural network architecture used here is identical to the one used for fitting the drift. The positivity of the diagonal terms of the lower triangular matrix $\mathbf{\Sigma}(\mathbf{x}(t))$ is enforced through a softplus activation function. Note that this (high-dimensional) multiplicative noise parametrization is only applied in the real-world-like case, where data are scarce: to prevent overfitting, we use early stopping during training, halting after only 5 epochs. The final fitted equation solved by the emulator is:
      $\begin{split}\mathbf{x}(t+1)=\mathbf{x}(t)+\mathbf{f}(\mathbf{x}(t))+\mathbf{\Sigma}(\mathbf{x}(t))\bm{\xi}(t)~{}\end{split}$ (17)
      where again $\bm{\xi}(t)\in\mathbb{R}^{N}$ is a Gaussian white noise with zero mean, unit variance.
Furthermore, even in the real-world case of sea surface temperature fields and radiative fluxes, reducing the dimensionality to a few, $N=21$, components led to a simplified computational problem where all neural networks could be trained on CPUs and on a personal computer.
A.2 Scalar data-driven stochastic models
Consider a trajectory of length $T$ of a $1$-dimensional dynamical system. As in the Section above, we assume that this scalar time series can be modeled as a discrete stochastic process with time-stepping $\Delta t=1$, therefore rescaling our time-stepping by the sampling interval. The scalar equation we aim to fit is therefore:
$\begin{split}\Delta x=x(t+1)-x(t)=f(x(t))+\sigma(x)\xi(t).\end{split}$ (18)
The training of the deterministic (i.e. drift) dynamics follows exactly the same procedure of the multivariate case. Even the network considered are the same, with one hidden layer and 30 neurons. So we refer the reader to the Section above regarding the fitting of the deterministic dynamics and proceed in focusing on the stochastic case.
Fitting the stochastic term.
  • Additive noise. We compute the residual time series $r(t)=\Delta x-f(x(t))$ from the drift term. We then proceed in computing the standard deviation $\sigma$ of such residual time series. The full fitted equation solved by the emulator can then be written as:
    $\begin{split}x(t+1)=x(t)+f(x(t))+\sigma\xi(t)~{}\end{split}$ (19)
    and we remind the reader that $\xi(t)$ is a scalar Gaussian white noise with zero mean, unit variance.
  • Multiplicative noise. We again consider the residual time series $r(t)$ from the drift term. We then compute at each time step $t$ the square of the residual at that time step, i.e. $r(t)^{2}$. We then train a neural network to have $g(x(t))\approx r(t)^{2}$ by minimizing the mean squared error [130]. The neural network is again a MLP with one hidden layer with 30 neurons. Differently, from the neural networks considered for the drift, here we added a softplus activation function after the last layer of the network to ensure that the variance estimates are nonnegative. The final fitted equation solved by the emulator is:
    $\begin{split}x(t+1)=x(t)+f(x(t))+\sqrt{g(x(t))}\xi(t).\end{split}$ (20)
Acknowledgements.
Most of the ideas discussed in this work stem from collaborations and exchanges with colleagues and friends. I am grateful for the supportive and engaging environment of the Zanna Group at Courant, as well as for interactions within the broader M2Lines project. I thank Laure Zanna and Carlos Fernandez-Granda for their feedback during early stages of the project. I am especially grateful to Pasha Perezhogin, Rory Basinski, Chris Pedersen, and Andre Souza for their thoughtful feedback and for many enriching discussions on data-driven modeling of the Climate system. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.

### v2

@@
 By real-world examples we refer to situations where (i) we observe only the temporal evolution of a subset of variables without access to the full state vector (i.e.
-a projection of the full dynamics) or (ii) where even in the presence of all variables the system encompasses a huge number of spatial and temporal scales (as for the climate and weather system).
+a projection of the full dynamics) or (ii) where even in the presence of all variables the system encompasses a huge number of spatial and temporal scales.
 Both cases (i) and (ii) lead to the need to formulate effective rather than primitive equations for both practical and conceptual reasons [40].
@@
 However, even when multiplicative noise is considered, the responses to impulse perturbations are still only qualitatively right.
-The example illustrates that accurately capturing stationary statistics does not guarantee faithful reproduction of perturbed statistics.
+We emphasize two key points:
+  1. i.
+    The analysis in Figure 4 highlights the importance of evaluating a model’s ability in predicting responses in higher-order moments, such as the variance.
+In the simple example discussed here, focusing solely on the ensemble mean response would lead to the incorrect conclusion that the two models respond to perturbations in the same way (see panels (b) and (e) in Figure 4).
+  2. ii.
+    Matching the stationary distribution quantitatively does not guarantee equally accurate responses to perturbations, as shown in Figure 4(d,e,f).
 Figure 4: First row: stationary probability distribution (panel (a)); ensemble mean response to an impulse perturbation (panel (b)); ensemble variance response to an impulse perturbation (panel (c)) for the variable $x_{1}$ in the triad system in Eq.
@@
 Ultimately, the remaining residuals become white in time, allowing the system to be closed using a Wiener process.
-Similarly, it could be possible to address the issue through recurrent neural networks [103].
 An alternative, practical approach for handling non-Markovian effects is to coarse-grain the data in the temporal dimension, for instance, by averaging over a time window.
-This suppresses fast fluctuations within that window leading to an effective “Markovianization” of the the dynamics.
-Due to its simplicity and ease of implementation, this method is commonly used in the literature (see for example [4]) and is also adopted in the present study.
+This suppresses fast fluctuations within that window leading to an effective “Markovianization” of the dynamics.
+Due to its simplicity and practicality, this method is commonly used in the literature (see e.g.
+[4]) and is also adopted in the present study.
 We therefore consider again the time series $x_{1}(t)$ from the long simulation of the triad system in Eq.
@@
 In contrast, the model with multiplicative noise not only captures the stationary distribution and further improves the mean response, but now also closely reproduces the response in variance, capturing the peak magnitude and deviating only slightly in its temporal decay.
-Note that the stationary response in ensemble mean and variance to impulse perturbations of the original degree of freedom $x_{1}$ are the same in both Figures 4 and 5, the only thing changing in the latter is that the reduced-order model is defined in a coarse-grained dynamics.
+Note that the response in ensemble mean and variance to impulse perturbations of the original degree of freedom $x_{1}$ are the same in both Figures 4 and 5, the only thing changing in the latter is that the reduced-order model is defined in a coarse-grained dynamics.
 Figure 5: Same as in Figure 4 however here the scalar emulators in Eq.
@@
 This leads to two major challenges: (i) the resulting data-driven model will depend heavily on the specific variable chosen for modeling, and (ii) once a suitable slow variable is identified, the cumulative effect of the unobserved degrees of freedom must be appropriately parametrized.
-These issues, demonstrated at the “practical level” within the context of a low-dimensional model, become even more severe in realistic high-dimensional systems, which often lack a clear time-scale separation.
+These issues, demonstrated at the “practical level” within the context of a low-dimensional model, become more severe in realistic high-dimensional systems, which often lack a clear time-scale separation.
 These challenges are intrinsic to data-driven approaches for dynamical systems, and simply extending time series length or training increasingly large neural networks is unlikely to resolve them.
-The core difficulty is conceptual: there is no rigorous mathematical framework for selecting the proper variables to model in real-world, high-dimensional dynamical systems.
-This poses fundamental limitations for purely data-driven approaches.
-In practice, such choices rely on physical intuition and prior knowledge.
-When possible, selecting the right variables amounts to identifying an appropriate coarse-grained description of the dynamics, including relevant fields and suitable spatial and temporal scales.
-Consequently, effective data-driven modeling of real-world complex systems hinges on deep physical insight and strong domain expertise.
+Crucially, proposing more sophisticated models can be effective only if the proper variables are first identified.
+The core difficulty is conceptual: there is no rigorous mathematical framework for selecting the proper variables to model from data in real-world, high-dimensional dynamical systems.
+In practice, such choices rely on physical insight and domain expertise.
+In a spatiotemporal system like climate, this involves identifying an appropriate coarse-grained description of the dynamics, including relevant fields and a limited range of spatial and temporal scales.
+In the context of climate dynamics, it is sometimes possible (depending on the specific problem) to leverage the governing equations to guide variable selection.
+However, even in this case, one faces the practical necessity of resolving only a limited range of scales and introducing parametrizations.
+Within this framework, stochastic emulators offer a promising approach for data-driven simulation of climate systems and their sensitivity to external perturbations.
 IV A real-world example: radiative responses to temperature perturbations in a climate model
@@
 Data.
-The focus in on a state-of-the-art coupled climate model GFDL-CM4 [112].
+The focus is on a state-of-the-art coupled climate model GFDL-CM4 [111].
 The model offers data from a long control run simulating the coupled climate system for 600 years with constant $\text{CO}_{2}$ fixed at pre-industrial level.
@@
 Stationary statistics.
-Given the $n+1$ state vector $\mathbf{x}(t)\in\mathbb{R}^{n+1}$ defined in the preprocessing step above, we can now train the emulator defined in Eq.
+Given the $n+1$ dimensional state vector $\mathbf{x}(t)\in\mathbb{R}^{n+1}$ defined in the preprocessing step above, we can now train the emulator defined in Eq.
 (10) on the long control run of the GFDL-CM4 model.
@@
 Addressing a specific, well-posed problem, proved crucial in guiding the construction of the model: it enabled us to build on previous literature, restrict attention to just two variables and a limited range of spatial and temporal scales.
-These simplifications led to a stable emulator that responds to external forcing and can be used to investigate causal relationships via perturbation experiments.
+These simplifications led to a stable and computationally efficient emulator that responds to external forcing and can be used to investigate causal relationships via perturbation experiments.
 The model reconstruct the net TOA radiative flux when forced with a pre-computed SST trajectory.
@@
 V Conclusions
-In the first part of this paper, we addressed, at a pedagogical level, foundational issues in neural emulation of dynamical systems with climate applications in mind.
+In the first part of this paper, we addressed, at a pedagogical level, fundamental issues in neural emulation of dynamical systems with climate applications in mind.
 We considered two scenarios: (i) when the full state vector is observed, and (ii) the more realistic case of partial observations.
@@
 A promising perspective is to reconcile neural networks with principled modeling strategies, such as reduced-order modeling and stochastic parameterizations, by focusing on coarse-grained dynamics rather than constructing deterministic models that attempt to resolve as many spatiotemporal scales as possible.
-In the second part of the paper, we built on these insights by constructing a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net top-of-atmosphere (TOA) radiative flux.
+In addition, we argued that impulse response functions are a building block that a model should capture in order to reproduce causal interactions across degrees of freedom and to accurately respond to external perturbations.
+This perspective enables us to move beyond evaluation metrics based solely on stationary statistics and instead consider the effect of external interventions.
+Assessing impulse responses is straightforward in the case of data-driven emulators built from simple numerical models, i.e.
+models that allow us to simulate thousands of ensemble members, as in the first part of this study.
+In such cases, one can actively perturb the reference model to generate a ground truth for the response operator.
+However, for models constructed from observational data, such direct perturbations are not possible.
+In this context, one could, in principle, leverage the Fluctuation-Dissipation Theorem (FDT) to estimate impulse response functions from data alone, at least at a qualitative level [38, 37, 35, 47, 62].
+In particular, it has been shown that even simple approximations of FDT, such as the quasi-Gaussian approximation [37], often yield qualitatively reliable estimates of impulse responses in the ensemble mean, without the need to perturb the system directly; see, for example, [122, 123, 104].
+That said, FDT itself is inferred from data and should therefore be viewed as a complementary, theory-based diagnostic (rather than a ground truth) allowing us to cross-validate the inferred responses of the emulator.
+In other words, consistency between the responses predicted by the emulator and those estimated via FDT can increase confidence in the model’s ability to capture the system’s response in the ensemble mean to external perturbations.
+In contrast, assessing impulse responses of higher-order moments (e.g., variance) via FDT remains substantially more challenging and calls for new methodological tools.
+A promising ongoing direction involves score-based generative modeling, as explored in [124, 125].
+Another fruitful avenue lies in data-driven frameworks at the intersection of Markov State Modeling and Koopman operator theory [126, 127].
+It is worth emphasizing that the central limitation discussed earlier, i.e.
+the need to choose the proper variables, remains relevant for the inference of response operators via FDT as well (and for any causal inference methodology); see the discussion in [47].
+Building on the theoretical insights introduced in the first part, the second part of the paper presents a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net top-of-atmosphere (TOA) radiative flux.
 The model was trained on long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies.
@@
 In this way, focusing on a few dominant modes, on a limited range of time scales, and using stochastic parametrizations allowed us to build a skillful and interpretable emulator.
+We emphasize that the fundamental limitations discussed here apply to data-driven models of complex dynamical systems that must both represent stationary statistics and respond faithfully to external perturbations.
+For Earth system models specifically, these considerations primarily concern climate emulators (particularly coupled atmosphere-ocean systems), and should not be conflated with the different challenge of weather forecasting using neural networks (e.g., [128]).
+Weather prediction focuses on short-term predictability where long-term stability and forced responses are of secondary importance.
 Modeling real-world dynamical systems from data alone is severely constrained by high dimensionality and the fact that we typically observe only a subset of relevant variables [46, 32, 40].
@@
 The early work of Charney and von Neumann, for example, demonstrated the conceptual and practical benefits of filtering out “meteorologically insignificant noise” and emphasized the need for interpretable and computationally efficient effective equations in metereology [129, 63, 43, 44].
-These insights remain broadly relevant when developing neural emulators, especially in future attempts of building atmosphere-ocean coupled emulators.
+These insights remain conceptually relevant, especially the focus on effective dynamics, in the development of neural emulators for climate, including coupled atmosphere–ocean systems.
 In purely data-driven settings, focusing on stochastic models that target coarse-grained processes offers a promising way forward by reconciling rigorous approaches such as reduced-order modeling with the flexibility and power of neural networks.

### v3

@@
-Neural models of multiscale systems: conceptual limitations, stochastic parametrizations, and a climate application
+Probing forced responses and causal mechanisms in large-scale climate dynamics with reduced-order neural models
 Abstract
-This work explores key conceptual limitations in data-driven modeling of multiscale dynamical systems, focusing on neural emulators and stochastic climate modeling.
-A skillful climate model should capture both stationary statistics and responses to external perturbations.
-While current autoregressive neural models often reproduce the former, they typically struggle with the latter.
-We begin by analyzing a low-dimensional dynamical system to expose, by analogy, fundamental limitations that persist in high-dimensional settings.
-Specifically, we construct neural stochastic models under two scenarios: one where the full state vector is observed, and another with only partial observations (i.e.
-a subset of variables).
-In the first case, the models accurately capture both equilibrium statistics and forced responses in ensemble mean and variance.
-In the more realistic case of partial observations, two key challenges emerge: (i) identifying the proper variables to model, and (ii) parameterizing the influence of unobserved degrees of freedom.
-These issues are not specific to neural networks but reflect fundamental limitations of data-driven modeling and the need to target the slow dynamics of the system.
-We argue that physically grounded strategies – such as coarse-graining and stochastic parameterizations – are critical, both conceptually and practically, for the skillful emulation of complex systems like the coupled climate system.
-Building on these insights, we turn to a more realistic application: a stochastic reduced neural model of the sea surface temperature field and the net radiative flux at the top of the atmosphere, assessing its stationary statistics, response to temperature forcing, and interpretability.
+A central challenge in climate science and applied mathematics is developing data-driven models of multiscale systems that capture not only stationary statistics but also responses to external perturbations.
+Current neural climate emulators seek to resolve the atmosphere–ocean system in all its complexity but often fail to reproduce forced responses, limiting their use in causal studies such as Green’s function experiments.
+Response theory offers a rigorous framework to explore these limitations, moving beyond stationary statistics, probing causal mechanisms, and guiding model design.
+Using a simplified multiscale system, we argue that the skill of purely data-driven models in reproducing perturbed statistics depends critically on (i) identifying the proper variables to model, including spatial and temporal scales, and (ii) accurate stochastic parameterizations of unresolved processes.
+For studies of low-frequency climate dynamics, these insights motivate reduced-order neural models, tailored to specific processes, as valuable alternatives to general-purpose emulators.
+We illustrate this perspective with a reduced-order neural stochastic model designed to investigate responses of radiative fluxes to surface temperature perturbations.
+The model largely reproduces stationary statistics and enables the study of forced responses in both ensemble mean and variance, shifting the study of climate feedbacks from single trajectories to responses of probability distributions.
+These results underscore reduced-order models, augmented with modern neural networks and evaluated within the framework of response theory, as a practical strategy for causal inference and attribution studies in large-scale climate dynamics.
 preprint: APS/123-QED
 I Introduction
-A key challenge in current applied mathematics is to build data-driven models from partial observations of complex physical systems.
+A key challenge in current applied mathematics is to build data-driven models from partial observations of multiscale physical systems.
 Among the most challenging examples is the climate system: a forced, dissipative, and chaotic system whose dynamics span at least 15 orders of magnitude across spatial and temporal scales [1].
@@
 [10, 11].
-More recent proposal also include dynamical Markov-chain models to simulate the evolution of probability distributions as in [12, 13, 14] as well as novel ideas from generative models [15].
-Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of the weather [16, 17, 18] and climate [19, 20, 21, 22, 23, 24, 25] systems.
-These tools build on earlier efforts in data-driven reduced-order modeling by considering new powerful architectures (i.e.
-neural networks) and by attempting to simulate the system in its full complexity and dimensionality.
-For example, [24] and [25] respectively seek to emulate the full ocean and atmosphere components of two state-of-the-art climate models.
-These models often excel at reproducing stationary statistics and generating forecasts across a range of time horizons, marking a promising direction for computationally efficient climate modeling.
-However, neural network emulators also exhibit several limitations.
-Long trajectory rollouts often suffer from error accumulation, which can drive the system toward instability.
-While promising approaches have been proposed to address this issue, e.g.
-[26, 27], these are currently research directions still to be explored across different models and scenarios.
-Instead current implementations maintain stability by prescribing part of the climate system as boundary conditions (e.g.
-[24] is an ocean emulator driven by pre-computed atmospheric fields as boundary conditions).
-We refer the reader to Pedersen et al.
-[26] for a detailed discussion of the stability challenges in autoregressive modeling.
-An additional issue, relevant for this study, is that skill in capturing unforced variability may not always guarantee accurate responses in perturbation experiments, limiting their use for causal mechanistic studies of climate dynamics.
-This leads to two important consequences: neural emulators (i) usually do not generalize in out-of-distribution climate change experiments and (ii) they may often not capture the right causal relations across variables.
-Note that the two issues (i) and (ii) above are likely related.
+Recent proposal include Markov-chain models as in [12, 13, 14] and novel ideas from generative models [15].
+Most prior work has focused on evaluating reduced-order models on equilibrium statistics and forecasting, with limited attention to responses to external perturbations, an exception being the contributions of Majda and collaborators (see e.g.
+[16]).
+Capturing such responses is key for building models that not only mimic the system’s dynamics but also faithfully capture causal relations across degrees of freedom [17, 18].
+Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of complex turbulent systems such as the weather [19, 20, 21, 22, 23] and climate [24, 25, 26, 27, 28, 29, 30, 31, 32, 33] systems.
+Unlike traditional reduced-order models, these neural architectures aim to emulate the full hierarchy of spatial and temporal scales present in the system.
+For example, [30] and [31] emulate the full ocean and atmosphere components of state-of-the-art climate models.
+Such emulators often reproduce stationary statistics and generate forecasts across a range of horizons, offering a promising direction for efficient climate modeling.
+However, several limitations remain.
+Long rollouts suffer from error accumulation that can drive instabilities [34, 35].
+Moreover, current emulators are unable to capture extreme events if not already present in the training set, preventing their use for rare-event analysis in a changing climate [36].
+More relevant here, skill in reproducing stationary statistics does not guarantee accurate responses in perturbation experiments, restricting their use for causal studies of climate dynamics.
+This has two consequences: (i) neural emulators often fail to generalize in out-of-distribution climate change scenarios, and (ii) they may not capture the correct causal relations across variables, two issues that are likely related.
 The work of Van Loon et al.
-[28] provides us with a clear and insightful example of this limitation in the context of the “pattern effect” problem [29, 30], i.e.
-the influence of sea surface temperature (SST) warming patterns on the Earth’s energy imbalance.
-Van Loon et al.
-[28] considered the ACE2 emulator [21] and built a Green’s function by iteratively applying warming perturbations in different locations in the sea surface temperature (SST) field and computing the responses in the radiative flux at the top of the atmosphere (TOA).
-In doing so, [28] showed how such Green’s function does not capture in a quantitative way the causal relations between the SST and the radiative fluxes.
-This is just one example of the limitations of neural emulators in capturing responses to perturbations.
-Even if such issues are addressed in specific cases, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent.
-Other examples related to capturing forced responses to changes in the boundary conditions can be found in [24, 25].
-Additionally, we note that current emulators usually focus on one climate component at a time, ocean or atmosphere, and the challenges described above, and also later in this paper, are expected to get worse when a fully coupled system will be considered.
-In general, limitations of neural emulators in capturing responses to external perturbations imply that such tools may not be yet ready to explore causal questions in climate dynamics.
+[37] illustrates this limitation in the context of the “pattern effect” problem [38, 39], i.e.
+the influence of sea surface temperature (SST) warming patterns on Earth’s energy imbalance.
+Using the ACE2 emulator [26] trained on ERA5 [40], they constructed a Green’s function by perturbing SSTs regionally and measuring the radiative response.
+The resulting Green’s function failed to capture the causal relation between SST and radiative fluxes in a quantitative way.
+This is just one example of limitations of neural emulators in capturing responses to perturbations with additional examples in [30, 31, 41].
+Even if such issues are addressed in specific cases, see [42] in the context of the pattern effect, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent.
+Moreover, most current emulators represent only a subsystem (e.g., atmosphere or ocean), and limitations in responding to perturbations are expected to worsen in coupled models.
+Overall, limitations in capturing forced responses suggest that neural emulators are not yet ready for causal applications in climate dynamics.
 While the field of neural emulation is advancing rapidly, it is useful to step back and examine its fundamental limitations in capturing forced responses, both from a theoretical standpoint and in controlled settings.
@@
 Vulpiani and collaborators on the limitations of data-driven modeling [43, 44, 45].
-We analyze a simplified stochastic dynamical system to illustrate fundamental limitations that are expected to persist in high-dimensional settings.
-Specifically, we ask the following question: Can neural autoregressive models, trained on (practically) infinitely long time series, accurately capture the response to general external perturbations?
-In addressing this question, we consider two scenarios:
-  • Fully observed state vector.
+We consider a simplified stochastic dynamical system and ask the following question: Can neural autoregressive models, trained on (practically) infinitely long time series, accurately capture the response to general external perturbations?
+In addressing this question, we focus on two relevant scenarios in data-driven-only modeling:
+  • Unknown equations, fully observed state vector.
 Data-driven modeling when the entire state vector is available.
 Here, we assume no knowledge of the underlying equations, but we have access to very long time series of the complete system state vector.
-  • Partially observed state vector.
+  • Unknown equations, partially observed state vector.
 Data-driven modeling in the presence of partial observations.
@@
 In this more realistic setting, we again assume no knowledge of the underlying equations and only partial access to the system’s state vector.
-To tackle such questions we proceed pedagocially by focusing on a “triad model” first proposed by Majda and collaborators in [34] and used as a testbed for Fluctuation-Dissipation theorem strategies in [35, 36].
-The model mimics the abstract structure of dynamical cores in large-scale climate models with dynamics characterized by a linear operator and an energy conserving nonlinear, quadratic operator.
+To tackle such questions we proceed pedagogically by focusing on a multiscale “triad model” first proposed by Majda and collaborators in [46, 47].
+The model mimics the abstract structure of dynamical cores in large-scale climate models with a linear operator and an energy conserving nonlinear, quadratic operator.
+It provides a simplified yet relevant mathematical framework to unambiguously explore key limitations in capturing forced responses in climate emulators.
 We address the questions above by training neural stochastic emulators on a long integration of the triad model and test the models’ performance on stationary and perturbed cases.
@@
 Nonetheless, the linear response to a general forcing can be computed by convolving the forcing with an impulse response operator [48, 49].
-Such operator can be retrieved from the model, and is here considered as the building block to get right for an emulator to, at least, capture the linear part of the response to general external perturbations.
-In the case of a fully observed state vector we show that the inferred neural model is able to reproduce stationary statistics of the system, i.e.
-the probability distributions and autocorrelation functions, as well as the response to small external perturbations.
-Therefore, there is no fundamental limitation for a neural model to respond to external perturbation, i.e.
-shifts in the input distribution, at least for perturbations of small amplitudes.
-However, in the case of partial observations, as in more realistic settings, we argue how the performance of a data-driven model necessarily depends on (i) the choice of variables to model and (ii) how to parametrize the unobserved degrees of freedom.
-The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [5, 39, 40, 41].
-Therefore, we will argue that a promising perspective is to build on such established approaches, emphasizing coarse-grained dynamics and stochastic parameterizations, while using neural networks as powerful tools within this framework, rather than as wholesale replacements.
+This operator can be inferred from the model and is a building block that an emulator must capture to reproduce the response to small, general external perturbations [48], as well as the cause–effect relations among variables [17].
+In the case of a fully observed state vector we show that the inferred neural model can reproduce stationary and perturbed statistics of the system.
+In the case of partial observations, as in realistic settings, the performance of a data-driven model necessarily depends on (i) the choice of proper variables to model and (ii) suitable parameterizations of unobserved processes.
+The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [5, 50, 51], though typically without addressing responses to external perturbations.
+Therefore, we will argue that, in the case of low-frequency climate dynamics, a promising perspective is to build on such established approaches, emphasizing coarse-grained dynamics and stochastic parameterizations, while using neural networks as powerful tools within this framework.
 We emphasize that the focus on a simplified dynamical system is not a limitation in this context.
-In fact, the goal is not to demonstrate the success of a method, but to illustrate, by analogy, fundamental issues that are already evident in low dimensions and are likely to become more pronounced in higher-dimensional settings.
+In fact, the goal is not to demonstrate the success of a method, but to illustrate, by analogy, fundamental issues of data-driven modeling that are already evident in low dimensions and are likely to become more pronounced in higher-dimensional settings.
 As Smith (2002) [52] argues, while it is unreasonable to expect solutions to low-dimensional problems to generalize to high dimensions, so too it is unlikely that limitations observed in low-dimensional systems will vanish in high-dimensional, operational models.
+The first part of this paper is pedagogical in nature, aiming to bridge perspectives across researchers in statistical physics, climate science, and machine learning, while also providing practical methodological tools for the design and evaluation of data-driven models.
 Building on these ideas, we conclude this study by considering a real-world example.
-Throughout this paper, we emphasize, where possible, the distinction between “real-world examples/systems” and merely high-dimensional dynamics.
+Throughout this paper, we emphasize, when possible, the distinction between “real-world examples/systems” and merely high-dimensional dynamics.
 By real-world examples we refer to situations where (i) we observe only the temporal evolution of a subset of variables without access to the full state vector (i.e.
@@
 (1950) [54] and the discussion in Dalmedico (2001) [51].
-The example we consider simulates the joint evolution of the sea surface temperature (SST) field and the net radiative flux at the top of the atmosphere (TOA).
-Our motivation lies in understanding the spatial structure of climate feedbacks by addressing the so-called pattern effect, i.e.
-the dependence of Earth’s energy imbalance on the spatial patterns of surface warming [30].
-To this end, we develop a simplified stochastic neural model based on 20 SST modes and a global mean, scalar time series for the radiative flux.
-Unobserved degrees of freedom are parametrized through multiplicative (state-dependent) noise.
-We demonstrate that the model captures both the stationary statistics and the radiative fluxes response to imposed temperature perturbations in two climate change experiments.
-Such model can then be used to explore the linkages between surface temperature warming patterns and radiative fluxes, as well as a “reconstruction” tool, by predicting the change in one variable given prescribed changes in the other.
-Finally, we apply ideas from response theory to infer direct causal relationships between surface temperature and TOA radiative flux.
-These results show qualitative agreement with perturbation experiments using atmospheric-only climate models and align with current physical understanding of radiative feedbacks in the climate system.
+The example we consider is the “pattern effect” introduced above, which requires Green’s function experiments to quantify the response of top-of-atmosphere (TOA) radiative fluxes to perturbations in the sea surface temperature (SST) field.
+To this end, we develop a simplified stochastic neural model that emphasizes a large-scale representation of SST and global-mean net radiative flux, rather than attempting to resolve all spatial and temporal scales of variability as in current emulators.
+We discuss the coarse-graining and preprocessing choices relevant for model construction.
+The model largely reproduces both stationary and forced statistics and, importantly, enables Green’s function–like experiments to infer direct causal relations between SST and TOA radiative flux, a task where general-purpose emulators can perform poorly [37].
+Crucially, we focus on responses in the ensemble mean and variance, rather than single trajectories as in most current studies, thereby extending the concept of climate feedback to include changes in distributional properties.
 The paper is organized as follows: in Section II we introduce the theoretical ideas underpinning this work.
@@
 We begin by outlining the main theoretical concepts that underpin this work.
-First, we briefly introduce general ideas for constructing Langevin-like equations from partial observations of multiscale systems.
-We then review key aspects of response theory, which offers a general framework for understanding and assessing the ability of models to respond to perturbations.
-II.1 Reduced stochastic models: theoretical and practical considerations
-Why reduced order modeling and effective equations?
-Consider a nonlinear dynamical system $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$.
-The main interest in data-driven modeling, is to estimate the operator $\mathbf{f}$ directly from data.
-In doing so we face a serious limitation: in many high-dimensional complex systems we have access only to partial observations of the state vector $\mathbf{x}$.
-A temptation could be to use Taken’s embedding theorem [45].
-While theoretically appealing, this strategy has significant practical limitations: it is not valid for stochastic systems and it becomes infeasible as the dimensionality of the system increases [46, 32, 40, 33, 47, 48].
-A more promising direction leverages the multiscale nature of real-world dynamical systems by focusing on coarse-grained, effective dynamics.
-In some cases, it is possible to identify slow variables exhibiting low-dimensional effective dynamics and treat the cumulative effect of unresolved fast variables on the slow ones as stochastic forcing terms [49, 50, 51, 3, 4, 37, 36, 2, 52].
-In these cases, the system state vector $\mathbf{x}$ can be decomposed into slow and fast components, $\mathbf{x}=(\mathbf{x}_{s},\mathbf{x}_{f})$, and motivates the derivation of an effective Langevin equation for the slow variables of the form:
-$d\mathbf{x}_{s}=\mathbf{f}_{eff}(\mathbf{x}_{s})dt+\mathbf{\Sigma}(\mathbf{x}_{s})d\mathbf{W},$ (1)
-where $\mathbf{f}_{eff}$ represents the deterministic “drift” dynamics, $\mathbf{x}_{s}$ are the slow variables and $\mathbf{W}$ is the standard Wiener process.
-The noise term parametrize the cumulative effect of unresolved fast processes $\mathbf{x}_{f}$ on the slow ones.
-In the general case the noise is “multiplicative”, i.e.
-state dependent: $\mathbf{\Sigma}(\mathbf{x}_{s})$.
-However, in many practical applications, additive noise can be preferred, with $\mathbf{\Sigma}$ independent of $\mathbf{x}_{s}$.
-In what follows, we will simplify notation by referring to $\mathbf{x}$ instead of $\mathbf{x}_{s}$, and we will clarify the precise meaning as needed in the text.
-The reduced-order modeling approach emerges naturally in the study of large-scale climate dynamics, where recurrent spatiotemporal patterns, i.e.
-modes of variability, dominate the system’s behavior [53, 54].
-Unlike traditional general circulation models (GCMs) that aim to simulate as many spatiotemporal scales as possible, reduced models target specific processes/problems, capturing a few key degrees of freedom while treating the rest as stochastic forcing.
-This not only reduces computational cost but also enhances mechanistic understanding, by isolating core processes and studying their qualitative behavior (see e.g.
-[55], as well as the introduction of [44]).
-Consequently, there is a growing need for systematic methods to reduce the dimensionality of high-dimensional systems and to quantify causal interactions across their degrees of freedom.
-This has fueled the development of low-dimensional stochastic models learned either from data [51, 3, 4, 56, 10, 7, 2] or from known equations [49, 50, 57, 58], as well as inference-based techniques to infer interactions across modes of variability [59, 60, 61, 62].
-In this paper, we will primarily focus on data-driven stochastic modeling from partial observations.
-Nevertheless, we emphasize that even when all the relevant variables of the system are known, in multiscale physical systems there is often a need to formulate new effective (even deterministic) equations of the slow variables rather than focusing on the primitive equations.
-A classic example in the weather/climate literature is given by Charney (1948,1949) [63, 43] as well as Charney et al.
-(1950) [44].
-Charney was among the first to recognize the importance of filtering out what he termed “meteorologically insignificant noise” from the primitive equations, leading to more interpretable and computationally efficient effective equations.
-Data-driven reduced-order modeling: practical considerations and prior work.
+First, we briefly introduce general ideas for reduced-order stochastic modeling relevant for climate dynamics.
+We then review key aspects of response theory, which offers a rigorous framework to move beyond the evaluation of stationary dynamics, probe causal mechanisms, and ultimately guide model design.
+II.1 Reduced-order stochastic climate models: theoretical and practical considerations
+A general dynamical system can be represented in the form
+$d\mathbf{u}/dt=\mathbf{F}+\mathbf{L}\mathbf{u}+\mathbf{n}(\mathbf{u}),$ (1)
+where $\mathbf{F}$, $\mathbf{L}$ and $\mathbf{n}(\mathbf{u})$ represent an external forcing, a linear and nonlinear operators, respectively.
+$\mathbf{u}$ is not necessarily finite dimensional, and in the context of coupled atmosphere-ocean variability, it could in principle represent the deviation of all fields, at all scales, from their climatology [4].
+However, for obvious practical reasons, Eq.
+(1) needs to be simulated on finite spatial and temporal resolutions.
+In contrast to numerical weather prediction, which demands high spatial and temporal resolution, low-frequency climate variability is governed by a few dominant large-scale modes [4, 7, 48, 47, 55, 18].
+The goal of reduced-order stochastic climate models is to develop low-dimensional models simulating large-scale, coarse-grained climate dynamics.
+The cumulative effect of unresolved fast variables, present at smaller spatial scales, is parametrized as stochastic terms.
+More formally, we decompose $\mathbf{u}$ into slow and fast components, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, leading to an effective Langevin equation of the form:
+$d\mathbf{x}=\mathbf{f}_{eff}(\mathbf{x})dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W},$ (2)
+where $\mathbf{f}_{eff}$ represents the deterministic “drift” dynamics, $\mathbf{x}$ are the slow variables and $\mathbf{W}$ is the standard Wiener process.
+The noise covariance matrix $\mathbf{\Sigma}$ can be independent of $\mathbf{x}$, corresponding to additive noise, or state-dependent $\mathbf{\Sigma}(\mathbf{x})$ corresponding to multiplicative noise.
+In the latter case, the equation is here interpreted in the Itô sense.
+Equation (2) is mathematically justified in the limit of infinite time-scale separation between fast and slow modes [1].
+In the most general (and often intractable) setting, the effective dynamics of the slow variables also involve memory effects induced by the fast degrees of freedom, leading to non-Markovian dynamics.
+Kondrashov et al.
+(2015) [7] addressed this problem in a practical and efficient way by introducing the framework of multilevel stochastic models.
+Nonetheless, appropriate coarse-graining procedures have been shown to effectively average out non-Markovian contributions in many climate applications, yielding useful Markovian approximations [3, 4].
+We refer the reader to [2] for a review on the role of non-Markovian contributions in effective climate equations and to [56, 57] for specific examples.
+Reduced-order stochastic climate models, as in Eq.
+(2), provide an efficient alternative to state-of-the-art climate models for studying low-frequency climate variability.
+They are much more computationally efficient and allow us for mechanistic understanding of climate dynamics by isolating core processes and examining their qualitative behavior.
+This has fueled the development of low-dimensional stochastic models learned either from data [58, 3, 4, 16, 10, 7, 2, 59] or from known equations [60, 61, 62, 63, 46], as well as inference-based techniques to infer interactions across modes of variability [64, 65, 66, 18].
 In practice, given a phenomenon of interest, we typically begin with spatiotemporal data consisting of $N$ time series of length $T$. We then often proceed as follows:
 Coarse-grained dynamics.
-In spatiotemporal systems like climate, coarse-graining involves: (i) selecting a few relevant modes, (ii) isolating specific temporal scales, and (iii) choosing a limited set of variables.
-Regarding point (i), we reduce the dimensionality of the system in terms of some bases, a common choice is the one of Empirical Orthogonal Functions (EOFs) [39, 8].
-We retain only the first $n<N$ of such bases, and project the data onto them to obtain the principal components (PCs), which serve as the reduced time series for model training.
-Although EOFs are convenient, they may not be always optimal, and alternatives such as those in [64, 62, 65] (among many others) may be more suitable depending on the application.
+In spatiotemporal dynamical systems like climate, coarse-graining involves: (i) choosing a limited set of climate fields (e.g.
+temperature, velocities etc.), (ii) retaining a few relevant modes or large-scale averages, (iii) restricting the range of temporal scales.
+Throughout the paper, we will often refer to the so-called proper variables as the coarse-grained spatiotemporal fields obtained via (i–iii).
+The relevant modes are often identified through Empirical Orthogonal Functions (EOFs) [67, 8], by retaining the first $n<N$ of such bases.
+The correspondent principal components (PCs), serve as the reduced time series for model training.
+Although EOFs are convenient, they may not be always optimal, and alternatives such as in [68, 18, 69] may be more suitable depending on the application.
 The performance of the final model is sensitive to all coarse-graining choices (i–iii); see, e.g., [70, 71].
-Considering the “right” coarse-grained procedures is perhaps the most difficult step in modeling from data.
-There is in fact no general mathematical guideline in high-dimensional dynamical systems for choosing the proper variables to model, an old problem [68] emphasized in many new contributions, such as [31, 69, 47, 48].
+Identifying the “right” coarse-grained representation remains one of the hardest challenges in data-driven modeling, as no general mathematical principle exists to guide such choices in high-dimensional dynamical systems [43, 17, 17].
 Inferring a stochastic model.
-Given a reduced order representation of the system, we are now left with the task of inferring a stochastic model of the form of Eq.
+Given the coarse-grained data $\mathbf{x}$, the goal is to infer a stochastic Markovian model as in Eq.
 (2).
-Many studies have shown that, for specific observables (e.g., sea surface temperature anomalies), a linear drift with additive noise provides forecasting skill comparable to full complexity models [4, 5, 70, 71].
-This leads to a Linear Inverse Model (LIM) [3] of the form
-    $d\mathbf{x}=\mathbf{L}\mathbf{x}dt+\mathbf{\Sigma}d\mathbf{W},$
-    where $\mathbf{L}$ is a linear operator, $\mathbf{\Sigma}$ is a noise matrix and $\mathbf{W}$ is the standard Wiener process.
-An interesting recent extension to tackle colored noise can be found in [72].
-The multilevel regression proposed by Kratsov et al.
-(2005) [6, 7] generalizes these ideas by incorporating quadratic nonlinearities and non-Markovian effects with a series of “matryoshka”-like levels for residuals.
-In many practical cases, however, temporal coarse-graining (e.g., averaging over a few time steps) can effectively remove memory effects, allowing a Markovian closure [73].
-In the framework proposed in [6, 7], the drift consists of both linear and quadratic nonlinear components.
-The inclusion of quadratic nonlinearities is physically motivated rather than ad-hoc.
-As shown in Section 2 of Lorenz (1963) [74], projecting forced-dissipative flows (in periodic domains) onto Fourier modes, yield a set of infinite coupled ODEs with linear and quadratic interactions for the modes’ amplitudes.
-In particular, the quadratic nonlinearity $\mathbf{B}(\mathbf{x},\mathbf{x})$ should be energy-conserving, satisfying $\mathbf{x}\cdot\mathbf{B}(\mathbf{x},\mathbf{x})=0$.
-This abstract representation underlies the starting point of the Majda-Timofeyev-Vanden Eijnden (MTV) mode reduction framework [49, 50, 57, 58].
-Additionally, Majda and Harlim focused on the energy conservation constraint above to propose physically constrained regression models [11].
-Guided by these insights, the neural models proposed in this work will have a deterministic dynamics specified by a linear term $\mathbf{L}$ and general nonlinear terms $\mathbf{n}(\mathbf{x})$; in formulae:
-    $\frac{d\mathbf{x}}{dt}=\mathbf{L}\mathbf{x}+\mathbf{n}(\mathbf{x}).$
-    It is in fact a sensible choice to guide the learning by specifying a decomposition into linear and nonlinear terms, instead of letting the neural network “figure out” the functional form by itself (see also [75]).
-Furthermore, we will move beyond additive noise and incorporate multiplicative (state-dependent) noise, both in one and higher dimensions.
-As we will show, incorporating state-dependent noise can be at times critical for accurately capturing the system’s response to external perturbations in the presence of partial observations.
+We summarize several traditional approaches that motivate our work, while acknowledging that many other methods exist beyond those discussed here.
+The simplest example is the Linear Inverse Model (LIM) proposed by Penland [3], where the drift is linear and the noise is additive, both estimated directly from data.
+LIMs have shown forecasting skill for sea surface temperature anomalies comparable to climate models [4, 5, 72, 73], and extensions exist to handle colored noise [74].
+Kravtsov et al.
+(2005) and Kondrashov et al.
+(2015) [6, 7] generalized this framework through multilevel regression, incorporating quadratic drift terms and non-Markovian effects.
+These models perform well for stationary statistics [8] and forecasting [75].
+The work of Majda and Harlim extended the multilevel regression framework by proposing stability constraints [11].
+A theoretically grounded alternative is the approach of Majda, Timofeyev, and Vanden-Eijnden [60, 61, 62, 63], dubbed MTV by the authors, which derives effective equations for the coarse-grained variables $\mathbf{x}$ rather than assuming them.
+MTV has been successful for simplified models, though less so for complex climate systems, likely due to choices of coarse-graining steps (i-iii) shown above rather than intrinsic limitations, see also discussion in [8].
+We will expand on this approach in Section III and Appendix B.
+Overall, these works demonstrated that robust reduced-order stochastic climate models can be built through relatively simple regression strategies.
+In this paper, we build on these insights and investigate how simple neural architectures can infer both the drift and state-dependent noise in Eq.
+(2), with minimal assumptions.
+Our neural models retain a decomposition into linear and nonlinear terms, $\mathbf{L}$ and $\mathbf{n}(\mathbf{x})$, as in the general representation of dynamical systems in Eq.
+(1).
+Furthermore, we will propose simple solutions to fit multiplicative (state-dependent) noise in high dimensions.
 II.2 Response theory: a (very) brief overview
-In this paper, we argue that a reduced-order model should capture not only the system’s stationary statistics but also its response to external perturbations.
-This naturally raises the question: How to assess skillful responses to general perturbations?
-Since it is impossible to test responses to the full space of possible forcings, linear response theory provides a principled framework for analyzing the system’s linear response to perturbations.
+In this paper, we argue that data-driven models should be tested on their ability to capture both stationary statistics and responses to external perturbations.
+Assessing responses to arbitrary forcings is ill-posed, as the space of possible perturbations is essentially infinite.
+A way forward is to consider small impulse perturbations, which allow us to define a response operator and derive general conclusions about causality and forced responses within the linear response regime.
 Impulse response operator.
@@
 $\delta\langle A(x_{k}(t))\rangle=\langle A(x_{k}(t))\rangle_{\text{p}}-\langle A(x_{k}(t))\rangle~,$ (3)
-[…]
+where $\langle\cdot\rangle$ refers to ensemble average and the subscript “p” refers to the perturbed system.
+In this paper we are going to focus on two different specific observables: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution.
+The two observables allow us to capture time-dependent responses in the mean and variance of the perturbed distribution respectively.
+The impulse response operator $\mathbf{R}(t)\in\mathbb{R}^{n,n}$ is then defined as:
 $R_{k,j}(t)=\lim_{\delta x_{j}(0)\to 0}\frac{\delta\langle A(x_{k}(t))\rangle}{\delta x_{j}(0)}.$ (4)
@@
 Understanding causal relations across degrees of freedoms in physical systems require us to study the effect of interventions on the system [77].
-The inferred response operator quantifies time-dependent cause-effect dependencies across degrees of freedoms $x_{j}$ and $x_{k}$; under the assumption that the reduced order models correctly capture the physics of the system.
-For more details to this physics-based approach to causal inference see [78, 79, 47, 80, 62].
+The inferred response operator quantifies time-dependent cause-effect dependencies across degrees of freedoms $x_{j}$ and $x_{k}$ [78, 79, 17, 80, 18].
+We refer to Appendix A for details on the relation between the response operator and causality.
   • Response to generic time-dependent forcing.
 The response operator $\mathbf{R}(t)$ can be used to predict the response of the system to small, but general, time-dependent external forcing $\mathbf{f}(t)\in\mathbb{R}^{n}$ applied directly to the system’s degrees of freedoms as:
-[…]
+    $\delta\langle A(x_{k}(t))\rangle=\sum_{j}\int_{0}^{t}R_{k,j}(\tau)f_{j}(t-\tau)d\tau.$ (5)
     Importantly, while we assume linearity in the system’s response, the formulas above are valid for both linear and nonlinear systems.
 A data-driven model capable of accurately capturing the impulse response operator will also capture the time-dependent causal interactions across degrees of freedom.
-Response theory can then be used to assess the skills of emulators by going beyond purely statistical metrics such as r-squared, by for example assessing the causal sensitivity of the system to external perturbation, see e.g.
-discussion in Falasca et al.
-(2024) [62].
-Since the linear response to small perturbations is given by the convolution formula in Eq.
+Response theory can then be used to assess the skills of emulators by going beyond purely statistical metrics such as r-squared, by for example assessing the causal sensitivity of the system to external perturbation [18].
+Moreover, since the linear response to small perturbations is given by the convolution formula in Eq.
 (5), recovering the impulse response operator ensures a faithful approximation of the system’s linear response to general forcings.
-For this reason, we view the impulse response operator as a fundamental building block that a data-driven model must reproduce to reliably capture at least the linear component of responses to general perturbations, and is necessary (but not sufficient) to capture nonlinear responses as well.
-Inferring the response operator in practice.
-Given a model $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W}$ we build the impulse response operator as follows:
-  1. 1.
-    We simulate a very long trajectory of the model, starting from a random initial condition and remove an initial transient.
-We then sample $N_{e}$ random points from the simulation to define an ensemble of $N_{e}$ initial conditions on the model’s attractor.
-$N_{e}$ should be very large (i.e.
-$N_{e}\gg 1$) in order to sample the whole attractor.
-  2. 2.
-    For each one of the $N_{e}$ initial conditions, we impose an impulse perturbation $x_{j}(0)\rightarrow x_{j}(0)+\Delta_{j}\delta(t)$ of size $\Delta_{j}$ on the $j$-th degree of freedom at time $t=0$.
-$\delta(t)$ is the Dirac delta function.
-The amplitude $\Delta_{j}$ should be theoretically infinitesimally small.
-In practice we do as follows: we consider the long time series of $x_{j}(t)$ from the long control integration above and define $\Delta_{j}=10^{-2}\sigma_{j}$, where $\sigma_{j}$ is the standard deviation of time series $x_{j}(t)$.
-  3. 3.
-    Therefore, for a given initial condition $\mathbf{x}(0)$, we simulate two trajectories: a control trajectory without perturbation and a perturbed one, where an impulse perturbation $\Delta_{j}$ has been imposed on the $j$-th degree of freedom at time $t=0$.
-This procedure is repeated in parallel for all initial conditions, resulting in an ensemble of $N_{e}$ paired of control and perturbed trajectories.
-  4. 4.
-    At each time $t$ we then estimate the time dependent ensemble average $\langle A(x_{k}(t))\rangle$ of an observable $A(x_{k}(t))$ for both the perturbed and unperturbed run.
-We refer to the perturbed and control ensemble average respectively as $\langle A(x_{k}(t))\rangle_{\text{p}}$ and $\langle A(x_{k}(t))\rangle$.
-The observables considered in this study are going to be: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the perturbed distribution, therefore quantifying the response in ensemble mean and variance.
-  5. 5.
-    We define the impulse response operator for observable $A(x_{k}(t))$ as
-    $R_{k,j}(t)=\frac{\delta A(x_{k}(t))}{\Delta_{j}}=\frac{\langle A(x_{k}(t))\rangle_{\text{p}}-\langle A(x_{k}(t))\rangle}{\Delta_{j}}.$
+For this reason, we view the impulse response operator as a fundamental building block that a data-driven model must reproduce to capture the response to small, general perturbations, and is necessary (but not sufficient) to capture nonlinear responses as well.
+Importantly, the response operator can be inferred from data alone through the Fluctuation-Dissipation Theorem [49], this is discussed in Appendix A.
 III Limitations of data-driven modeling by analogy: the case of partial observations
-As argued by Cecconi et al.
-(2012) [31], the main challenge in modeling real-world dynamical systems from data stems from their high dimensionality.
-Learning a dynamical system from data requires sufficiently long time series to sample many recurrences.
-However, the recurrence time, i.e.
-the time needed for a trajectory to return to a previous region in state space, scales exponentially with the system’s dimensionality [31, 81].
-The success of data-driven methods in real-world systems is therefore often attributed to the presence of an effective low-dimensional dynamics.
-In other words, the dynamics of the system usually lives on a low-dimensional manifold embedded in a higher-dimensional space.
-This so-called “manifold hypothesis” underpins research in fields ranging from spatiotemporal chaos [82, 83, 84, 85] to neuroscience [86, 87, 88, 89, 90], biophysics [91, 92], artificial neural networks [93, 94] and climate dynamics [74, 1, 95, 96, 54].
-In practice, however, even when a system exhibits low effective dimensionality, we still face several difficult challenges.
-First, it is often unclear how to construct an optimal basis for efficient modeling in the system’s intrinsic coordinates [64].
-Second, and more fundamentally, we typically observe only a subset of relevant variables of a complex, multiscale physical system.
-Here we illustrate how learning on only partial observations, as is common in real-world applications [10, 33], can hinder data-driven modeling, even when the underlying dynamics is low-dimensional and seemingly tractable.
-Importantly, while the discussion is framed in terms of neural emulators, motivated by current challenges in climate science, the issues we highlight are general and inherent to data-driven approaches for real-world dynamical systems.
-III.1 A triad model as a prototype stochastic climate model
-We consider a triad model originally introduced by Majda and collaborators in the context of low-frequency climate dynamics [34, 36]. The model is:
-$\begin{split}dx_{1}&=(L_{11}x_{1}+L_{12}x_{2}+L_{13}x_{3}+Ix_{1}x_{2})dt\\ dx_{2}&=(-L_{12}x_{1}-Ix_{1}^{2}-\gamma_{2}\epsilon^{-1}x_{2})dt+\sigma_{2}\epsilon^{-1/2}dW_{2}\\ dx_{3}&=(-L_{13}x_{1}-\gamma_{3}\epsilon^{-1}x_{3})dt+\sigma_{3}\epsilon^{-1/2}dW_{3}\\ \end{split}$ (5)
-This idealized system mimics the functional structure of a reduced stochastic climate model, featuring linear couplings $L_{11},L_{12},L_{13}$ and energy-conserving quadratic nonlinearities.
-The parameter $I$ controls the strength of the nonlinearity: when $I=0$, the model reduces to purely linear dynamics.
+A fundamental limitation of purely data-driven modeling of multiscale climate systems is that we cannot observe all degrees of freedom.
+Even when governing equations are known, as in meteorology, climate fields are only available at finite temporal and spatial resolution.
+Dynamics at smaller and faster scales are therefore missing from the data and often remain unparameterized in emulators, unlike in full climate models.
+An even greater challenge arises in many areas where we lack governing equations altogether to guide emulator construction.
+In both cases, data-driven approaches can mimic the observed variability of the retained variables, reproducing probability distributions, autocorrelation functions, and providing reasonable forecasts.
+Yet testing responses to perturbations, where unobserved dynamics becomes relevant, often exposes structural deficiencies that are otherwise hidden.
+The goal of this section is to illustrate such limitations within a controlled mathematical setting, where they can be made explicit and unambiguous.
+III.1 A triad model as a simplified prototype for coarse-grained, stochastic climate modeling
+We focus on a simplified yet relevant dynamical system that retains essential features of high-dimensional geophysical flows, making the conclusions of the Section meaningful for climate applications.
+The model was introduced by Majda and collaborators [46] and later used in [47] in the context of climate dynamics.
+In what follows, we present the model and refer to Appendix B for an in-depth discussion on its relevance as a building block for complex turbulent flows.
+The triad model considered in this study takes the following form:
+[…]
+This idealized system mimics the functional structure of turbulent geophysical flows, featuring linear couplings $L_{11},L_{12},L_{13}$ and energy-conserving quadratic nonlinearities.
+The parameter $I$ controls the strength of the nonlinearity: when $I=0$, the model reduces to purely linear dynamics and Gaussian statistics.
+The quadratic interactions, here between $x_{1}$ and $x_{2}$, typically arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs [16].
 The parameter $\epsilon$ sets the time-scale separation between the slow mode $x_{1}$ and the fast modes $(x_{2},x_{3})$, allowing exploration of multiscale dynamics in a controlled setting.
 The terms $W_{2}$ and $W_{3}$ represent independent Wiener processes, introducing stochastic forcing into the second and third equations.
-This stochastic forcing, together with the linear damping terms $\gamma_{2}$ and $\gamma_{3}$, serves as a linear proxy for the nonlinear self-interactions of the fast variables (see Eq.
-(4) in [49] or Section 3.3 in [37] for further details).
-Importantly, the model generates relatively small, but significant, deviations from Gaussianity which is a relevant feature in the study of low-frequency climate dynamics [34] and found in both general circulation models and observational data, see e.g.
+The model generates unimodal distributions with relatively small, but significant, deviations from Gaussianity, which is a relevant feature in the study of low-frequency climate variability [81] and found in both general circulation models and observational data, see e.g.
 [82, 83, 84].
-III.2 Data-driven modeling when we observe the full state vector
+III.2 The case of unknown equations but a fully observed state vector
 We consider the relatively simple case of a low-dimensional dynamical system with only one characteristic time scale (i.e.
@@
 $10^{5}$ Model Time Units).
-As we will show later, given the system’s decorrelation time scale, the resulting time series can be effectively treated as infinitely long.
-With this long trajectory of the full state vector $\mathbf{x}(t)=[x_{1},x_{2},x_{3}](t)$ in hand, we now attempt to learn data-driven models directly from the time series.
+Given the system’s decorrelation time scale (shown later), the resulting time series can be effectively treated as infinitely long.
+Given this long trajectory of the full state vector $\mathbf{x}(t)=[x_{1},x_{2},x_{3}](t)$, we learn data-driven models directly from the time series.
 We construct both linear and nonlinear stochastic models.
-We specify the functional forms of these models and assess their ability to reproduce both the stationary and perturbed statistics of the original system defined in Equation (5).
+We specify the functional forms of these models and assess their ability to reproduce both the stationary and perturbed statistics of the original triad system.
 III.2.1 Functional form of the data-driven models considered
-We consider two data-driven models and report their functional forms below.
-To keep the main text focused, we leave the technical details of the fitting procedures, along with details of the neural networks, in Appendix A.
+We consider two data-driven models. To keep the main text focused, we leave the technical details of the fitting procedures, along with details of the neural networks, in Appendix C.
   • Linear model. We consider the following linear model:
@@
 Crucially, the term $(\mathbf{L}+\mathbf{n}(\mathbf{x}))$ is fitted jointly.
-As for the linear model, the $\mathbf{\Sigma}\mathbf{\Sigma}^{\text{T}}$ is the noise covariance matrix and $\mathbf{W}$ is the standard Wiener process.
 In what follows we are going to often refer to this model as “neural model” or “neural emulator”.
@@
 The linear method is considered for two reasons.
-First, we show that linear inverse modeling (LIM), as originally proposed in [3], serves as a strong baseline for neural autoregressive models, successfully capturing both the mean and variance of the system’s stationary distribution, the autocorrelation functions, and the ensemble-mean response to external perturbations.
+First, we show that linear inverse modeling (LIM), as originally proposed in [3], serves as a strong baseline for neural autoregressive models, successfully capturing both the mean and variance of the system’s stationary distribution, the autocorrelation functions, and the ensemble mean response to external perturbations.
 Second, and consistent with theoretical expectations [47], we demonstrate that linear models are fundamentally limited in capturing changes in the full probability distribution: for instance, they cannot represent any response in the variance to external forcings.
-This limitation restricts their usefulness for studying forced changes in distributions beyond the mean, which, dependent on the research question, can be a central objective in climate science.
+This limitation restricts their usefulness for studying forced changes in distributions beyond the mean.
 III.2.2 Stationary statistics
@@
 We then compare the modeled and original probability density functions (PDFs) and autocorrelation functions (ACFs), as shown in Figure 1.
-The neural model described in Equation (7) successfully captures the stationary distributions of all three modes $x_{1}(t),x_{2}(t)$ and $x_{3}(t)$.
-In particular, it outperforms the linear model by accurately reproducing the non-Gaussian features of the first mode $x_{1}$.
+The neural model described in Equation (8) successfully captures the stationary distributions of $x_{1}(t),x_{2}(t)$ and $x_{3}(t)$.
+In particular, it outperforms the linear model by accurately reproducing the non-Gaussian features of $x_{1}$.
 Both models also reproduce the autocorrelation functions well, as seen in the second row of Figure 1.
-In summary, the neural emulator defined in Equation (7) is capable of capturing both the stationary statistics and temporal correlations of the full state vector $\mathbf{x}(t)$.
+In summary, the neural emulator defined in Equation (8) is capable of capturing both the PDF and autocorrelations of the full state vector $\mathbf{x}(t)$.
 At the same time, the linear model serves as a strong baseline for evaluating more sophisticated data-driven approaches.
@@
 III.2.3 Response of probability distributions to external perturbations
-Impulse response functions
+Impulse response functions.
 We now examine the impulse response of the original triad model (Equation (6)), the linear model (Equation (7)), and the nonlinear neural model (Equation (8)).
 Specifically, we analyze the time-dependent response of observables $A(x_{k}(t))$, with $k=1,2,3$, to an impulse perturbation imposed on $x_{1}(0)$ at time $t=0$.
-The observables considered here are the mean and variance of the time-dependent perturbed distribution.
-The ensemble size considered here is $N_{e}=10^{5}$ trajectories.
+We choose as observables the mean and variance of the time-dependent perturbed distribution and refer to Appendix A for details on how to compute the response operator.
 The results, shown in Figure 2, reveal that the nonlinear neural model captures the mean response of the distribution quantitatively, while the linear model captures it only qualitatively.
-The most significant differences arise in the response of the variance: the linear model yields exactly zero variance response, clearly highlighting the limitations of linear dynamics.
-This is expected, as a linear model with a Gaussian invariant measure produces zero response to any quadratic functionals, such as the variance, consistent with the theoretical results in [36].
+The most significant differences arise in the response of the variance: the linear model yields exactly zero variance response, clearly highlighting the limitations of linear dynamics [47].
 In contrast, the neural model captures the variance response with relatively high fidelity.
-As argued in Section II.2, the impulse response operator plays a foundational role in determining a model’s ability to capture the linear part of the response to general perturbations.
-We therefore consider the neural model presented here as a skillful model both in terms of reproducing the stationary distribution as well as the response to forcing.
+We therefore consider the neural model presented here as a skillful model both in terms of reproducing the stationary distribution as well as the response operator.
 Figure 2: Impulse response functions as modeled by the original triad system in Eq.
@@
 Responses to a specific logarithmic forcing.
-Having evaluated the ability of the data-driven neural model to capture the impulse response operator, we now consider a more targeted test involving a specific, time-dependent forcing.
-We introduce a logarithmic forcing term to the right-hand side of the original triad model (Equation (5)), the linear model (Equation (6)), and the neural model (Equation (7)).
-The forcing is applied only to the first mode $x_{1}$ and takes the form $(0.01log(1+t),0,0)$ per time step.
-Each model is integrated for 2500 time steps (25 Model Time Units), and we compute the difference between the forced and unforced simulations in terms of ensemble-mean and ensemble-variance responses.
+we now consider a more targeted test involving a specific, time-dependent forcing.
+We introduce a logarithmic forcing term to the right-hand side of the original and data-driven triad model.
+The forcing is applied only to $x_{1}$ and takes the form $(0.01log(1+t),0,0)$ per time step.
+Each model is integrated for 2500 time steps (25 Model Time Units), and we compute the difference between the forced and unforced simulations in terms of ensemble mean and variance responses.
 Results are presented in Figure 3.
-The neural model accurately reproduces the forced responses across all three components.
-The linear model performs qualitatively well but exhibits significant deviations in the ensemble-mean response of the second mode $x_{2}$ as shown in Figure 3(b).
+The neural model (Eq.
+(8)) accurately reproduces the forced responses across all three components.
+The linear model (Eq.
+(7)) performs qualitatively well but exhibits significant deviations in the ensemble mean response of the second mode $x_{2}$ as shown in Figure 3(b).
 This is consistent with the earlier Section: the largest error in the linear model’s impulse response also occurred in $x_{2}$ (see Figure 2(b)).
 This supports the interpretation that accurately capturing the impulse response operator is critical for a model’s reliability in predicting responses to general, time-dependent forcings.
-As expected from the analysis of the response operator, the linear model is not able to capture the response in ensemble variance.
-The neural model shows skillful response, capturing the time-dependent changes in ensemble variance relatively well, with the largest error in the second mode $x_{2}$.
+As expected, the linear model is not able to capture the response in ensemble variance.
+The neural model shows skillful response, capturing the time-dependent changes in ensemble variance relatively well, with the largest error in $x_{2}$.
 In general, considering that the changes in variance are very small, we conclude that the neural emulator is skillful in responding to the imposed logarithmic forcing.
@@
 III.2.4 Main takeaways
-Given a (practically) infinitely long trajectory of a stochastic dynamical system, and assuming the ideal condition where the full state vector is observed, it is possible to construct a neural stochastic model that is stable, reproduces the stationary statistics, and responds accurately to small but general external perturbations and forcings.
+Given a (practically) infinitely long trajectory of a stochastic dynamical system, and assuming the ideal condition where the full state vector is observed, it is possible to construct a neural stochastic model that is stable, reproduces the stationary statistics, and responds accurately to small but general external perturbations.
+The analysis does not rule out that other systems may struggle to capture responses even with full observations; however, these issues are here regarded as technical rather than fundamental.
+Such technical issues can be addressed by considering more complex architectures or by adding constraints withing a chosen architecture.
 While these results are encouraging, they remain far from any real-world scenario.
-In practice, even when a low-dimensional effective dynamics exists, the main challenge is that we typically have access only to partial observations of a complex, interconnected system.
-Such challenges lead to the important problem of having to choose the proper variables to model as well as the proposal of stochastic parametrizations to account for the unobserved degrees of freedom.
-III.3 Data-driven modeling when we do not observe the full state vector
+In practice, we typically have access only to partial observations of a complex, interconnected system.
+III.3 The case of unknown equations and a partially observed state vector
 We now return to the triad model defined in Eq.
@@
   1. i.
-    It must reproduce the statistics of the observed component;
+    It must reproduce the stationary statistics of the observed component;
   2. ii.
-    It must correctly captures its response to perturbations as in the original triad system.
-Criterion (ii) is particularly demanding because, in the triad system defined by Eq.
-(5), perturbations can propagate through the network of interacting variables, inducing cascades of feedbacks.
-These time-dependent feedback loops must be effectively parametrized within the scalar model, despite the absence of direct observations of some components.
+    It must correctly capture its response to perturbations as in the original triad system.
+Criterion (ii) is particularly demanding because, in the triad model defined by Eq.
+(6), perturbations can propagate through the system, inducing cascades of feedbacks.
+These time-dependent feedbacks must be effectively parametrized within the scalar model, despite the absence of direct observations of some components.
 While clearly nontrivial, this problem is representative of real-world data-driven modeling scenarios, where models are often built from partial observations of high-dimensional systems.
@@
   • Choice of variable to model.
-The derived reduced order model is now dependent on the variable we choose to model.
-Not all components can be modeled to meet the above criteria.
-In the simple case considered here the choice of $x_{2}$ or $x_{3}$ will be wrong as these two modes represent the fast dynamics forced by the much slower mode $x_{1}$.
-Therefore, it would be impossible to parametrize the effect of $x_{1}$ from information of $x_{2}$ or $x_{3}$ alone.
-The only possible way forward is to choose the slow mode $x_{1}$.
-Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model in the presence of only partial observations [37, 31, 52, 69].
+The reduced-order model depends critically on the choice of the modeled variable.
+Not all components can be selected to satisfy criteria (i) and (ii) discussed above.
+The only viable option is to model the slow mode $x_{1}$, since the cumulative effect of the much faster $x_{2}$ and $x_{3}$ can be represented stochastically.
+Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model under partial observations [48, 43, 85, 86].
+In realistic spatiotemporal climate dynamics, this choice translates into selecting appropriate climate fields together with a narrow range of spatial and temporal scales as we will show in Section IV.
   • The need for suitable parametrizations.
@@
 As we will show, the success of the resulting model, particularly in capturing forced responses, hinges critically on the quality of this parametrization.
+We note that for the specific case of the triad model in Eq.
+(6), Majda and collaborators showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ [46, 47]; see discussion in Appendix B.
+Our aim here is not to test such analytical strategies, which require full knowledge of the governing equations and are often limited in real-world settings; see footnote  [^1] Even moderately more complex systems already expose the practical limitations of rigorous stochastic mode reduction procedures; see Strounine et al.
+(2008) in the context of quasi-geostrophic modeling [8]..
+Rather, we use the triad model as a simple conceptual tool to highlight key challenges of data-driven modeling of multiscale systems.
+In realistic settings, both the governing equations and the full set of relevant variables are typically unknown [91].
+Even before specifying a model structure, identifying the slow variables is nontrivial and highly problem-dependent.
+In what follows, we assume that the proper slow variable, $x_{1}$, has been correctly identified.
+Given this, we examine how the choice of stochastic parametrization and temporal coarse-graining influences the ability of neural models to emulate both stationary statistics and responses to perturbations.
+This discussion serves as a precursor to the more challenging real-world application in Section IV, where identifying the relevant variables requires a series of nontrivial coarse-graining decisions.
 III.3.1 Functional form of the reduced order models considered
-Given the correct choice of variable $x_{1}$ to model we can now proceed in proposing the two data-driven scalar models considered.
-As in the previous Section, to keep the main text focused and readable, we leave the technical details of the fitting procedures for Appendix A.
-We consider two nonlinear, neural models differing with the stochastic parametrization considered:
-  • Neural model with additive noise.
-We consider the following neural model fitted using an artificial neural network, i.e.
-a Multilayer Perceptron (MLP) with one hidden layer.
-The model takes the following form:
+Given the correct choice of variable $x_{1}$ to model we now proceed in proposing two data-driven scalar models.
+Technical details of the fitting procedures are left to Appendix C.
+We consider two neural models:
+  • Neural model with additive noise. We consider the following neural model fitted using an artificial neural network, i.e. a Multilayer Perceptron (MLP). The model takes the following form:
     $dx=(Lx+n(x))dt+\sigma dW,$ (9)
-    $L$ is a linear operator (here just a scalar), $n(x)$ is a nonlinear operator and $\sigma^{2}$ is the noise variance.
-$W$ is the standard Wiener process.
-Note: the whole drift $(Lx+n(x))$ is fitted by the neural network.
+    $L$ is a linear operator (here just a scalar), $n(x)$ is a nonlinear operator and $\sigma^{2}$ is the noise variance. $W$ is the standard Wiener process.
   • Neural model with multiplicative noise. We consider a neural model as before, but now defined with multiplicative rather than additive noise:
     $dx=(Lx+n(x))dt+\sigma(x)dW.$ (10)
-    Differently from the previous model in Eq.
-(8), now the noise is also state dependent.
-The multiplicative noise $\sigma(x)$ is fitted with an additional neural network, see Appendix A.
-As before, we are going to often refer to the two models above as “neural model” or “neural emulator”.
+    The multiplicative (state dependent) noise $\sigma(x)$ is fitted with an additional neural network, see Appendix C.
 To train the two models above we consider a long simulation of the triad system in Eq.
 (6) using the set of parameters reported in Section III.3.
-We then extract only the time series of $x_{1}$.
-The two models in Eq.
-(8) and (9) are then trained on this long, scalar time series of $x_{1}(t)$ alone.
+We then extract only the time series of $x_{1}$ and train the models above on this long, scalar time series $x_{1}(t)$.
 III.3.2 Stationary and perturbed statistics of the scalar model
-We integrate the two data-driven scalar models and specified in Eq.
-(8) and (9) for $T=10^{7}$ time steps, yielding time series of the same length as the time series of $x_{1}(t)$ obtained from the long integration of the triad system used for training.
-To evaluate the models’ ability to capture both stationary and perturbed statistics, we examine the stationary probability distribution and the impulse response functions.
-Specifically, we focus on responses in the ensemble mean and variance by studying the time dependent change of the perturbed probability distribution $P_{p}(x,t)$ (where the subscript “$p$” stands for “perturbed”) to impulse perturbations in the variable $x$ itself, see Section II.2 for more details.
+We integrate the two data-driven scalar models in Eq.
+(9) and (10) for $T=10^{7}$ time steps, yielding time series of the same length as the time series of $x_{1}(t)$ used for training.
+To evaluate the models’ ability to capture both stationary and perturbed statistics, we examine the stationary probability distribution and the impulse response operator in ensemble mean and variance.
 The neural model with additive noise (Figure 4(a)) fails to accurately capture the stationary statistics in a quantitative way.
-Moreover, the ensemble mean response to an impulse perturbation decays too slowly, and notably, the model fails to reproduce any significant response in the ensemble variance.
+Moreover, the ensemble mean response to an impulse perturbation decays too slowly, and notably, the model fails to reproduce significant responses in the ensemble variance.
 In contrast, incorporating multiplicative noise (Eq.
-(9)) leads to a substantially improved representation of the stationary distribution, as seen in Figure 4(e), which closely matches the true distribution.
+(10)) leads to a substantially improved representation of the stationary distribution, see Figure 4(e).
 The mean response remains largely unchanged, decaying too slowly, but the model now captures a qualitatively correct signal in the ensemble variance, albeit with a similarly slow decay.
 These results indicate that the inclusion of multiplicative noise improves the reduced-order model’s ability to represent both stationary statistics and variance responses to perturbations.
-In the full triad system, a perturbation to $x_{1}$ at time $t=0$ leads to changes in $x_{1}$, $x_{2}$ and $x_{3}$ at the next time step $dt$.
+In the full triad system, perturbing $x_{1}$ at time $t=0$ leads to changes in $x_{1}$, $x_{2}$ and $x_{3}$ at the next time step $dt$.
 Consequently, $x_{2}$ and $x_{3}$ feedback on $x_{1}$ at the consequent time step.
-This leads to a cascade of state-dependent interactions until the system relaxes back to stationarity.
-Such complex, state-dependent feedbacks cannot be represented by additive noise alone and require the inclusion of state dependent noise $\sigma(x)$, effectively parametrizing such interactions.
+Such time-dependent feedbacks require the inclusion of state dependent noise $\sigma(x)$ to effectively parametrize such interactions.
+Moreover, as discussed in Appendix B, multiplicative noise is not an ad-hoc addition but is theoretically justified and explicitly predicted by rigorous mode reduction strategies [60].
 However, even when multiplicative noise is considered, the responses to impulse perturbations are still only qualitatively right.
@@
 (10).
-Impulse response functions are computed using an ensemble of $N_{e}=10^{5}$ members.
+Impulse response functions are computed using an ensemble of $N_{e}=10^{6}$ members.
 The presence of slowly decaying responses in the ensemble mean and variance to perturbations is indicative of a system that is no longer Markovian.
@@
 One strategy for addressing such memory effects is the multilevel regression approach first introduced by Kratsov, Kondrashov and Ghil (2005) [6] and later generalized by Kondrashov, Checkroun and Ghil (2015) [7].
-The multilevel regression strategy method augments the reduced-order model with a hierarchy of linear equations solving for the residuals from the deterministic dynamics, and can be thought as a practical implementation of the Mori–Zwanzig formalism in statistical mechanics [7, 2].
-Ultimately, the remaining residuals become white in time, allowing the system to be closed using a Wiener process.
-An alternative, practical approach for handling non-Markovian effects is to coarse-grain the data in the temporal dimension, for instance, by averaging over a time window.
+The multilevel regression strategy augments the reduced-order model with a hierarchy of linear equations solving for the residuals from the deterministic dynamics and it can be thought as a practical implementation of the Mori–Zwanzig formalism in statistical mechanics [7, 2].
+An alternative, pragmatic approach for handling non-Markovian effects is to coarse-grain the data in the temporal dimension, for instance, by averaging over a time window.
 This suppresses fast fluctuations within that window leading to an effective “Markovianization” of the dynamics.
@@
 [4]) and is also adopted in the present study.
-We therefore consider again the time series $x_{1}(t)$ from the long simulation of the triad system in Eq.
+We therefore analyze the time series $x_{1}(t)$ from the long simulation of the triad system in Eq.
 (6).
-As an additional step, we average the time series over a short temporal window of 10 time steps (recall that $dt=0.01$, so this is equivalent to $0.1$ Model Time Units).
-Importantly, this is not equivalent to sub-sampling the original dynamics by taking every 10th data point.
-Instead, averaging over a 10-step window smooths out fast variability and effectively defines a new, coarse-grained “more Markovian” system.
-We then re-train the neural models with additive and multiplicative noise on the new coarse-grained data.
-Next, we simulate the neural models over long timescales to analyze their stationary distributions, as well as their ensemble mean and variance responses to impulse perturbations.
+As an additional step, we average $x_{1}(t)$ over a short temporal window of 10 time steps (equivalent to $0.1$ Model Time Units).
+This averaging smooths out fast variability and effectively defines a coarse-grained system that is “more Markovian”.
+We then retrain the neural models with additive and multiplicative noise on the coarse-grained data.
 Results are shown in Figure 5.
-As before, the stochastic model with additive noise fails to capture the correct stationary statistics (Figure 5(a)).
-While the response in the ensemble mean improves relative to the non-coarse-grained case, the response in ensemble variance remains suppressed.
-In contrast, the model with multiplicative noise not only captures the stationary distribution and further improves the mean response, but now also closely reproduces the response in variance, capturing the peak magnitude and deviating only slightly in its temporal decay.
-Note that the response in ensemble mean and variance to impulse perturbations of the original degree of freedom $x_{1}$ are the same in both Figures 4 and 5, the only thing changing in the latter is that the reduced-order model is defined in a coarse-grained dynamics.
+As before, the stochastic model with additive noise fails to capture the correct stationary and perturbed statistics.
+By contrast, the model with multiplicative noise not only reproduces the stationary distribution and improves the mean response, but also closely matches the variance response, capturing its peak magnitude and deviating only slightly in its temporal decay.
+One could, in principle, tune the averaging window to match the autocorrelation of the emulated $x(t)$ with that of the original $x_{1}(t)$, thereby ensuring closer agreement in the decay of responses (Figure 5(e,f)).
+We view this as a methodological refinement for future work, and leave it outside the scope of this study.
 Figure 5: Same as in Figure 4 however here the scalar emulators in Eq.
@@
 Even with access to very long time series, modeling real-world dynamical systems from data alone remains fundamentally constrained by observing only a subset of the full state vector.
-This leads to two major challenges: (i) the resulting data-driven model will depend heavily on the specific variable chosen for modeling, and (ii) once a suitable slow variable is identified, the cumulative effect of the unobserved degrees of freedom must be appropriately parametrized.
-These issues, demonstrated at the “practical level” within the context of a low-dimensional model, become more severe in realistic high-dimensional systems, which often lack a clear time-scale separation.
-These challenges are intrinsic to data-driven approaches for dynamical systems, and simply extending time series length or training increasingly large neural networks is unlikely to resolve them.
-Crucially, proposing more sophisticated models can be effective only if the proper variables are first identified.
-The core difficulty is conceptual: there is no rigorous mathematical framework for selecting the proper variables to model from data in real-world, high-dimensional dynamical systems.
-In practice, such choices rely on physical insight and domain expertise.
-In a spatiotemporal system like climate, this involves identifying an appropriate coarse-grained description of the dynamics, including relevant fields and a limited range of spatial and temporal scales.
-In the context of climate dynamics, it is sometimes possible (depending on the specific problem) to leverage the governing equations to guide variable selection.
-However, even in this case, one faces the practical necessity of resolving only a limited range of scales and introducing parametrizations.
-Within this framework, stochastic emulators offer a promising approach for data-driven simulation of climate systems and their sensitivity to external perturbations.
+This creates two major challenges: (i) the model’s skill depends critically on the variables chosen, and (ii) once these variables are fixed, the influence of unobserved degrees of freedom must be adequately parametrized.
+While such issues are illustrated with low-dimensional models, they become far more acute in realistic spatiotemporal climate systems, where time-scale separation is weak and variable selection involves nontrivial choices of fields, dimensionality reduction methods, and spatiotemporal scales.
+There is no rigorous mathematical framework to guide these choices from data alone.
+Even in cases where the governing equations are known, such as in meteorology, and where large computational resources are available, emulators still operate on finite-resolution data, leading to subgrid processes.
+The cumulative effect of such unobserved processes still require parametrizations.
+This highlights that the problem of parameterizations, traditionally associated with numerical modeling [92, 93], is equally central to data-driven approaches with neural emulators.
+These considerations highlight intrinsic limitations of general-purpose emulators in capturing responses to external perturbations and in supporting Green’s function–like experiments [38].
+In contrast, reduced-order neural models are often tailored to specific problems, allowing one to leverage prior literature, insights, and results to guide variables and coarse-graining choices.
+In the next section, we present a concrete example where a simple reduced-order neural model succeeds in a task where a state-of-the-art neural emulator showed limitations [37].
 IV A real-world example: radiative responses to temperature perturbations in a climate model
-Building on the ideas discussed above, we present a stochastic reduced-order model tailored to a real-world application.
-The emulator is designed to capture the large-scale dynamics of sea surface temperature (SST) and the global mean net radiative flux at the top of the atmosphere (TOA).
-To achieve this, we restrict our focus to a limited range of spatiotemporal scales by selecting a few dominant modes, using monthly-averaged data, and high-pass filtering to remove multidecadal variability.
-The cumulative impact of fast, unresolved processes on the slow dynamics is parametrized by multiplicative, state-dependent noise.
-We then evaluate the model’s response to temperature changes and assess its utility in improving our understanding of radiative feedbacks.
+Building on the insights discussed above, we present a stochastic reduced-order model tailored to a specific real-world test-case.
+We first introduce the scientific problem at hand and the form of the proposed emulator to tackle this specific question.
+We then highlight the coarse-graining choices defining the proper variables to model.
+Finally, we showcase the skill of this emulator to study causal mechanisms in large-scale climate dynamics.
 IV.1 Scientific problem at hand
-Radiative forcing (e.g., as resulting from a change in $\text{CO}_{2}$ concentration) drives warming in the Earth system, leading to changes in sea surface temperatures (SSTs) and associated radiative feedbacks.
-Recent studies have shown that even given the same global mean temperature warming, different spatial patterns of SST warming can produce markedly different radiative feedbacks.
-For example, warming concentrated in the eastern Pacific tends to produce positive radiative feedbacks, thereby amplifying the initial temperature change.
-In contrast, greater warming in the western Pacific or warm-pool region tends to have a stabilizing effect [29, 30].
-This sensitivity of the Earth’s energy imbalance to the spatial pattern of surface temperature change is commonly referred to as the “pattern effect”.
-The pattern effect captures the fast response of top-of-atmosphere (TOA) radiative fluxes to spatial variations in temperature.
-It is theoretically defined in the framework of energy balance models by Taylor expanding the radiative response over the steady state temperature (see for example Section 2 of Meyssignac et al.
-(2023) [103]).
-A general data-driven framework to explore radiative responses from data has recently been proposed in [104], with the pattern effect arising as a limiting case in the case of fast responses.
-The framework in [104] focuses on both direct causal links between SST and radiative fluxes as well as cumulative (in time) links.
-A widely used approach to study the relationship between SST patterns and TOA radiative fluxes is through Green’s function experiments [105], in which the radiative response to localized SST perturbations is quantified using atmosphere-only models [106, 107, 108, 30, 109].
-These experiments are extremely valuable for rigorously testing model sensitivities.
-However, they are computationally expensive and typically require large SST perturbations (e.g., 1–4 K) to generate a measurable response within short simulations.
-These large perturbations can lead to nonlinear responses, invalidating the linearity assumptions adopted in the Green’s function framework [110].
-There is therefore a need for fast and interpretable data-driven models capable of performing large-ensemble experiments at low computational cost and with small perturbation amplitudes.
-Moreover, climate models are known to be biased, while data-driven emulators offer, in principle, a way to learn climate dynamics directly from observations.
-There are two broad strategies to build such emulators.
-One is to construct a general-purpose emulator of the entire climate system and then use it to study specific problems.
+The “pattern effect” problem.
+Radiative forcing (e.g., from changes in $\text{CO}_{2}$) drives surface warming, which alters sea surface temperatures (SSTs) and associated radiative feedbacks.
+Recent studies show that even for the same global-mean warming, different SST warming patterns can yield markedly different feedbacks.
+For example, eastern Pacific warming tends to amplify warming through positive feedbacks, while western Pacific (warm pool) warming has a stabilizing effect [38, 39].
+This sensitivity of Earth’s energy imbalance to spatial SST patterns is known as the “pattern effect.” The pattern effect describes the fast response of top-of-atmosphere (TOA) radiative fluxes to spatial temperature variations and can be formalized in energy balance models via a Taylor expansion around the steady-state temperature [94].
+A common approach to study the pattern effect is to relate SST patterns to TOA radiative fluxes using Green’s function experiments [95], where the radiative response to localized SST perturbations is quantified with atmosphere-only models [96, 97, 98, 39, 99].
+These experiments are valuable for testing model sensitivities but are computationally costly and typically require large SST perturbations (1–4 K) to obtain a measurable response within short simulations.
+Such large perturbations can trigger nonlinear behavior, undermining the linearity assumptions of the Green’s function framework [100].
+This motivates the need for fast, interpretable data-driven models that can simulate responses of large ensembles at low cost and with small perturbation amplitudes.
+Furthermore, since climate models are known to be biased, data-driven emulators offer, in principle, a complementary way to learn climate dynamics directly from observations.
+Possible emulator strategies.
+Two broad strategies exist for building emulators.
+One is to construct a general-purpose emulator of the climate system and then apply it to specific problems.
 However, as shown by Van Loon et al.
-(2025) in the context of the ACE2 emulator [28], this strategy may not always succeed.
-A second, more traditional and targeted approach is to design an emulator focused on the problem at hand, e.g.
+(2025) [37] in the context of the ACE2 emulator and the pattern effect, this strategy may not always succeed.
+A second, more targeted approach for large-scale climate dynamics is to design an emulator tailored to a given problem, e.g.
 [5].
-While this approach sacrifices generality, it offers important advantages: the focus on a specific problem often comes together with a greater understanding of what are the proper variables to model as well as the spatial and temporal scales to consider.
-These choices, typically informed by physical intuition and prior literature, can directly shape the architecture and inputs of the emulator.
-Moreover, by isolating only core relevant processes for the specific problem considered, such an emulator can remain interpretable, thereby helping to bridge the gap between simulation and understanding [55].
+While this sacrifices generality, it offers key advantages: clarifying the choice of variables and scales to model, grounding the design in physical intuition and prior literature, and isolating only the processes most relevant to the problem.
+This focus also enhances interpretability and helps bridging the gap between simulation and understanding [101].
 IV.2 A reduced order model of surface temperature and radiative fluxes
-Given the general problem outlined above, we aim to build a skillful emulator of sea surface temperature and radiative fluxes that allows us to study stationary and perturbed statistics.
-Building on the previous Sections we consider a functional form of the neural model defined follows:
+We aim to build a skillful emulator of large-scale sea surface temperature (SST) and radiative fluxes capable of capturing both stationary and perturbed statistics.
+Building on the previous sections, we adopt the following functional form for the neural model:
 $d\mathbf{x}=(\mathbf{L}\mathbf{x}+\mathbf{n}(\mathbf{x}))dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W},$ (11)
-Where $\mathbf{W}$ is the standard Wiener process.
-As in the previous Sections, we use a neural network to fit the drift part $(\mathbf{L}+\mathbf{n}(\mathbf{x}))$ and another network to fit the multiplicative noise $\mathbf{\Sigma}(\mathbf{x})$.
-As before, to improve readability we refer to Appendix A for technical details.
-We now face the challenging task of selecting appropriate preprocessing steps to define the coarse-grained dynamics to be modeled.
-These coarse-graining choices are crucial, as the real climate system involves not only SST and radiative fluxes, but also many other interacting variables, i.e.
-this is a practical example of the case of a partially observed state vector.
-We now describe the dataset used for this experiment, along with the preprocessing steps taken.
-The preprocessing procedure largely follows that described in [104] and we will refer the reader to that work for additional details and clarifications where needed.
-IV.2.1 Data, preprocessing and coarse-graining
+Where $\mathbf{W}$ is the standard Wiener process. As before, to improve readability we refer to Appendix C for technical details.
+As outlined throughout the paper, a central challenge lies in selecting preprocessing steps to define the coarse-grained dynamics $\mathbf{x}$ to be modeled.
+These choices are critical for identifying the proper variables, since the climate system involves not only SST and radiative fluxes but also many other interacting variables.
+This is thus a practical instance of a partially observed state vector.
+Below, we describe the dataset and preprocessing procedure, which largely follows [102], with further details provided in that work.
+IV.2.1 Data, preprocessing and the choice of proper variables
 Data.
-The focus is on a state-of-the-art coupled climate model GFDL-CM4 [111].
-The model offers data from a long control run simulating the coupled climate system for 600 years with constant $\text{CO}_{2}$ fixed at pre-industrial level.
-The ocean component is the MOM6 ocean model [112] with a horizontal grid spacing of 0.25 and 75 vertical layers.
-The atmospheric component is the AM4 model [113, 114] with a horizontal grid spacing of roughly 100 km and 33 vertical layers.
-The model considered in Eq.
-(10) will be trained on the long control run.
-The variables considered are two: SST and the global mean net radiative flux at the top of the atmosphere (hereafter, often referred to as TOA).
-The focus on global mean net radiative flux, rather than its spatiotemporal representation is motivated by the fact that such variable does not show large scale patterns of variability.
-It is in fact a fast variable in respect to SST and the main signal we are interested in is its global average.
-The SST field is considered only in the latitudinal range $[-45,45]$, a choice informed by earlier studies highlighting the prominence of radiative feedbacks in the tropics and subtropics [30, 107].
-Preprocessing and coarse-graining.
-Climate dynamics is not only encompassing the SST and radiative fluxes.
-Therefore, we aim in building a reduced order model focusing on the slow dynamics and parametrizing unobserved degrees of freedom as noise.
-We consider the control run and proceed in coarse-graining both the spatial and temporal scales.
-First, starting from daily data, we compute monthly averages.
-We then remove the seasonal cycle from both the SST and radiative fluxes, allowing us to focus on seasonal anomalies, rather than periodic signals driven by the annual solar forcing.
-We then decompose the SST spatiotemporal field in a set of $n$ modes, i.e.
-empirical orthogonal functions (EOFs).
-Projecting the computed EOFs onto the original data give us $n$ Principal Components (PCs).
-In formulae, given the SST field $\mathbf{X}(t)\in\mathbb{R}^{N}$, we decompose it as:
-$\begin{split}\mathbf{X}(t)=\sum_{i=1}^{n}x_{i}(t)\mathbf{e}_{i}+\sum_{j=n+1}^{N}y_{j}(t)\mathbf{e}_{j},\end{split}$ (11)
-and retain only the first $n=20$ modes $x_{i}(t)$, while the remaining components $y_{j}(t)$ represent the fast modes.
-$\mathbf{e}$ are orthonormal bases (i.e.
-the EOFs).
-Therefore, the time series $x_{i}(t)$ will be used to train the emulator in Eq.
-(10).
-These two steps, i.e.
-one month averaging and the projection over EOFs, attempt to define a new coarse-grained system for the slow variability of the SST field.
-Note that Penland and Sardeshmukh [4] and Penland (2006) [5] already showed how a similar preprocessing would lead to a skillful linear Markov model of the tropical Pacific SST.
-These SST principal components are then concatenated with the scalar net radiative flux time series: to simplify the notation, we are going to refer to the state vector of the reduced order model as simply as $\mathbf{x}(t)\in\mathbb{R}^{n+1}$.
-Is a 600 year-long control run a long-enough simulation to learn a model of SST and radiative fluxes?
-Whether this is a sufficiently long simulation depend on whether the time series considered sample enough recurrences, see Cecconi et al.
-(2012) [31].
-In a multiscale system like climate, even in the presence of constant external forcing, we observe processes such as multidecadal oscillations which are only sampled a few times in a 600-year long experiment.
-Therefore, we make the following choice: we high-pass filter the data with a cut-off frequency of $10^{-1}$ years.
-This removes all low-frequency variability which we would have little chance to learn from data in the first place, and focus on variability inside 10 years only.
-This choice involves a trade-off: our model will not be able to capture multidecadal variability and we will restrict ourselves to explore subdecadal variability only.
-At the same time this leads to an effectively larger sample size than we would have had by keeping all time scales.
-Additionally, this step also allow us to remove drifts present in the control run of the CM4 run, mainly coming from an ocean component that still did not equilibrate.
-In this paper we will focus on testing the ability to “generalize” to unseen climate scenarios: we will fix the SST variable with a precomputed trajectory coming from two, 150 years long forced runs, namely a $1$pctCO2 and 4×4$\times$4 ×CO2, and use this information to reconstruct the change in global mean TOA in these simulations.
-The $1$pctCO2 is an idealized experiment simulating the climate system under a 1% $\text{CO}_{2}$ increase per year starting from pre-industrial conditions.
-The 4×4$\times$4 ×CO2 is an idealized experiment simulating the climate system under an abrupt quadrupling of $\text{CO}_{2}$ concentration from its pre-industrial level at time $t=0$.
-Importantly, considering the $1$pctCO2 and 4×4$\times$4 ×CO2 simulations leads to an additional challenge: the radiative fluxes are directly influenced by an increase in atmospheric $\text{CO}_{2}$.
-However, our reduced order model only has access to SST and radiative fluxes and it cannot explicitly include $\text{CO}_{2}$ concentration.
-To address this, we remove the portion of the net TOA radiative flux that is attributable solely to rising $\text{CO}_{2}$.
+We consider outputs from the state-of-the-art coupled climate model GFDL-CM4 [103], which provides a 600-year pre-industrial control run with constant $\text{CO}_{2}$.
+The ocean component is MOM6 [104] with 0.25 horizontal resolution and 75 vertical layers, and the atmosphere is AM4 [105, 106] with  100 km resolution and 33 vertical layers.
+The model in Eq.
+(11) is trained on this control run.
+We consider two climate fields: sea surface temperature (SST) and the global mean net radiative flux at the top of the atmosphere (TOA).
+The SST field is restricted to latitudes $45^{\circ},45^{\circ}]$, following earlier studies that highlight the importance of radiative feedbacks in the tropics and subtropics [39, 97].
+Choosing the proper variables.
+Climate dynamics involves many interacting fields beyond SST and radiative fluxes.
+Our goal is a reduced-order model that captures the coarse-grained dynamics while representing unresolved processes as noise.
+We coarse-grain the GFDL-CM4 control run in space and time, averaging daily data to monthly means and removing the seasonal cycle to focus on anomalies rather than periodic signals driven by solar forcing.
+Considering 1 month averages in TOA fluxes is not arbitrary: the immediate response of TOA fluxes to perturbations in SST is fast (e.g., $\sim 20$ days in tropical large-scale convective overturning motion in the GFDL-CM4 model) [102].
+This averaging filters out higher-frequency processes that mediate surface–TOA coupling, yielding a system that is more Markovian.
+Given so, direct causal links between TOA fluxes at SST patterns can be detected as one time step responses; see Appendix A.
+We then high-pass filter with a cut-off of $10^{-1}$ years, removing (i) low-frequency (multidecadal) oscillations, which are poorly sampled even in 600-year runs, and (ii) slow drifts from incomplete ocean equilibration.
+This choice sacrifices the ability to capture multidecadal processes but ensures robust sampling of subdecadal statistics, yielding a larger effective sample size.
+Implicitly, the choice assumes that the response of the net radiative flux to SST perturbations occurs within a 10-year window.
+Next, we decompose the SST field $\mathbf{X}(t)\in\mathbb{R}^{N}$ into empirical orthogonal functions (EOFs):
+$\mathbf{X}(t)=\sum_{i=1}^{n}x_{i}(t)\mathbf{e}_{i}+\sum_{j=n+1}^{N}y_{j}(t)\mathbf{e}_{j},$ (12)
+and retain the first $n=20$ modes $x_{i}(t)$.
+The combined steps of monthly averaging, high-pass filtering and projection over EOF bases define a coarse-grained system for the slow SST variability.
+This choice follows earlier work showing the skill of linear Markov models built on similar preprocessing [4, 5].
+For radiative fluxes, we focus on the global mean TOA net flux, as the TOA flux exhibits little spatial structure and we are mainly interest in its average value.
+The state vector is then defined as $\mathbf{x}(t)\in\mathbb{R}^{n+1}$, consisting of the $n$ SST principal components concatenated with the global mean TOA flux.
+Forced simulations
+We test the model’s ability to reconstruct changes in global-mean TOA radiative flux given prescribed SST trajectories from two 150-years long forced runs: 11pctCO2 and $4\timesCO$2.
+The 11pctCO2 experiment imposes a 1% $\text{CO}_{2}$ increase per year starting from pre-industrial conditions, while the $4\timesCO$2 experiment abruptly quadruples $\text{CO}_{2}$ at $t=0$.
+A complication is that TOA fluxes respond directly to rising $\text{CO}_{2}$, whereas our reduced-order model only uses SST and fluxes and does not explicitly include $\text{CO}_{2}$.
+To address this, we remove the portion of change of the net TOA radiative flux that is attributable solely to rising $\text{CO}_{2}$.
 This allow us to isolate and study the component of radiative feedback that is driven by changes in SST.
-This correction relies on the well established fact that the global-mean radiative forcing from $\text{CO}_{2}$ alone increases approximately logarithmically with its concentration [115, 116].
-As is standard, for the 4×4$\times$4 ×CO2 experiment, we remove a standard constant of $8Wm^{-2}$ [117].
-For the $1$pctCO2 run we instead apply a time-dependent correction of the form $\alpha\log[C_{t}/C_{0}]$; $C_{0}$ and $C_{t}$ respectively representing the concentration of $\text{CO}_{2}$ in the control run and forced experiments.
-We note that considering 1 month averages in TOA fluxes is not arbitrary: the immediate response of TOA fluxes to perturbations in SST is fast (e.g., $\sim 20$ days in tropical large-scale convective overturning motion in the GFDL-CM4 model).
-Therefore, reconstructing the net radiative fluxes given information of the SST in the two forced runs above is a relatively easy task for this system and it converts to a 1 time step prediction.
-For this point we refer the reader to the conceptual discussion in Section 7 of [104].
-In other words, the preprocessing considered here maximizes the chance of having a data-driven model that correctly reconstruct the TOA variable in unseen climates.
+This correction exploits the logarithmic dependence of global-mean forcing on $\text{CO}_{2}$ concentration [107, 108].
+As is standard, we subtract $8\text{Wm}^{-2}$ in the $4\timesCO$2 run [109], and apply a time-dependent correction $\alpha\log(C_{t}/C_{0})$ in the 11pctCO2 run, where $C_{t}$ and $C_{0}$ denote $\text{CO}_{2}$ concentrations in the forced and control simulations.
 IV.2.2 Stationary statistics and forced responses
 Stationary statistics.
-Given the $n+1$ dimensional state vector $\mathbf{x}(t)\in\mathbb{R}^{n+1}$ defined in the preprocessing step above, we can now train the emulator defined in Eq.
-(10) on the long control run of the GFDL-CM4 model.
-The trained emulator is then used to simulate a 600 years long climate trajectory (same length as the one used in the original model GFDL-CM4).
-The stationary probability distributions of the first SST mode and of the net radiative flux at the TOA are found to be qualitatively well captured, see Figure 6(a,b).
-The largest discrepancies between the emulator and the original model are found for the autocorrelation functions.
-Autocorrelation functions are shown for the first SST mode (largely corresponding to the El Niño Southern Oscillation (ENSO) [118]) and for the TOA variable.
+Given the $n+1$ dimensional state vector $\mathbf{x}(t)\in\mathbb{R}^{n+1}$ defined above, we train the emulator in Eq.
+(11) on the long control run of GFDL-CM4.
+The trained model is then used to simulate a trajectory of equal length (i.e.
+600 years at monthly time step).
+The stationary probability distributions of the leading SST mode and the net radiative TOA flux are well reproduced (Fig.
+6a,b).
+The main discrepancies arise in the autocorrelation functions, shown for the first SST mode (closely linked to ENSO [110]) and the TOA flux.
 Forced responses.
-The pattern effect problem discussed in IV.1 concerns identifying a direct causal link between patterns of SST warming and changes in radiative fluxes at the TOA.
-Before attempting any mechanistic interpretation or physical understanding, we first assess whether the neural model we have built possesses predictive skill in capturing forced responses.
-Specifically, we test if the neural model can be used to reconstruct the global mean radiative flux from a pre-computed, forced sea surface temperature field in forced experiments.
-In doing so, we also test an important question in neural models: is the neural network model able to generalize to out-of-distribution predictions?
-To answer these question we consider the two, 150-years long, forced runs of the GFDL-CM4 climate model introduced in Section IV.2.1.
-We extract the sea surface temperature from these forced runs and project it onto the set of $n=20$ EOF bases found in the control run.
-We refer to such forced modes as $\mathbf{x}_{\text{SST}}^{\text{f}}$.
-We then simulate the neural emulator for 150 years by constraining the SST degrees of freedom to be exactly equal to $\mathbf{x}_{\text{SST}}^{\text{f}}$.
-Given $\mathbf{x}_{\text{SST}}^{\text{f}}(t)$ the model returns the TOA variable, $x_{\text{TOA}}(t+1)$, at the next time step, starting from initial condition $x_{\text{TOA}}(0)=0$.
-The results is shown in Figure 6(e,f).
-In both the case of $1$pctCO2 or 4×4$\times$4 ×CO2 runs, the model is able to reconstruct the change in radiative fluxes at the TOA in a quantitative way.
-We highlight that our preprocessing allowed us to reduce this reconstruction step into a prediction of the $21$-st entry of the state vector at time $t+1$, given information of the first 20 entries at the previous time $t$.
-This is only possible because of our focus on 1-month averages: the radiative fluxes at the TOA largely respond already at a sub-monthly time scales to temperature perturbations, therefore allowing us to focus on a direct statistical link $\mathbf{x}_{\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$.
-We also note that it is perhaps surprising that the neural network is able to respond faithfully even when for nonstationary inputs and with temperature values far out of the training distribution (in physical units, the mean temperature anomaly shifts from a $0K$ to $5K$ from the control in both forced simulations).
+The pattern-effect problem (Section IV.1) concerns identifying causal links between SST warming patterns and TOA flux changes.
+Before using our emulator for Green’s function–like studies to investigate the pattern-effect, we first assess its skill as a reconstruction tool.
+Specifically, we reconstruct global-mean TOA flux from prescribed SST trajectories in the two forced experiments, i.e.
+11pctCO22 and $4\timesCO$22.
+Forced SSTs are projected onto the first $n=20$ EOFs from the control run, yielding coarse-grained trajectories $\mathbf{x}_{\text{SST}}^{f}(t)$, $f$ denotes “forced.” Given $\mathbf{x}_{\text{SST}}^{f}(t)$, the emulator returns the TOA variable $x_{\text{TOA}}(t+1)$ at the next time step, initialized with $x_{\text{TOA}}(0)=0$.
+Results (Fig.
+6e,f) show that the emulator largely reproduces TOA flux changes in both scenarios.
+This reconstruction is possible because we focus on one-month averages, which reduce the time-dependent TOA response, driven by an atmospheric reorganization at sub-monthly scales, into a one-step prediction.
+For the 1pctCO2 experiment, the linear trend in net TOA flux is $0.043Wm^{-2}yr^{-1}$ in GFDL-CM4 and $0.034Wm^{-2}yr^{-1}$ in the emulator.
+We then compute correlations between reconstructed and simulated radiative flux at the TOA, after removing a linear trend in 1pctCO2 and an initial 50-year transient in 4xCO2.
+At monthly resolution, correlations are $r=0.37$ (1pctCO2) and $r=0.43$ (4xCO2).
+If yearly averages are considered (as in the current literature, see e.g.
+[38]) these increase to $r=0.70$ (1pctCO2) and $r=0.74$ (4xCO2).
+All correlations are significant at the $5$ sigma level under a null hypothesis of zero correlation, tested via bootstrap resampling with replacement.
 Figure 6: Panel (a): stationary distributions obtained for the 1st principal component of the sea surface temperature (SST) field as in the original GFDL-CM4 climate model and as emulated by the neural emulator with multiplicative noise in Eq.
@@
 (11) after constraining the SST degrees of freedom to follow a precomputed trajectory in two forced experiments in the GFDL-CM4 model.
-The forced experiments considered are the $1$pctCO2 and 4×4$\times$4 ×CO2, see Section IV.2.1 for details.
+The forced experiments considered are the 11pctCO2 and $4\timesCO$2, see Section IV.2.1 for details.
 Note: the net flux at the TOA variable is negative upward (i.e., leaving the planet).
-IV.2.3 Interpretability and understanding
-In the previous Section we showed that the neural emulator proposed in this study can in fact be used to reconstruct the instantaneous change in radiative feedback at the top of the atmosphere given pre-computed changes in the temperature field.
-As emphasized before, the task of the emulator in the reconstruction/prediction task reduces to the following mapping: $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\in\mathbb{R}^{n}$, $x_{\text{TOA}}(t)\in\mathbb{R}$ and with initial condition $x_{\text{TOA}}(0)=0$.
+IV.2.3 Interpretability and understanding: ensemble mean and variance responses
+In the previous section, we showed that the reduced-order neural emulator can reconstruct the net radiative flux at the top of the atmosphere (TOA) from prescribed changes in the sea surface temperature (SST) field.
+This amounts to the mapping $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $\mathbf{x}_{\text{SST}}(t)\in\mathbb{R}^{n}$, $x_{\text{TOA}}(t)\in\mathbb{R}$, and $x_{\text{TOA}}(0)=0$.
 We now ask the following question: What are the SST patterns that contribute the most to this skillfull prediction?
-To answer this question, we must go beyond the prediction task presented above and identify which components of the sea surface temperature input contributes the most to changes in radiative flux.
-We explore this question through the framework of linear response theory, by capturing the direct causal link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $x_{j,\text{SST}}(t)$ represents the $j$-th component of the surface temperature field at time $t$.
-We briefly describe the strategy to infer this direct causal link with response theory, and refers to Section II.2 for details.
-We consider the neural emulator in Eq.
-(10) and iteratively perturb the $j$-th sea surface temperature component of the state vector at time $t=0$ with a small impulse perturbation $x_{j}(0)+\Delta_{j}\delta(t)$.
-$\delta(t)$ being a Dirac delta and $\Delta_{j}$ being a perturbation of very small amplitude.
-The amplitude $\Delta_{j}$ depends on the component $x_{j,\text{SST}}$ itself (Section II.2).
-We then compute the ensemble mean response of the net radiative flux at the TOA at the next time step $t=1$, referred here as $\delta\langle x_{\text{TOA}}(1)\rangle$.
-To estimate the ensemble mean response we consider $10^{3}$ initial conditions.
-We write this response as follows:
-$\begin{split}R_{\text{TOA},j}(1)=\frac{\delta\langle x_{\text{TOA}}(1)\rangle}{\Delta_{j}\delta(t)},\end{split}$ (12)
-ad remind the reader that the index $j$ only considers the SST degrees of freedom.
-Therefore, the response at 1-time-step $R_{\text{TOA},j}(1)$ defines the direct causal link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$ between the $j$-th degree of freedom in SST and the net radiative flux at the TOA.
-The vector $\mathbf{R_{\text{TOA}}}(1)\in\mathbb{R}^{n}$ is projected back to the higher, original $N$-dimensional space.
-To do so we consider the previously computed EOFs, i.e.
-the orthonormal vectors $\mathbf{e}\in\mathbb{R}^{n,N}$ and project the 1-time-step responses as follows:
-$\begin{split}\mathbf{S}=\mathbf{R_{\text{TOA}}}(1)\cdot\mathbf{e}=\sum_{j}^{n}R_{\text{TOA},j}(1)~{}\mathbf{e}_{j},\end{split}$ (13)
-where $\mathbf{S}\in\mathbb{R}^{N}$ is a sensitivity map defined in the high-dimensional $N$ dimensional space.
-Such map approximates at each grid point $i$ in the ocean, the sensitivity $\frac{\partial x_{\text{TOA}}}{\partial x_{\text{SST},i}}$, where $x_{\text{TOA}}$ is the global mean radiative flux at the TOA and $x_{\text{SST},i}$ is the SST at grid point $i$.
-The sensitivity map $\mathbf{S}$ obtained using the neural emulator in Eq.
-(10) is shown in Figure 7.
-The obtained sensitivity map is in qualitative agreement with previous results in response theory (Figure 2(a) in [104]), with other data-driven approaches (see for example Figure 7(b) in [119] and Figure 2 in [120]), but, most importantly, with the sensitivity map obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [29], see Figure 5(a) in [30]).
-Note that in the latter case, the maps exhibit fewer spatial features due to the coarse resolution of the perturbation patches, which span tens of degrees.
-The marked feature of this sensitivity map is the sensitivity dipole found the tropical Pacific, with negative values on the western side of the basin and positive values on the eastern side, see for example [30].
-This means that an initial warming in the eastern Pacific basin will result in a direct, positive radiative feedback, further amplifying the imposed temperature change; the opposite is true for the western Pacific.
-The mechanistic explanations for such mechanisms can be found for example in [121, 110] and we refer the interested reader to these papers for details.
-In brief, the positive sensitivity in $\mathbf{S}$ in the eastern Pacific is associated with a high-climatology of low-level clouds: a warming perturbation is expected to decrease the cloud cover, leading to anomalies in incoming shortwave radiation and therefore a positive sensitivity.
-On the other hand, the western Pacific is a region dominated by deep convective phenomena.
-Because of this, a surface temperature warming in the western Pacific basin tend to travel vertically through the troposphere and then horizontally through gravity waves.
-This process strengthen the inversion layer in the eastern Pacific basin (and in other regions with high-climatology of low-level clouds), leading to a larger low-level cloud fraction and increasing the reflected shortwave at the top of the atmosphere.
-To conclude this Section, the analysis in Figure 7 confirms that the good reconstruction observed in the forced regime (Figure 6(e,f)) arises for the physically correct reasons.
-Figure 7: Sensitivity map obtained using the neural emulator in Eq.
+Addressing this question requires moving beyond prediction to identify the causal influence of individual SST components on radiative flux.
+We frame the problem in terms of linear response theory, focusing on the direct link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $x_{j,\text{SST}}(t)$ denotes the $j$-th SST component.
+Response in ensemble mean.
+To quantify this link, we perturb the emulator in Eq.
+(11) following the procedure in Appendix A.
+Briefly, at $t=0$, we apply a small impulse $\Delta_{j}=10^{-2}\sigma_{j}$ to the $j$-th SST component, with $\sigma_{j}$ its stationary standard deviation.
+We then compute the ensemble-mean response of the net TOA flux at $t=1$, denoted $\delta\langle x_{\text{TOA}}(1)\rangle$.
+This quantity, referred to as $R^{\mu}_{\text{TOA},j}(1)$, represents the ensemble-mean response ($\mu$ indicates the mean) and is estimated over $10^{6}$ initial conditions to ensure robust statistics
+Response in ensemble variance.
+While the ensemble-mean response is informative, it captures only part of the adjustment.
+Large-scale climate responses can involve shifts of probability distributions.
+To partially account for this, we quantify the response in the variance of the net TOA radiative flux to surface warming perturbations.
+As above, the estimate is based on an ensemble of $10^{6}$ trajectories.
+For consistency in physical units, we express the response in terms of standard deviation rather than variance, denoting it by $R^{\sigma}_{\text{TOA},j}(1)$ ($\sigma$ indicates the standard deviation).
+This quantity measures the change in ensemble spread following a perturbation of the $j$-th SST component.
+Conceptually, this procedure provides a computationally efficient analogue of atmospheric-only Green’s function experiments, while shifting the focus from individual trajectories to full probability distributions.
+Sensitivity maps.
+The vectors of one-time-step responses in mean and standard deviation, $\mathbf{R}^{\mu}$ and $\mathbf{R}^{\sigma}$, are projected back to the full $N$-dimensional SST field using the EOFs, $\mathbf{e}\in\mathbb{R}^{n,N}$:
+$\begin{split}\mathbf{S}^{\mu,\sigma}=\mathbf{R^{\mu,\sigma}_{\text{TOA}}}(1)\cdot\mathbf{e}=\sum_{j}^{n}R^{\mu,\sigma}_{\text{TOA},j}(1)~\mathbf{e}_{j},\end{split}$ (13)
+where $\mathbf{S}^{\mu},\mathbf{S}^{\sigma}\in\mathbb{R}^{N}$ represent sensitivity maps in mean and standard deviation, respectively.
+At each ocean grid point $i$, these maps approximate the sensitivities $\frac{\partial x^{\mu,\sigma}_{\text{TOA}}}{\partial x_{\text{SST},i}}$, where $x^{\mu}_{\text{TOA}}$ and $x^{\sigma}_{\text{TOA}}$ respectively represent the ensemble mean and standard deviation of the mean, net radiative flux and $x_{\text{SST},i}$ is the SST at location $i$.
+The sensitivity maps in mean ($\mathbf{S}^{\mu}$) and standard deviation ($\mathbf{S}^{\sigma}$) are shown in Figure 7.
+The obtained sensitivity map $\mathbf{S}^{\mu}$ is in qualitative agreement with previous results in response theory (Figure 2(a) in [102]), with other data-driven approaches (see Figure 7(b) in [111] and Figure 2 in [112]), but, most importantly, with the sensitivity map obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [38], see Figure 5(a) in [39]).
+In the latter case, the maps exhibit fewer spatial features due to the coarse resolution of the perturbation patches (spanning tens of degrees) in numerical models experiments.
+Figure 7: Sensitivity maps obtained using the neural emulator in Eq.
 (11).
-At each grid point $i$ we are plotting the 1 time step response in the global mean, net radiative flux at the top of the atmosphere (TOA) driven by an impulse perturbation in the sea surface temperature (SST) at that grid point.
-Note: the computation is done in a 21 dimensional space spanned by a set of empirical orthogonal functions and this map is a projection of the results obtained in the low-dimensional space.
-The response has been obtained by averaging over an ensemble of $10^{3}$ initial conditions.
+At each SST grid point $i$, we show the one–time-step response in the ensemble mean ($\mathbf{S}^{\mu}$) and ensemble standard deviation ($\mathbf{S}^{\sigma}$) of the global-mean net radiative flux at the TOA following an impulse perturbation in SST at that location.
+The computation is carried out in a 21-dimensional reduced space spanned by empirical orthogonal functions, and the maps represent the projection of these low-dimensional results back to the full grid.
+Responses are estimated from an ensemble of $10^{6}$ initial conditions.
+The dominant feature of $\mathbf{S}^{\mu}$ is a dipole in the tropical Pacific: negative sensitivity in the west and positive sensitivity in the east [39].
+This implies that eastern Pacific warming produces a positive radiative feedback, while western Pacific warming produces the opposite.
+The mechanisms are well established [113, 100]: in the east, high climatological low-level cloud cover decreases under warming, reducing reflected shortwave radiation and yielding positive sensitivity.
+In the west, deep convection transports surface warming vertically and then horizontally via gravity waves, strengthening the inversion over the eastern Pacific and other stratocumulus regions, increasing low-level cloud fraction and reflected shortwave radiation, thus producing a negative feedback.
+This analysis confirms that the accurate reconstructions in the forced regime (Fig.
+6e,f) arise from physically consistent mechanisms, and that coarse-grained neural models can recover such causal linkages.
+Finally, the variance response $\mathbf{S}^{\sigma}$ (Fig.
+7b) highlights regions where perturbations alter the spread of radiative responses across ensemble members.
+Perturbations in stratocumulus-dominated regions (northeast Pacific, off the coast of South America) yield negative variance responses: the ensemble distribution of net radiative flux narrows relative to the control, highlighting small uncertainties in the mean response studied above.
+By contrast, surface warming perturbations in the central and western Pacific lead to positive variance responses, reflecting a broader spread of the net radiative flux distribution after SST perturbations.
+To our knowledge, such variance responses have not been quantified before.
+Unlike the response in the mean, which has been examined extensively and shows broad agreement across atmospheric models [38], the response in ensemble variance has no established benchmark.
+Because it reflects changes in variability, it may differ significantly across atmospheric models.
+Moreover, estimating such responses requires very large ensembles, which are prohibitively expensive in conventional climate models.
+For this reason, the results presented here should be viewed as an exploratory step: they highlight potentially important regional structures, but a full climate interpretation will require systematic intercomparison and validation against other modeling frameworks, such as novel data-driven tools to estimate response through the fluctuation-dissipation theorem [114].
 IV.2.4 Main takeaways
@@
 These simplifications led to a stable and computationally efficient emulator that responds to external forcing and can be used to investigate causal relationships via perturbation experiments.
-The model reconstruct the net TOA radiative flux when forced with a pre-computed SST trajectory.
-Beyond reconstruction, we illustrated that the model can be used to perform perturbation experiment, allowing us to enhance mechanistic understanding of the underlying physical processes.
-Findings are in qualitative agreement with results from traditional perturbation experiments in numerical models, giving additional confidence in the emulator’s ability to capture key interactions across variables.
-V Conclusions
-In the first part of this paper, we addressed, at a pedagogical level, fundamental issues in neural emulation of dynamical systems with climate applications in mind.
-We considered two scenarios: (i) when the full state vector is observed, and (ii) the more realistic case of partial observations.
-Our findings can be summarized as follows:
+Such experiments focused on very large ensembles ($10^{6}$ trajectories), currently impossible to perform with climate models.
+These results demonstrate that reduced-order neural emulators can extend beyond reproducing stationary climate dynamics to enable efficient, ensemble-based perturbation experiments.
+V Discussion and Conclusion
+In summary, the first part of this work discussed the limitations of purely data-driven models of complex climate systems.
+We highlighted the potential of combining traditional reduced-order modeling approaches, including state-dependent (multiplicative) stochastic parameterizations, with recent advances in neural networks to study large-scale, low-frequency climate dynamics.
+We emphasized the need to evaluate such models against both stationary and perturbed statistics.
+In this context we discussed how response theory provides rigorous guidance to move beyond stationary evaluation and probe causal mechanisms.
+These ideas were then illustrated through a climate-dynamics problem focusing on inferring causal relations between surface warming and changes in net radiative fluxes.
+Given careful variable selection, reduced-order neural stochastic models can capture both stationary and perturbed statistics in a computationally efficient way.
+Below, we expand on these points and outline directions for future work.
+V.1 Limits of data-driven-only climate emulators and the role of stochastic approaches
+In the first part of this paper, we distinguished two possible scenarios in data-driven modeling:
   1. i.
-    Fully observed state vector.
-In this idealized case, it is possible to infer a dynamical model from data alone that reproduces the system’s stationary statistics and responds accurately to small, external perturbations.
+    Unknown equations, fully observed state vector.
+In this idealized setting, it is possible to infer a model from data alone that reproduces the system’s stationary statistics and responds accurately to small external perturbations.
   2. ii.
-    Partially observed state vector.
+    Unknown equations, partially observed state vector.
 This more realistic case presents fundamental difficulties, even with access to high-quality data and powerful neural models architectures.
 The model skill now depends critically on two factors: (a) the choice of variables to model and (b) how unobserved degrees of freedom are parameterized.
-These results highlight intrinsic limitations of purely data-driven modeling in high-dimensional systems, where the full set of relevant variables is typically unknown and there is no rigorous mathematical framework for selecting the proper variables to model.
-The identification of such variables, when possible, is inherently problem-dependent and often leads to adopting a coarse-grained representation of the dynamics.
-The next challenge is to parameterize the cumulative effect of unobserved degrees of freedom.
-This final step underscores that the issue of parameterization, traditionally associated with numerical modeling, is equally central to data-driven approaches using neural emulators.
-A promising perspective is to reconcile neural networks with principled modeling strategies, such as reduced-order modeling and stochastic parameterizations, by focusing on coarse-grained dynamics rather than constructing deterministic models that attempt to resolve as many spatiotemporal scales as possible.
-In addition, we argued that impulse response functions are a building block that a model should capture in order to reproduce causal interactions across degrees of freedom and to accurately respond to external perturbations.
+These results highlight intrinsic limitations of purely data-driven-only modeling in high-dimensional climate systems, where the full set of relevant variables is typically unknown and no rigorous framework exists for selecting the proper variables to model.
+By variable selection we mean the choice of climate fields as well as their spatial and temporal scales (also including dimensionality reduction methodologies).
+Such choices are guided by the physical process of interest, and in the case of low-frequency climate dynamics, they motivate the development of reduced-order neural emulators tailored to specific phenomena as efficient alternative to general purpose emulators.
+A central challenge is to parameterize the cumulative effect of unresolved degrees of freedom.
+Model performance (especially in capturing responses to external forcings) depends critically on this choice and we showed that multiplicative (state-dependent) noise is preferable to additive noise.
+We emphasize that the analysis presented here applies to data-driven models of turbulent geophysical systems that must capture both stationary statistics and responses to external perturbations, as in (causal) Green’s function experiments [38].
+For Earth system models, these issues mainly concern autoregressive climate emulators, particularly coupled atmosphere–ocean systems, and should not be confused with short-term weather forecasting using neural networks (e.g., [115, 23]), where long-term stability and forced responses are less relevant.
+Our study also does not address statistics-based emulators for climate change projections, which have proven effective for investigating responses of key climate variables to CO2 forcing [116, 117]; see [116] for an example where simple approaches, such as linear pattern scaling, can outperform more complex models.
+Future work.
+The discussion above suggests two possible directions for future research in autoregressive emulators of climate.
+First, a stochastic formulation of neural autoregressive climate emulators is preferable to a purely deterministic one.
+Probabilistic formulations align more closely with the inherently stochastic and uncertain nature of climate dynamics [118].
+Recent advances in weather emulation already point in this direction such as FourecastNet3 [23], GenCast [23] and the novel work of Alet et al.
+(2025) [119].
+Second, the ability of climate emulators in capturing forced responses in the full probability distribution, akin to Green’s function experiments, remains largely unexplored.
+A promising way forward is to evaluate novel architectures on their ability to respond to external perturbations on a hierarchy of simplified models relevant to geophysical fluid dynamics, such as the Lorenz–96 system [120].
+These controlled testbeds enable rigorous quantification of forced responses and could inform the design of future architectures for data-driven climate modeling.
+V.2 Response theory as a theoretical framework for model evaluation
+We argued that the impulse response operator is a building block that a model should capture in order to reproduce causal interactions across degrees of freedom and to accurately respond to external perturbations (Section II.2 and Appendix A).
 This perspective enables us to move beyond evaluation metrics based solely on stationary statistics and instead consider the effect of external interventions.
@@
 However, for models constructed from observational data, such direct perturbations are not possible.
-In this context, one could, in principle, leverage the Fluctuation-Dissipation Theorem (FDT) to estimate impulse response functions from data alone, at least at a qualitative level [38, 37, 35, 47, 62].
-In particular, it has been shown that even simple approximations of FDT, such as the quasi-Gaussian approximation [37], often yield qualitatively reliable estimates of impulse responses in the ensemble mean, without the need to perturb the system directly; see, for example, [122, 123, 104].
-That said, FDT itself is inferred from data and should therefore be viewed as a complementary, theory-based diagnostic (rather than a ground truth) allowing us to cross-validate the inferred responses of the emulator.
-In other words, consistency between the responses predicted by the emulator and those estimated via FDT can increase confidence in the model’s ability to capture the system’s response in the ensemble mean to external perturbations.
-In contrast, assessing impulse responses of higher-order moments (e.g., variance) via FDT remains substantially more challenging and calls for new methodological tools.
-A promising ongoing direction involves score-based generative modeling, as explored in [124, 125].
-Another fruitful avenue lies in data-driven frameworks at the intersection of Markov State Modeling and Koopman operator theory [126, 127].
-It is worth emphasizing that the central limitation discussed earlier, i.e.
+In this context, one could, in principle, leverage the Fluctuation-Dissipation Theorem (FDT) to predict the response operator from data alone [81, 17].
+Traditionally, the quasi-Gaussian approximation have yield reliable estimates of impulse responses in the ensemble mean [121, 47].
+Recent advances in score-based generative models have shown the potential to extend these predictions to responses in higher-order moments [114].
+Future work.
+In the absence of direct perturbation experiments, FDT provides a rigorous bridge between stationary and forced dynamics [49, 48].
+With sufficiently long integrations of current climate emulators, one can readily focus on representative large-scale modes and estimate their impulse responses from the stationary record, thereby extending evaluation beyond stationary statistics with minimal computational cost.
+A recent example illustrating this approach in the climate context can be found in [18].
+Moreover, the response operator at the shortest time step encodes direct causal links, i.e.
+the input–output relationships across variables in a reduced-order emulator (see Appendix A).
+Future work will leverage this information to suppress spurious, non-causal input-output dependencies during training, similarly to the recent work of Chen and Liu (2024) [59] on causally constrained, sparse reduced-order models.
+V.3 A realistic application: radiative responses to temperature perturbations
+Building on the theoretical insights introduced in the first part, the second part of the paper presented a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net radiative flux at the top-of-atmosphere (TOA).
+The model was trained on a long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies.
+The emulator focuses on the evolution of 20 principal components of the SST field and on the scalar global mean, net radiative flux; unobserved degrees of freedom are parametrized through multiplicative noise through an additional neural network.
+The emulator largely reproduces the stationary statistics of the system and, more importantly, allows us to reconstruct TOA radiative fluxes in two forced simulations.
+We inferred the direct causal links between temperature modes and TOA fluxes through a perturbation approach, akin to Green’s function-like experiments with climate models [38], targeting the so-called “pattern effect” [39].
+Crucially, the analyses targeted both responses in the ensemble mean and variance.
+Such responses are then projected back onto the EOFs to obtain maps of sensitivity in mean and variance change of the radiative flux at the TOA to SST warming perturbations.
+We argued that the sensitivity map for the ensemble mean is consistent with recent studies, indicating that the emulator’s skill in reconstructing TOA fluxes under forcing arises from physically meaningful mechanisms.
+In contrast, the sensitivity map of the ensemble variance is a novel result, not previously examined in the literature.
+It highlights the potential for uncertainty quantification in perturbation experiments but it requires further investigation to better understand the underlying physical mechanisms.
+These findings highlight the potential of reduced-order neural emulators to perform thousands of perturbation experiments, shifting the focus of climate feedback analysis from mean responses to the responses of probability distributions.
+Modeling the climate system from data alone is constrained by observing only a subset of variables of a multiscale, high-dimensional dynamical system.
+This raises the fundamental challenge of identifying the appropriate variables (both fields and spatiotemporal scales) a fundamental problem in physics, emphasized in classical work [122] and revisited in recent studies [123].
+However, large-scale climate dynamics is often well characterized by a few recurrent spatiotemporal patterns, or modes of variability.
+In this case, the multiscale structure of the climate system can be an advantage allowing us to derive data-driven models of the coarse-grained dynamics while treating unresolved, fast processes stochastically.
+Traditional reduced-order models stand out from general-purpose emulators by targeting specific climate problems from the outset, guided by prior literature, physical insight, and a clear understanding of the relevant variables and scales at play.
+Such data-driven modeling approach has a long history in climate science [2], though most efforts have emphasized reproducing stationary statistics or short-term forecasting [3, 4, 75] rather than forced responses.
+A promising perspective is to build on such established reduced-order approaches, focusing on powerful neural networks to fit deterministic and stochastic components and on tools from response theory, such as the FDT, to guide model development.
+Together, these strategies can enable efficient algorithms to study how full probability distributions respond to external perturbations, opening new avenues for causal inference and attribution in climate science.
+Appendix A Response operator: practical estimation and causal interpretation
+A.1 Inferring the response operator in practice
+Given a model $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W}$ we build the impulse response operator as follows:
+  1. 1.
+    We simulate a very long trajectory of the model, starting from a random initial condition and remove an initial transient.
+We then sample $N_{e}$ random points from the simulation to define an ensemble of $N_{e}$ initial conditions on the model’s attractor.
+$N_{e}$ should be very large (i.e.
+$N_{e}\gg 1$) in order to sample the whole attractor and approximate averages over the invariant distribution.
+  2. 2.
+    For each one of the $N_{e}$ initial conditions, we impose an impulse perturbation $x_{j}(0)\rightarrow x_{j}(0)+\Delta_{j}\delta(t)$ of size $\Delta_{j}$ on the $j$-th degree of freedom at time $t=0$.
+$\delta(t)$ is the Dirac delta function.
+The amplitude $\Delta_{j}$ should be theoretically infinitesimally small, ensuring linearity of the response even in nonlinear systems.
+In practice we do as follows: we consider the long time series of $x_{j}(t)$ from the long control integration above and define $\Delta_{j}=10^{-2}\sigma_{j}$, where $\sigma_{j}$ is the standard deviation of time series $x_{j}(t)$.
+  3. 3.
+    Therefore, for a given initial condition $\mathbf{x}(0)$, we simulate two trajectories: a control trajectory without perturbation and a perturbed one, where an impulse perturbation $\Delta_{j}$ has been imposed on the $j$-th degree of freedom at time $t=0$.
+This procedure is repeated in parallel for all initial conditions, resulting in an ensemble of $N_{e}$ paired of control and perturbed trajectories.
+  4. 4.
+    At each time $t$ we then estimate the time dependent ensemble average $\langle A(x_{k}(t))\rangle$ of an observable $A(x_{k}(t))$ for both the perturbed and unperturbed run.
+We refer to the perturbed and control ensemble average respectively as $\langle A(x_{k}(t))\rangle_{\text{p}}$ and $\langle A(x_{k}(t))\rangle$.
+The observables considered in this study are going to be: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution, therefore quantifying the response in ensemble mean and variance.
+  5. 5.
+    We define the impulse response operator for observable $A(x_{k}(t))$ as
+    $R_{k,j}(t)=\frac{\langle A(x_{k}(t))\rangle_{\text{p}}-\langle A(x_{k}(t))\rangle}{\Delta_{j}}.$
+A.2 Causal interpretation of the response operator
+Causality in natural systems is inferred through external interventions: an agent perturbs the system and the response is measured [77, 78].
+Climate modeling follows the same principle through Green’s function experiments, where one variable is perturbed and the response of another is evaluated [38, 124].
+Linear response theory formalizes this principle, providing a rigorous framework for causal analysis both in models and directly from data [17, 18, 114].
+Because responses are inherently time-dependent, causal links through responses can be distinguished in three forms, all encoded in the time-dependent response operator: (i) direct links, (ii) indirect links, and (iii) cumulative links (or causal sensitivity).
+In what follows, we use the triad system in Eq.
+(6) as a simple example to illustrate these ideas in a clear and unambiguous way.
+Finally, we summarize recent approaches for estimating responses directly from data and refer to Baldovin et al.
+(2020) [17] for details on the theoretical foundations of this approach.
+Consider a $N$ dimensional dynamical system of the kind:
+$\begin{split}\frac{dx_{i}}{dt}=f_{i}(x_{1},x_{2},...,x_{N});~\text{with}~i=1,2,...,N.\end{split}$ (14)
+For stochastic systems, causal links are contained in the deterministic drift $f_{i}$, so the discussion below applies to both deterministic and stochastic dynamics.
+In this setting, the existence of a direct causal link $x_{j}\rightarrow x_{k}$ is determined by whether $x_{j}$ appears in the arguments of $f_{k}$.
+More formally, the absence of a direct causal link $x_{j}\rightarrow x_{k}$ is expressed by:
+$\frac{\partial f_{k}}{\partial x_{j}}=0.$
+As a concrete example, we return to the triad system in Eq.
+(6), for which the underlying direct causal links are known exactly.
+This example provides a straightforward, intuitive illustration of how perturbation experiments reveal causal interactions.
+Figure 8 reports both (a) its functional form and (b) its causal graph, which encode the same information.
+$\frac{dx_{1}}{dt}$ $=f_{1}(x_{1},x_{2},x_{3})$
+$\frac{dx_{2}}{dt}$ $=f_{2}(x_{1},x_{2})$
+$\frac{dx_{3}}{dt}$ $=f_{3}(x_{1},x_{3})$
+(a) Functional form of the triad system in Eq. (6).
+(b) Causal graph of the triad system.
+Figure 8: Panel (a) Functional form of the triad system in Eq.
+(6).
+Panel (b) Graph structure of the triad system.
+Note from panel (b) that causality in dynamical systems can involve cycles, in contrast with the acyclic structures often assumed in applications outside physics, see e.g.
+[125].
+Causal relations can then be probed by adding independent impulse perturbations to each degree of freedom and record the system’s response.
+Responses are computed up to a long time horizon $\tau_{\max}$, with discrete times $t=0,dt,2dt,\dots,\tau_{max}$.
+Since we are interested in causal relations, it suffices to consider the identity observable $A(x_{k}(t))=x_{k}(t)$, i.e.
+the response of the ensemble mean [17].
+We now clarify the three different types of causal links mentioned at the start of this Section.
+Direct causal links.
+The direct causal links are encoded in the response operator $\mathbf{R}(dt)$ evaluated at the shortest time step $t=dt$.
+The intuition follows directly from the graph representation in Fig.
+8(b) and is rather straightforward.
+  • Any perturbation applied to $x_{2}$ will affect $x_{2}$ itself and $x_{1}$ after one time step, but not $x_{3}$. Hence, we infer $\partial f_{3}/\partial x_{2}=0$.
+  • A perturbation applied to $x_{3}$ will affect $x_{3}$ and $x_{1}$, but not $x_{2}$, so $\partial f_{2}/\partial x_{3}=0$.
+  • By contrast, a perturbation to $x_{1}$ propagates to all three variables within a single step, indicating direct causal links from $x_{1}$ to both $x_{2}$ and $x_{3}$.
+Indirect causal links.
+Causal relations in physical systems extend beyond direct interactions.
+A variable $x_{j}$ may affect another variable $x_{k}$ indirectly through an intermediate variable $x_{i}$, forming a pathway $x_{j}\rightarrow x_{i}\rightarrow x_{k}$.
+Such indirect links are central in climate dynamics [126] and in complex systems more broadly [127].
+Within the response framework, these pathways are revealed by the time-dependent operator $R_{k,j}(t)$: the first non-zero response identifies the shortest causal path in the graph connecting $x_{j}$ to $x_{k}$.
+Crucially, if $R_{k,j}(t)=0$ for all time $t$, then no external intervention on $x_{j}$ can ever influence the behavior of $x_{k}$.
+Cumulative causal links or causal sensitivity.
+Finally, Baldovin et al.
+[17] introduced a cumulative (in time) measure of causation between two variables, defined as $D_{k,j}=\int_{0}^{\tau_{max}}R_{k,j}(t)dt$.
+Such cumulative link, can be also referred to as “causal sensitivity”.
+Such a measure is especially relevant in practical experiments with high-dimensional models where interactions unfold across multiple spatial and temporal scales.
+A concrete example using climate models, is provided by the approach of Bloch-Johnson et al.
+(2024) [38], where causal links are inferred between patterns of surface warming and the net radiative flux at the top of the atmosphere.
+Their method reconstructs responses to perturbations by: (i) applying surface warming perturbations, (ii) integrating an atmospheric-only model with and without the perturbation, and (iii) computing the average (in time) difference between the two runs after removing an initial transient.
+In this case, the causal link emerges precisely as a cumulative (in time) atmospheric response to the surface perturbations.
+We emphasize the crucial role of ensemble averages in establishing the general validity of response theory.
+Evaluating responses along single trajectories of nonlinear systems is strongly affected by chaos, restricting the applicability of these ideas to short times and very small perturbations.
+In contrast, focusing on observables, i.e.
+quantities averaged over the invariant density, ensures the robustness of the theory and largely removes the dependence on chaos.
+For a detailed discussion of this point, we refer the reader to [128].
+Future efforts to estimate responses in climate science, which often rely on perturbing single trajectories, would therefore benefit from focusing on ensemble-averaged observables.
+Response operator from data alone.
+Given the assumption of Markovianity it is possible to infer the system’s response operator from a long stationary trajectory using the Fluctuation-Dissipation Theorem (FDT) [17].
+The FDT formula provides a link between stationary and perturbed dynamics by predicting forced responses from stationary statistics alone [129].
+Specifically:
+$R_{k,j}(t)=-\Big\langle A(x_{k}(t))\frac{\partial\ln\rho(\mathbf{x})}{\partial x_{j}}\Big|_{\mathbf{x}(0)}\Big\rangle,$ (15)
+where $A(x_{k}(t))$ is a general observable, $\rho(\mathbf{x})$ is the stationary distribution.
+$\langle\cdot\rangle$ refers to ensemble average, approximated through temporal averages.
+$\frac{\partial\ln\rho(\mathbf{x})}{\partial x_{j}}\Big|_{\mathbf{x}(0)}$ represents the so-called “score function” evaluated at the initial state, i.e.
+on the attractor [15].
+Therefore, in principle the response operator can be computed from data if the score $\nabla\ln\rho(\mathbf{x})$ is available.
+In practice, the score is often inaccessible, and most applications rely on the quasi-Gaussian (qG) approximation, by replacing $\rho(\mathbf{x})$ with a Gaussian distribution $\rho^{G}(\mathbf{x})$ with same mean and covariances [130, 48].
+For responses in the ensemble mean—which are sufficient to infer causal links—the qG approximation achieves high skill even for nonlinear systems [121, 17, 114].
+Recent work [114] has demonstrated that generative modeling techniques can be used to efficiently estimate the score directly from data, leading to high skill in computation of higher-order responses (up to the 5th moment) in strongly nonlinear regimes.
+Other fruitful avenues lie in frameworks at the intersection of Markov State Modeling and Koopman operator theory such as [131].
+It is worth emphasizing that the central limitation discussed in the main text, i.e.
 the need to choose the proper variables, remains relevant for the inference of response operators via FDT as well (and for any causal inference methodology); see the discussion in [17].
-Building on the theoretical insights introduced in the first part, the second part of the paper presents a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net top-of-atmosphere (TOA) radiative flux.
-The model was trained on long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies.
-We showed that the emulator reproduces the stationary statistics of the system and, more importantly, allows us to reconstruct TOA radiative fluxes in forced simulations when provided with forced SST trajectories.
-To deepen this analysis, we applied linear response theory to extract the direct causal links between temperature modes and TOA fluxes, targeting the so-called “pattern effect” [30, 29].
-These links, when projected back onto the high-dimensional state space, were found to be consistent with recent studies.
-This suggests that the emulator’s skill in reconstructing TOA fluxes under forcing arises for the right physical reasons.
-In this way, focusing on a few dominant modes, on a limited range of time scales, and using stochastic parametrizations allowed us to build a skillful and interpretable emulator.
-We emphasize that the fundamental limitations discussed here apply to data-driven models of complex dynamical systems that must both represent stationary statistics and respond faithfully to external perturbations.
-For Earth system models specifically, these considerations primarily concern climate emulators (particularly coupled atmosphere-ocean systems), and should not be conflated with the different challenge of weather forecasting using neural networks (e.g., [128]).
-Weather prediction focuses on short-term predictability where long-term stability and forced responses are of secondary importance.
-Modeling real-world dynamical systems from data alone is severely constrained by high dimensionality and the fact that we typically observe only a subset of relevant variables [46, 32, 40].
-This leads to the fundamental challenge of identifying, when possible, the proper variables to model, a problem recognized since the work of Onsager and Machlup (1953) [68] and highlighted in more recent studies [47, 48].
-These challenges are especially pronounced in climate dynamics, where observations are always partial, and the underlying system is both high-dimensional and multiscale.
-In this context, the multiscale structure of the climate system can be an advantage, allowing us to model coarse-grained dynamics while treating unresolved processes stochastically.
-The need to formulate models tackling the effective dynamics of slow variables, rather than the full set of primitive equations, has a long history in climate science, even in non-data-driven settings where all governing equations are known.
-The early work of Charney and von Neumann, for example, demonstrated the conceptual and practical benefits of filtering out “meteorologically insignificant noise” and emphasized the need for interpretable and computationally efficient effective equations in metereology [129, 63, 43, 44].
-These insights remain conceptually relevant, especially the focus on effective dynamics, in the development of neural emulators for climate, including coupled atmosphere–ocean systems.
-In purely data-driven settings, focusing on stochastic models that target coarse-grained processes offers a promising way forward by reconciling rigorous approaches such as reduced-order modeling with the flexibility and power of neural networks.
-Possible strategies include learning stochastic parameterizations directly from data using deep learning (as in this paper), guiding variable selection through causal inference [130, 131], incorporating information theory and response theory into the training phase [56], studying the evolution of the climate system in its “latent” dynamics [65] (i.e.
-on the attractor) and incorporating empirical knowledge and physics constraints in the learning phase [11, 7, 132, 25].
-Finally, the rather hard difficulty (at times irreducible) of data-driven modeling for the coupled atmosphere–ocean system can at times be ameliorated by building models tailored to solve specific problems, by building on prior literature and insights, and developed with a clear understanding of the relevant variables and the spatiotemporal scales at play.
-Appendix A Technical and practical details on the model formulations
-A.1 Multivariate data-driven stochastic models
+Appendix B Relevance of the the triad system for geophysical turbulence
+Here, we motivate the choice of the triad model analyzed in Section III as a building block for geophysical turbulent flows.
+We further present and test the associated analytical stochastic model for the slow variable $x_{1}$ (as discussed in Section III.3.1), illustrating that Markovian reduced-order models can sometimes be derived from partial observations.
+For detailed mathematical derivations and additional examples, we refer to the original work by Majda and collaborators [46, 60, 61, 48, 16].
+B.1 Link to geophysical turbulence
+Projecting many geophysical turbulent flows, such as quasigeostrophic models, onto a complete set of orthogonal basis functions (spectral or empirical modes such as EOFs), give rise to the following generic form for the mode amplitudes:
+$d\mathbf{u}/dt=\mathbf{F}+\mathbf{L}\mathbf{u}+\mathbf{B}(\mathbf{u},\mathbf{u}),$ (16)
+where $\mathbf{F}$ is an external forcing, $\mathbf{L}$ is a linear operator and $\mathbf{B}(\mathbf{u},\mathbf{u})$ is a nonlinear, quadratic operator.
+The linear operator $\mathbf{L}=\mathbf{L}^{\prime}+\mathbf{D}$, is formed by a skew-symmetric operator $\mathbf{L}^{\prime}$ and by a diagonal dissipation matrix $\mathbf{D}$ with negative real entries.
+The quadratic nonlinearity should be energy-conserving, satisfying $\mathbf{u}\cdot\mathbf{B}(\mathbf{u},\mathbf{u})=0$ [132, 60, 48].
+Eq.
+(16) is the typical abstract form of dry cores of atmospheric models [67].
+Galerkin truncations of turbulent systems often reduce to this structure, as in the Lorenz ’63 and Charney–DeVore models [132, 133, 68].
+$\mathbf{u}\in\mathbb{R}^{N}$ represents the full turbulent geophysical system, with $N$ infinite in principle and finite but very large in applications ($N\gg 1$).
+The state $\mathbf{u}$ is then decomposed into slow and fast components, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, i.e.
+$\mathbf{x}$ will represent the first $n$ components of $\mathbf{u}$ and $\mathbf{y}$ the next NnN-n.
+We consider the simple case of $N=3$: one large-scale resolved mode $x_{1}$ and two fast modes $\mathbf{y}=(x_{2},x_{3})$.
+Majda and collaborators [46] considered an exact projection of the slow mode $x_{1}$ and the two fast modes $(x_{2},x_{3})$ onto the dynamics in Eq.
+(16).
+This gives rise to three equations, defined by a forcing term, linear and quadratic self- and cross-interaction across slow and fast time scales.
+The triad model in Eq.
+(6) is then obtained by approximating the quadratic self-interactions of the fast modes as a damping and stochastic forcing, as: $\gamma_{i}\epsilon^{-1}x_{i}+\sigma_{i}\epsilon^{-1/2}\xi_{i}$, with $i=2,3$.
+$\xi_{i}$ is a standard Gaussian white noise process.
+$\epsilon$ sets the time-scale separation between the slow and fast modes.
+The key assumption is that the detailed nonlinear self-interactions among fast modes are not important if the goal is to capture the coarse-grained slow dynamics.
+This approximation transforms the original deterministic equations into a stochastic system.
+The specific version used in Eq.
+(6) is further simplified by setting to zero a few interactions, such as the linear cross- and self-interactions of the fast modes as well as the external forcing terms.
+Therefore, the relevance of the triad system (Eq.
+(6)) to this study is that it provides a minimal multiscale setting with linear and quadratic nonlinear interaction of the kind that naturally arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs.
+For additional details see also Section 4 of [16].
+B.2 Analytical stochastic model for
+the slow variable $x_{1}$
+Starting from the triad model in Eq.
+(6), Majda and collaborators [60] showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ using standard projection techniques [48, 46, 47].
+The mode reduction predicts the functional form of the dynamics governing the resolved, slow variable [60, 61, 62, 63].
+The resulting stochastic equation is
+$dx=(\tilde{F}+ax+bx^{2}-cx^{3})\,dt+(A-Bx)\,dW_{2}+\sigma\,dW_{3},$ (17)
+where the coefficients are expressed in terms of parameters of the triad system in Eq.
+(6) as: $a=L_{11}+\epsilon(\frac{I^{2}\sigma_{2}^{2}}{2\gamma^{2}_{2}}-\frac{L_{12}^{2}}{\gamma_{2}}-\frac{L_{13}^{2}}{\gamma_{3}})$; $b=-2\epsilon\frac{L_{12}I}{\gamma_{2}}$; $c=\epsilon\frac{I^{2}}{\gamma_{2}}$; $B=-I\epsilon^{1/2}\frac{\sigma_{2}}{\gamma_{2}}$; $L=-B\frac{L_{12}}{I}$; $\tilde{F}=-\frac{AB}{2}$; $\sigma=\epsilon^{1/2}L_{13}\frac{\sigma_{3}}{\gamma_{3}}$.
+$W_{2}$ and $W_{3}$ are independent Wiener processes.
+Interestingly, the equation for the coarse-grained variables $x_{1}$ display a structure richer than that of the full triad system in Eq.
+(6).
+In addition to the original linear and quadratic interactions, the reduced equation includes new terms: an effective forcing $\tilde{F}$, a cubic nonlinearity, and both additive and multiplicative noise.
+The additive noise results from the linear interactions across fast modes, while the multiplicative noise reflects the advection of the slow mode by the fast modes [67].
+Thus, Eq.
+(17) demonstrates that multiplicative noise arises from theoretical considerations, rather than being introduced ad hoc, providing a firm basis for its use in reduced-order stochastic climate models.
+The scalar model above has served as a prototype in several contexts: it has been used to test the Fluctuation–Dissipation Theorem in climate dynamics [81, 114] as well as for inverse modeling of the leading principal component of a quasi-geostrophic model and the observed variability associated with the North Atlantic Oscillation (NAO), demonstrating its relevance for theoretical studies in climate dynamics.
+Figure 9: In each panel: $x_{1}$ Triad” represent statistics of $x_{1}$ in the original triad model in Eq.
+(6); $x_{1}$ Scalar” represent statistics of $x_{1}$ in the scalar reduced model in Eq.
+(17).
+Panel (a): stationary probability distribution.
+Panel (b): response in ensemble mean.
+Panel (c): response in ensemble variance.
+Impulse response functions are computed using an ensemble of $N_{e}=10^{7}$ members.
+We evaluate the model in Eq.
+(17) as done in Section III, so to reproduce the stationary statistics of the slow variable $x_{1}$ as well as its response to perturbations as in the original triad system in Eq.
+(6).
+The results are reported in Figure 9.
+The analysis show that, given $\epsilon\ll 1$, it is possible to build a Markovian model of the slow variable $x_{1}$ reproducing both stationary and perturbed statistics as in the original system.
+Finally, as shown in this Section, rigorous model reduction strategies such as those developed by Majda and collaborators [61] and by Roberts [134, 135] (among others [2]) offer invaluable theoretical guidelines for multiscale climate modeling.
+While such strategies where traiditionally developed in the case when the equation of motions are known, they can possibly serve as inspiration and leading to new directions in future work involving data-driven emulators.
+Finally, as shown in this Section, rigorous model reduction frameworks, such as those developed by Majda and collaborators [61] and by Roberts [134, 135], among others [2], provide invaluable theoretical guidance for multiscale climate modeling.
+While these strategies were traditionally developed for systems with known equations of motion, they may also inspire future work aimed at incorporating such principles into data-driven-only emulators.
+Appendix C Technical and practical details on the model formulations
+C.1 Multivariate data-driven stochastic models
 Consider a $N$-dimensional dynamical system described by a stochastic differential equation $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W}$.
 This equation can be discretized by the forward Euler method as follows:
-$\begin{split}\Delta\mathbf{x}=\mathbf{x}(t+\Delta t)-\mathbf{x}(t)=\mathbf{f}(\mathbf{x}(t))\Delta t+\mathbf{\Sigma}(\mathbf{x})\sqrt{\Delta t}\bm{\xi}(t)~{}\end{split}$ (14)
+[…]
 where $\bm{\xi}(t)$ is a Gaussian white noise with zero mean, unit variance, and $\sqrt{\Delta t}\bm{\xi}(t)$ is the increment of the Wiener process $\Delta\mathbf{W}=\mathbf{W}(t+\Delta t)-\mathbf{W}(t)$.
@@
 The strategy is then to fit a rescaled version of Eq.
-14 from observational data alone.
+(18) from observational data alone.
 Such rescaled equation takes the following form:
@@
 The full fitted equation solved by the emulator can then be written as:
-      $\begin{split}\mathbf{x}(t+1)=\mathbf{x}(t)+\mathbf{f}(\mathbf{x}(t))+\mathbf{\Sigma}\bm{\xi}(t)~{}\end{split}$ (16)
+[…]
       and we remind the reader that $\bm{\xi}(t)\in\mathbb{R}^{N}$ is a Gaussian white noise with zero mean, unit variance.
@@
 The final fitted equation solved by the emulator is:
-      $\begin{split}\mathbf{x}(t+1)=\mathbf{x}(t)+\mathbf{f}(\mathbf{x}(t))+\mathbf{\Sigma}(\mathbf{x}(t))\bm{\xi}(t)~{}\end{split}$ (17)
+[…]
       where again $\bm{\xi}(t)\in\mathbb{R}^{N}$ is a Gaussian white noise with zero mean, unit variance.
 Furthermore, even in the real-world case of sea surface temperature fields and radiative fluxes, reducing the dimensionality to a few, $N=21$, components led to a simplified computational problem where all neural networks could be trained on CPUs and on a personal computer.
-A.2 Scalar data-driven stochastic models
+C.2 Scalar data-driven stochastic models
 Consider a trajectory of length $T$ of a $1$-dimensional dynamical system.
@@
 The full fitted equation solved by the emulator can then be written as:
-    $\begin{split}x(t+1)=x(t)+f(x(t))+\sigma\xi(t)~{}\end{split}$ (19)
+[…]
     and we remind the reader that $\xi(t)$ is a scalar Gaussian white noise with zero mean, unit variance.
@@
 I am especially grateful to Pasha Perezhogin, Rory Basinski, Chris Pedersen, and Andre Souza for their thoughtful feedback and for many enriching discussions on data-driven modeling of the Climate system.
+I acknowledge an interesting email exchange with Prof.
+Tony Roberts, who drew my attention to recent advances on stochastic model reduction.
 This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.

### v4

@@
-Probing forced responses and causal mechanisms in large-scale climate dynamics with reduced-order neural models
+Probing forced responses and causality in data-driven climate emulators: conceptual limitations and the role of reduced-order models
 Abstract
-A central challenge in climate science and applied mathematics is developing data-driven models of multiscale systems that capture not only stationary statistics but also responses to external perturbations.
-Current neural climate emulators seek to resolve the atmosphere–ocean system in all its complexity but often fail to reproduce forced responses, limiting their use in causal studies such as Green’s function experiments.
-Response theory offers a rigorous framework to explore these limitations, moving beyond stationary statistics, probing causal mechanisms, and guiding model design.
-Using a simplified multiscale system, we argue that the skill of purely data-driven models in reproducing perturbed statistics depends critically on (i) identifying the proper variables to model, including spatial and temporal scales, and (ii) accurate stochastic parameterizations of unresolved processes.
-For studies of low-frequency climate dynamics, these insights motivate reduced-order neural models, tailored to specific processes, as valuable alternatives to general-purpose emulators.
-We illustrate this perspective with a reduced-order neural stochastic model designed to investigate responses of radiative fluxes to surface temperature perturbations.
-The model largely reproduces stationary statistics and enables the study of forced responses in both ensemble mean and variance, shifting the study of climate feedbacks from single trajectories to responses of probability distributions.
-These results underscore reduced-order models, augmented with modern neural networks and evaluated within the framework of response theory, as a practical strategy for causal inference and attribution studies in large-scale climate dynamics.
+A central challenge in climate science and applied mathematics is developing data-driven models of multiscale systems that capture both stationary statistics and responses to external perturbations.
+Current neural climate emulators aim to resolve the atmosphere–ocean system in all its complexity but often fail to reproduce forced responses, limiting their use in causal studies such as Green’s function experiments.
+To investigate the origin of these limitations, we first focus on a simplified dynamical system that retains key features of climate variability.
+We interpret the results through linear response theory, providing a rigorous framework to evaluate neural models beyond stationary statistics and probe causal mechanisms.
+We argue that the ability of multiscale systems’ emulators to reproduce perturbed statistics depends critically on (i) identifying an appropriate coarse-grained representation and (ii) careful parameterizations of unresolved processes.
+For low-frequency climate dynamics, these insights highlight reduced-order models, tailored to specific processes and scales, as valuable alternatives to general-purpose emulators.
+We next consider a real-world application, developing a neural model to investigate the joint variability of the surface temperature field and radiative fluxes.
+The model infers a multiplicative noise process directly from data, largely reproduces the system’s probability distribution, and enables causal studies through forced responses.
+We discuss its limitations and outline directions for future work.
+These results expose fundamental challenges in data-driven modeling of multiscale physical systems and underscore the value of coarse-grained, stochastic approaches, with response theory as a principled framework to guide model design.
 preprint: APS/123-QED
@@
 A key challenge in current applied mathematics is to build data-driven models from partial observations of multiscale physical systems.
-Among the most challenging examples is the climate system: a forced, dissipative, and chaotic system whose dynamics span at least 15 orders of magnitude across spatial and temporal scales [1].
-This problem has been extensively studied over the past three decades, mainly in the realm of reduced-order models [2].
-Seminal contributions include the work of Penland and collaborators [3, 4, 5], the multilevel modeling strategies of Kravtsov and colleagues [6, 7, 8, 9], and the contributions of Majda et al.
-[10, 11].
-Recent proposal include Markov-chain models as in [12, 13, 14] and novel ideas from generative models [15].
-Most prior work has focused on evaluating reduced-order models on equilibrium statistics and forecasting, with limited attention to responses to external perturbations, an exception being the contributions of Majda and collaborators (see e.g.
-[16]).
-Capturing such responses is key for building models that not only mimic the system’s dynamics but also faithfully capture causal relations across degrees of freedom [17, 18].
-Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of complex turbulent systems such as the weather [19, 20, 21, 22, 23] and climate [24, 25, 26, 27, 28, 29, 30, 31, 32, 33] systems.
+One of the most complex examples is the climate system: a forced, dissipative, and chaotic system whose dynamics spans at least 15 orders of magnitude across spatial and temporal scales [GhilLucarini].
+This problem has been extensively studied over the past three decades, mainly in the realm of reduced-order models [LucariniChekroun].
+Seminal contributions include the work of Penland and collaborators [Penland89, Penland95, PENLAND1996534], the strategy of Kravtsov and colleagues [KRAVTSOV, KONDRASHOV201533, STROUNINE2010145, Agarwal], and the contributions of Majda et al.
+[majdaVsGhil, majdaConstraints].
+Recent proposals include Markov-chain and generative approaches [Andre1, Andre2, GIORGINI2024134393, GiorginiScore, CycloLudoAndre], as well as autoencoder-based, latent-space formulations [PierreLatent].
+Most prior work has focused on evaluating reduced-order models on equilibrium statistics and forecasting, with limited attention to responses to external perturbations.
+Notable exceptions include the contributions of Majda and Qi [majdaSIAM, QiMajdaChaos].
+Capturing such responses is key for building models that not only mimic the system’s dynamics but also faithfully capture causal relations across degrees of freedom [Baldovin, FabCausal].
+Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of complex turbulent systems such as the weather [Pathak2022, gencast, Lam2023, Bi2023, fourcastnet3] and climate [Kochkov2023, Watt-Meyer2023, Watt-Meyer2024, Lucie, OceanNet, Bire, Samudra, Camulator, lucie3D, SamudraACE].
 Unlike traditional reduced-order models, these neural architectures aim to emulate the full hierarchy of spatial and temporal scales present in the system.
-For example, [30] and [31] emulate the full ocean and atmosphere components of state-of-the-art climate models.
-Such emulators often reproduce stationary statistics and generate forecasts across a range of horizons, offering a promising direction for efficient climate modeling.
-However, several limitations remain.
-Long rollouts suffer from error accumulation that can drive instabilities [34, 35].
-Moreover, current emulators are unable to capture extreme events if not already present in the training set, preventing their use for rare-event analysis in a changing climate [36].
-More relevant here, skill in reproducing stationary statistics does not guarantee accurate responses in perturbation experiments, restricting their use for causal studies of climate dynamics.
+For example, [Samudra] and [Camulator] emulate the full ocean and atmosphere components of state-of-the-art climate models.
+Such emulators often reproduce stationary statistics and generate forecasts across a range of horizons.
+They can potentially be used for tuning and sensitivity analysis [M2Lines], and to improve the study of rare events [Lancelin], offering a promising direction for fast and efficient tools to study climate.
+However, several limitations remain: for example, long rollouts can suffer from error accumulation that can lead to instabilities [Chris, PedramStability]; rare events are not captured if absent from the training set, limiting their applicability for rare-event analysis in a changing climate [PedramPNAS]; key atmospheric modes, such as the Quasi-Biennial Oscillation and the Southern Annular Mode, are not well represented [Baxter].
+A limitation particularly relevant for climate science, as opposed to weather forecasting, is that high skill in reproducing stationary statistics does not ensure accurate responses in perturbation experiments, thereby limiting the use of such models for causal studies of climate dynamics.
 This has two consequences: (i) neural emulators often fail to generalize in out-of-distribution climate change scenarios, and (ii) they may not capture the correct causal relations across variables, two issues that are likely related.
 The work of Van Loon et al.
-[37] illustrates this limitation in the context of the “pattern effect” problem [38, 39], i.e.
-the influence of sea surface temperature (SST) warming patterns on Earth’s energy imbalance.
-Using the ACE2 emulator [26] trained on ERA5 [40], they constructed a Green’s function by perturbing SSTs regionally and measuring the radiative response.
+[Senne] illustrates this limitation in the context of the “pattern effect” problem [blochjohnsonetal2024, Dong2019AttributingPacific], i.e.
+the influence of sea surface temperature (SST) warming patterns on the Earth’s energy imbalance.
+Using the ACE2 emulator [Watt-Meyer2024] trained on the ERA5 reanalysis [era5], they constructed a Green’s function by perturbing SSTs regionally and measuring the radiative response.
 The resulting Green’s function failed to capture the causal relation between SST and radiative fluxes in a quantitative way.
-This is just one example of limitations of neural emulators in capturing responses to perturbations with additional examples in [30, 31, 41].
-Even if such issues are addressed in specific cases, see [42] in the context of the pattern effect, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent.
-Moreover, most current emulators represent only a subsystem (e.g., atmosphere or ocean), and limitations in responding to perturbations are expected to worsen in coupled models.
-Overall, limitations in capturing forced responses suggest that neural emulators are not yet ready for causal applications in climate dynamics.
-While the field of neural emulation is advancing rapidly, it is useful to step back and examine its fundamental limitations in capturing forced responses, both from a theoretical standpoint and in controlled settings.
-In the first part of this paper, we build on a series of previous works by A.
-Vulpiani and collaborators on the limitations of data-driven modeling [43, 44, 45].
-We consider a simplified stochastic dynamical system and ask the following question: Can neural autoregressive models, trained on (practically) infinitely long time series, accurately capture the response to general external perturbations?
+This is just one example of limitations of neural emulators in capturing responses to perturbations with additional examples in [Samudra, Camulator, ZhangACE].
+Even if such issues are addressed in specific cases, as shown by Wu et al (2025) [Wu] in the context of the pattern effect, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent.
+Most current climate emulators represent only a subsystem (e.g.
+atmosphere or ocean) and limitations in capturing forced responses are likely to become more pronounced in coupled settings.
+Additionally, emulators’ skill in capturing forced responses is often assessed using single trajectories.
+In climate applications, however, the relevant quantities are probability distributions, particularly the responses of the mean and variance.
+We will show that focusing on the skill in capturing variance responses can reveal important limitations otheriwse hidden when examining only the mean response.
+Taken together, these issues indicate that neural emulators may not yet be reliable general-purpose tools for causal studies of climate dynamics.
+While the field of neural emulation is advancing rapidly and yielding impressive results, it is useful to step back and examine its fundamental limitations in capturing forced responses, both from a theoretical standpoint and in controlled settings.
+In the first part of this paper, we build on a series of previous works by Vulpiani and collaborators on the limitations of data-driven modeling [FutureFromPast, Hosni, BaldovinEntropy].
+We consider a simplified stochastic dynamical system and ask the following question: Can neural autoregressive models, trained on (practically) infinitely long time series, accurately capture responses to general external perturbations?
 In addressing this question, we focus on two relevant scenarios in data-driven-only modeling:
@@
 Data-driven modeling in the presence of partial observations.
-This reflects common constraints in modeling high-dimensional systems from data, where the full set of relevant variables is typically unknown.
+This reflects common constraints in modeling high-dimensional systems from data, where the full set of relevant variables is typically unobserved [e20070509], or even unknown [BaldovinEntropy].
 In this more realistic setting, we again assume no knowledge of the underlying equations and only partial access to the system’s state vector.
-To tackle such questions we proceed pedagogically by focusing on a multiscale “triad model” first proposed by Majda and collaborators in [46, 47].
-The model mimics the abstract structure of dynamical cores in large-scale climate models with a linear operator and an energy conserving nonlinear, quadratic operator.
-It provides a simplified yet relevant mathematical framework to unambiguously explore key limitations in capturing forced responses in climate emulators.
+To tackle such questions we proceed pedagogically by focusing on a multiscale “triad model” first proposed by Majda and collaborators in [NormalForms, Majda2010].
+The model mimics the abstract structure of dynamical cores of large-scale climate models with a linear operator and energy conserving quadratic nonlinearities.
+It provides a simplified yet relevant mathematical framework to unambiguously explore fundamental (rather than technical) limitations in capturing forced responses in climate emulators.
 We address the questions above by training neural stochastic emulators on a long integration of the triad model and test the models’ performance on stationary and perturbed cases.
 In doing so, we face the following issue: quantifying how an emulator respond to general forcing is an ill-posed question, as there are infinite possible forcings.
-Nonetheless, the linear response to a general forcing can be computed by convolving the forcing with an impulse response operator [48, 49].
-This operator can be inferred from the model and is a building block that an emulator must capture to reproduce the response to small, general external perturbations [48], as well as the cause–effect relations among variables [17].
+Nonetheless, the linear response can be computed by convolving the forcing with an impulse response operator [MajdaBook, Marconi].
+This operator can be inferred from the model and is a building block that an emulator must capture to reproduce the response to general external perturbations.
+Therefore, linear response theory is here proposed as a rigorous framework to evaluate neural models beyond stationary statistics and probe causal mechanisms [Baldovin].
 In the case of a fully observed state vector we show that the inferred neural model can reproduce stationary and perturbed statistics of the system.
 In the case of partial observations, as in realistic settings, the performance of a data-driven model necessarily depends on (i) the choice of proper variables to model and (ii) suitable parameterizations of unobserved processes.
-The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [5, 50, 51], though typically without addressing responses to external perturbations.
+The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [PENLAND1996534, Gammaitoni, Dalmedico], though typically without addressing responses to external perturbations.
 Therefore, we will argue that, in the case of low-frequency climate dynamics, a promising perspective is to build on such established approaches, emphasizing coarse-grained dynamics and stochastic parameterizations, while using neural networks as powerful tools within this framework.
 We emphasize that the focus on a simplified dynamical system is not a limitation in this context.
-In fact, the goal is not to demonstrate the success of a method, but to illustrate, by analogy, fundamental issues of data-driven modeling that are already evident in low dimensions and are likely to become more pronounced in higher-dimensional settings.
-As Smith (2002) [52] argues, while it is unreasonable to expect solutions to low-dimensional problems to generalize to high dimensions, so too it is unlikely that limitations observed in low-dimensional systems will vanish in high-dimensional, operational models.
+In fact, the goal is not to demonstrate the success of a method, but to illustrate, by analogy, fundamental issues of data-driven modeling that are already evident in low dimensions and are likely to become more pronounced in higher-dimensional settings [Lenny2002].
 The first part of this paper is pedagogical in nature, aiming to bridge perspectives across researchers in statistical physics, climate science, and machine learning, while also providing practical methodological tools for the design and evaluation of data-driven models.
@@
 By real-world examples we refer to situations where (i) we observe only the temporal evolution of a subset of variables without access to the full state vector (i.e.
-a projection of the full dynamics) or (ii) where even in the presence of all variables the system encompasses a huge number of spatial and temporal scales.
-Both cases (i) and (ii) lead to the need to formulate effective rather than primitive equations for both practical and conceptual reasons [50].
-For an early instance of case (ii), see the work of Charney (1948) [53], Charney et al.
-(1950) [54] and the discussion in Dalmedico (2001) [51].
+a projection of the full dynamics) or (ii) when even in the presence of all variables the system encompasses a huge number of spatial and temporal scales.
+Both cases (i) and (ii) lead to the need to formulate effective rather than primitive equations for both practical and conceptual reasons [Gammaitoni].
+For an early instance of case (ii), see the work of Charney (1948) [charney1949], Charney et al.
+(1950) [neumann] and the discussion in Dalmedico (2001) [Dalmedico].
 The example we consider is the “pattern effect” introduced above, which requires Green’s function experiments to quantify the response of top-of-atmosphere (TOA) radiative fluxes to perturbations in the sea surface temperature (SST) field.
 To this end, we develop a simplified stochastic neural model that emphasizes a large-scale representation of SST and global-mean net radiative flux, rather than attempting to resolve all spatial and temporal scales of variability as in current emulators.
-We discuss the coarse-graining and preprocessing choices relevant for model construction.
-The model largely reproduces both stationary and forced statistics and, importantly, enables Green’s function–like experiments to infer direct causal relations between SST and TOA radiative flux, a task where general-purpose emulators can perform poorly [37].
-Crucially, we focus on responses in the ensemble mean and variance, rather than single trajectories as in most current studies, thereby extending the concept of climate feedback to include changes in distributional properties.
+We discuss the coarse-graining choices relevant for model construction, as well as a practical and efficient way to fit multiplicative noise from data.
+The model largely reproduces stationary statistics, especially the most energetic SST mode, corresponding to the El Niño Southern Oscillation [ENSOcomplexity], and it enables Green’s function–like experiments to infer direct causal relations between SST and TOA radiative flux.
+The discussion also addresses the model’s limitations and potential methodological developments.
 The paper is organized as follows: in Section II we introduce the theoretical ideas underpinning this work.
-In Section III we discuss key conceptual difficulties in data-driven modeling of real-world dynamical systems.
-In Section IV we present a real-world application by focusing on radiative feedbacks to temperature perturbations.
-Conclusions then follow in Section V.
+In Section III we discuss key conceptual limitations in data-driven modeling of real-world dynamical systems.
+In Section IV we present a real-world application.
+A summary of the main findings and proposals follow in Section V and Section VI.
 II Theoretical Framework: Data-driven stochastic climate modeling and response theory
@@
 $d\mathbf{u}/dt=\mathbf{F}+\mathbf{L}\mathbf{u}+\mathbf{n}(\mathbf{u}),$ (1)
-where $\mathbf{F}$, $\mathbf{L}$ and $\mathbf{n}(\mathbf{u})$ represent an external forcing, a linear and nonlinear operators, respectively.
-$\mathbf{u}$ is not necessarily finite dimensional, and in the context of coupled atmosphere-ocean variability, it could in principle represent the deviation of all fields, at all scales, from their climatology [4].
+where $\mathbf{F}$ is an external forcing, and $\mathbf{L}$, $\mathbf{n}(\mathbf{u})$ represent linear and nonlinear operators, respectively.
+$\mathbf{u}$ is not necessarily finite dimensional, and in the context of coupled atmosphere-ocean variability, it could in principle represent the deviation of all fields, at all scales, from their climatology [Penland95].
 However, for obvious practical reasons, Eq.
 (1) needs to be simulated on finite spatial and temporal resolutions.
-In contrast to numerical weather prediction, which demands high spatial and temporal resolution, low-frequency climate variability is governed by a few dominant large-scale modes [4, 7, 48, 47, 55, 18].
-The goal of reduced-order stochastic climate models is to develop low-dimensional models simulating large-scale, coarse-grained climate dynamics.
+In contrast to numerical weather prediction, which demands high spatial and temporal resolution, low-frequency climate variability is governed by a few dominant large-scale modes [Penland95, KONDRASHOV201533, MajdaBook, Majda2010, Dubrulle, FabCausal].
+The goal of reduced-order stochastic climate models is to capture the evolution of these modes and their interactions.
+This modeling choice allows us to move away from a grid-based representation of the system, emphasizing resolution-independent modes as fundamental components of the system.
 The cumulative effect of unresolved fast variables, present at smaller spatial scales, is parametrized as stochastic terms.
@@
 In the latter case, the equation is here interpreted in the Itô sense.
-Equation (2) is mathematically justified in the limit of infinite time-scale separation between fast and slow modes [1].
+Equation (2) is mathematically justified in the limit of infinite time-scale separation between fast and slow modes [GhilLucarini].
 In the most general (and often intractable) setting, the effective dynamics of the slow variables also involve memory effects induced by the fast degrees of freedom, leading to non-Markovian dynamics.
 Kondrashov et al.
-(2015) [7] addressed this problem in a practical and efficient way by introducing the framework of multilevel stochastic models.
-Nonetheless, appropriate coarse-graining procedures have been shown to effectively average out non-Markovian contributions in many climate applications, yielding useful Markovian approximations [3, 4].
-We refer the reader to [2] for a review on the role of non-Markovian contributions in effective climate equations and to [56, 57] for specific examples.
+(2015) [KONDRASHOV201533] addressed this problem in a practical and efficient way by introducing the framework of multilevel stochastic models.
+Nonetheless, appropriate coarse-graining procedures have been shown to effectively average out non-Markovian contributions in many climate applications, yielding useful Markovian approximations [Penland89, Penland95].
+We refer the reader to [LucariniChekroun] for a review on the role of non-Markovian contributions in effective climate equations and to [Wouters, Vissio2018] for specific examples.
 Reduced-order stochastic climate models, as in Eq.
@@
 They are much more computationally efficient and allow us for mechanistic understanding of climate dynamics by isolating core processes and examining their qualitative behavior.
-This has fueled the development of low-dimensional stochastic models learned either from data [58, 3, 4, 16, 10, 7, 2, 59] or from known equations [60, 61, 62, 63, 46], as well as inference-based techniques to infer interactions across modes of variability [64, 65, 66, 18].
+This helps bridging the gap between simulations and understanding [HeldGap].
+This has fueled the development of low-dimensional stochastic models learned either from data [Hasselmann, Penland89, Penland95, majdaSIAM, majdaVsGhil, KONDRASHOV201533, LucariniChekroun, nanChen] or from known equations [MTV1, MTV2, MTV3, MTV4, NormalForms], as well as inference-based techniques to infer interactions across modes of variability [Kane, Donges, Ilias_1, FabCausal].
 In practice, given a phenomenon of interest, we typically begin with spatiotemporal data consisting of $N$ time series of length $T$. We then often proceed as follows:
@@
 temperature, velocities etc.), (ii) retaining a few relevant modes or large-scale averages, (iii) restricting the range of temporal scales.
-Throughout the paper, we will often refer to the so-called proper variables as the coarse-grained spatiotemporal fields obtained via (i–iii).
-The relevant modes are often identified through Empirical Orthogonal Functions (EOFs) [67, 8], by retaining the first $n<N$ of such bases.
+Throughout the paper, we will often refer to the so-called proper variables as the coarse-grained representation obtained via (i–iii).
+The relevant modes are often identified through Empirical Orthogonal Functions (EOFs) [MajdaPerspective, STROUNINE2010145], by retaining the first $n<N$ of such bases.
 The correspondent principal components (PCs), serve as the reduced time series for model training.
-Although EOFs are convenient, they may not be always optimal, and alternatives such as in [68, 18, 69] may be more suitable depending on the application.
-The performance of the final model is sensitive to all coarse-graining choices (i–iii); see, e.g., [70, 71].
-Identifying the “right” coarse-grained representation remains one of the hardest challenges in data-driven modeling, as no general mathematical principle exists to guide such choices in high-dimensional dynamical systems [43, 17, 17].
+Although EOFs are convenient, they may not be always optimal, and alternatives such as principal interaction patterns [CROMMELIN], community detection [FabCausal] and autoencoders [latentPierre, PierreLatent] may be more suitable depending on the application.
+The performance of the final model is sensitive to all coarse-graining choices (i–iii); see, e.g., [HeldFDT, Pedram2].
+Identifying the “right” coarse-grained representation depends on the specific goals of the researcher and remains one of the most difficult challenges in data-driven modeling, as no general mathematical principle exists to guide such choices from data alone in high-dimensional dynamical systems [FutureFromPast, Baldovin, Baldovin].
 Inferring a stochastic model.
@@
 (2).
-We summarize several traditional approaches that motivate our work, while acknowledging that many other methods exist beyond those discussed here.
-The simplest example is the Linear Inverse Model (LIM) proposed by Penland [3], where the drift is linear and the noise is additive, both estimated directly from data.
-LIMs have shown forecasting skill for sea surface temperature anomalies comparable to climate models [4, 5, 72, 73], and extensions exist to handle colored noise [74].
+We summarize several traditional approaches that motivated our work, while acknowledging that many other methods exist beyond those discussed here.
+The simplest example is the Linear Inverse Model (LIM) proposed by Penland [Penland89], where the drift is linear and the noise is additive, both estimated directly from data.
+LIMs have shown forecasting skill for sea surface temperature anomalies comparable to climate models [Penland95, PENLAND1996534, Penland_Fowler, Capotondi].
+Recent developments extended the LIM framework to include colored noise [coloredLIM] and nonlinearities [LIM_nonlinear].
 Kravtsov et al.
 (2005) and Kondrashov et al.
-(2015) [6, 7] generalized this framework through multilevel regression, incorporating quadratic drift terms and non-Markovian effects.
-These models perform well for stationary statistics [8] and forecasting [75].
-The work of Majda and Harlim extended the multilevel regression framework by proposing stability constraints [11].
-A theoretically grounded alternative is the approach of Majda, Timofeyev, and Vanden-Eijnden [60, 61, 62, 63], dubbed MTV by the authors, which derives effective equations for the coarse-grained variables $\mathbf{x}$ rather than assuming them.
-MTV has been successful for simplified models, though less so for complex climate systems, likely due to choices of coarse-graining steps (i-iii) shown above rather than intrinsic limitations, see also discussion in [8].
+(2015) [KRAVTSOV, KONDRASHOV201533] generalized this framework through multilevel regression, incorporating quadratic drift terms and non-Markovian effects.
+These models perform well for stationary statistics [STROUNINE2010145] and forecasting [KONDRAENSO].
+The work of Majda and Harlim extended the multilevel regression framework by proposing stability constraints [majdaConstraints].
+A theoretically grounded alternative is the approach of Majda, Timofeyev, and Vanden-Eijnden [MTV1, MTV2, MTV3, MTV4], dubbed MTV by the authors, which derives effective equations for the coarse-grained variables $\mathbf{x}$ rather than assuming them.
+MTV has been successful for simplified models, though less so for complex climate systems, likely due to choices of coarse-graining steps (i-iii) shown above rather than intrinsic limitations, see also discussion in [STROUNINE2010145].
 We will expand on this approach in Section III and Appendix B.
@@
 In this paper, we build on these insights and investigate how simple neural architectures can infer both the drift and state-dependent noise in Eq.
-(2), with minimal assumptions.
+(2) with minimal assumptions.
 Our neural models retain a decomposition into linear and nonlinear terms, $\mathbf{L}$ and $\mathbf{n}(\mathbf{x})$, as in the general representation of dynamical systems in Eq.
 (1).
-Furthermore, we will propose simple solutions to fit multiplicative (state-dependent) noise in high dimensions.
 II.2 Response theory: a (very) brief overview
@@
 Impulse response operator.
-Given a $n$-dimensional coarse-grained state vector $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$, we perturb a degree of freedom $x_{j}$ with a small, step function perturbation $x_{j}(0)\rightarrow x_{j}(0)+\delta x_{j}(t)$, with $\delta x_{j}(t)=\Delta_{j}\delta(t)$, where $\delta(t)$ is a delta function and $\Delta_{j}$ is a small perturbation amplitude [76, 17].
+Given a $n$-dimensional coarse-grained state vector $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$, we perturb a degree of freedom $x_{j}$ with a small, step function perturbation $x_{j}(0)\rightarrow x_{j}(0)+\delta x_{j}(t)$, with $\delta x_{j}(t)=\Delta_{j}\delta(t)$, where $\delta(t)$ is a delta function and $\Delta_{j}$ is a small perturbation amplitude [Risken, Baldovin].
 We then consider a general observable $A(x_{k}(t))$ (i.e.
@@
 where $\langle\cdot\rangle$ refers to ensemble average and the subscript “p” refers to the perturbed system.
-In this paper we are going to focus on two different specific observables: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution.
+In this paper we are going to focus on two observables: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution.
 The two observables allow us to capture time-dependent responses in the mean and variance of the perturbed distribution respectively.
@@
 $R_{k,j}(t)=\lim_{\delta x_{j}(0)\to 0}\frac{\delta\langle A(x_{k}(t))\rangle}{\delta x_{j}(0)}.$ (4)
-The utility of the inferred impulse response operator is two-fold:
+We refer to Appendix A for details on how to infer $R_{k,j}(t)$ in practice. The utility of the inferred impulse response operator is two-fold:
   • Causal inference.
-Understanding causal relations across degrees of freedoms in physical systems require us to study the effect of interventions on the system [77].
-The inferred response operator quantifies time-dependent cause-effect dependencies across degrees of freedoms $x_{j}$ and $x_{k}$ [78, 79, 17, 80, 18].
+Understanding causal relations across degrees of freedoms in physical systems requires to study the effect of interventions on the system [IsmaelNew].
+The inferred response operator quantifies time-dependent cause-effect interactions across degrees of freedoms $x_{j}$ and $x_{k}$ [Aurell, Lucarini2018, Baldovin, Valerio2024, FabCausal].
 We refer to Appendix A for details on the relation between the response operator and causality.
   • Response to generic time-dependent forcing.
-The response operator $\mathbf{R}(t)$ can be used to predict the response of the system to small, but general, time-dependent external forcing $\mathbf{f}(t)\in\mathbb{R}^{n}$ applied directly to the system’s degrees of freedoms as:
+The response operator $\mathbf{R}(t)$ can be used to predict the response of the system to small, but general, time-dependent external forcing $\mathbf{f}(t)\in\mathbb{R}^{n}$ as:
     $\delta\langle A(x_{k}(t))\rangle=\sum_{j}\int_{0}^{t}R_{k,j}(\tau)f_{j}(t-\tau)d\tau.$ (5)
@@
 A data-driven model capable of accurately capturing the impulse response operator will also capture the time-dependent causal interactions across degrees of freedom.
-Response theory can then be used to assess the skills of emulators by going beyond purely statistical metrics such as r-squared, by for example assessing the causal sensitivity of the system to external perturbation [18].
-Moreover, since the linear response to small perturbations is given by the convolution formula in Eq.
-(5), recovering the impulse response operator ensures a faithful approximation of the system’s linear response to general forcings.
-For this reason, we view the impulse response operator as a fundamental building block that a data-driven model must reproduce to capture the response to small, general perturbations, and is necessary (but not sufficient) to capture nonlinear responses as well.
-Importantly, the response operator can be inferred from data alone through the Fluctuation-Dissipation Theorem [49], this is discussed in Appendix A.
+Response theory can then be used to assess the skills of emulators by going beyond purely statistical metrics such as r-squared, assessing the causal sensitivity of the system to external perturbation [FabCausal].
+Thus, this approach opens new, rigorous avenues for evaluating and studying data-driven models.
+Moreover, recovering the impulse response operator ensures a faithful approximation of the system’s linear response to small general forcings.
+For this reason, we view the impulse response operator as a fundamental building block that a data-driven model must reproduce to capture the responses to general perturbations.
+Importantly, the response operator can be inferred from data alone through the Fluctuation-Dissipation Theorem [Marconi], as discussed in Section V and in Appendix A.
 III Limitations of data-driven modeling by analogy: the case of partial observations
-A fundamental limitation of purely data-driven modeling of multiscale climate systems is that we cannot observe all degrees of freedom.
-Even when governing equations are known, as in meteorology, climate fields are only available at finite temporal and spatial resolution.
-Dynamics at smaller and faster scales are therefore missing from the data and often remain unparameterized in emulators, unlike in full climate models.
-An even greater challenge arises in many areas where we lack governing equations altogether to guide emulator construction.
-In both cases, data-driven approaches can mimic the observed variability of the retained variables, reproducing probability distributions, autocorrelation functions, and providing reasonable forecasts.
-Yet testing responses to perturbations, where unobserved dynamics becomes relevant, often exposes structural deficiencies that are otherwise hidden.
+A fundamental limitation of purely data-driven modeling of real-world, multiscale climate systems is that many degrees of freedom remain unobserved.
+Even when the governing equations are known, as in meteorology, climate fields are only sampled at finite temporal and spatial resolutions, which requires data-driven closures for unresolved fast and small-scale processes.
+The challenge is even greater in settings where governing equations are unknown and provide no guidance for emulator design.
+In both cases, data-driven approaches can reproduce the variability of the retained variables, including their probability distributions, autocorrelation functions, and short-term forecasts.
+However, the response of probability distributions to perturbations, particularly the mean and variance in realistic climate applications, can uncover structural deficiencies stemming from unresolved dynamics that would otherwise remain hidden from stationary statistics alone.
 The goal of this section is to illustrate such limitations within a controlled mathematical setting, where they can be made explicit and unambiguous.
 III.1 A triad model as a simplified prototype for coarse-grained, stochastic climate modeling
-We focus on a simplified yet relevant dynamical system that retains essential features of high-dimensional geophysical flows, making the conclusions of the Section meaningful for climate applications.
-The model was introduced by Majda and collaborators [46] and later used in [47] in the context of climate dynamics.
+We focus on a simplified yet relevant dynamical system that retains essential features of high-dimensional geophysical flows, making the conclusions of this section meaningful for climate applications.
+The model was introduced by Majda and collaborators [NormalForms] and later used in [Majda2010] in the context of climate dynamics.
 In what follows, we present the model and refer to Appendix B for an in-depth discussion on its relevance as a building block for complex turbulent flows.
@@
 The parameter $I$ controls the strength of the nonlinearity: when $I=0$, the model reduces to purely linear dynamics and Gaussian statistics.
-The quadratic interactions, here between $x_{1}$ and $x_{2}$, typically arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs [16].
+The quadratic interactions typically arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs [majdaSIAM].
 The parameter $\epsilon$ sets the time-scale separation between the slow mode $x_{1}$ and the fast modes $(x_{2},x_{3})$, allowing exploration of multiscale dynamics in a controlled setting.
-The terms $W_{2}$ and $W_{3}$ represent independent Wiener processes, introducing stochastic forcing into the second and third equations.
-The model generates unimodal distributions with relatively small, but significant, deviations from Gaussianity, which is a relevant feature in the study of low-frequency climate variability [81] and found in both general circulation models and observational data, see e.g.
-[82, 83, 84].
+The terms $W_{2}$ and $W_{3}$ represent independent Wiener processes.
+The model generates smooth, unimodal distributions with relatively small (but significant) deviations from Gaussianity, and it exhibits fast decay of correlations.
+Both features are relevant in the study of low-frequency climate variability [MajdaStructuralStability] and found in both general circulation models and observational data, see e.g.
+[FranzkeMajda, Judith, Sura].
 III.2 The case of unknown equations but a fully observed state vector
@@
 The linear method is considered for two reasons.
-First, we show that linear inverse modeling (LIM), as originally proposed in [3], serves as a strong baseline for neural autoregressive models, successfully capturing both the mean and variance of the system’s stationary distribution, the autocorrelation functions, and the ensemble mean response to external perturbations.
-Second, and consistent with theoretical expectations [47], we demonstrate that linear models are fundamentally limited in capturing changes in the full probability distribution: for instance, they cannot represent any response in the variance to external forcings.
+First, we show that linear inverse modeling (LIM), as originally proposed in [Penland89], serves as a strong baseline for neural autoregressive models, successfully capturing both the mean and variance of the system’s stationary distribution, the autocorrelation functions, and the ensemble mean response to external perturbations.
+Second, and consistent with theoretical expectations [Majda2010], we demonstrate that linear models are fundamentally limited in capturing changes in the full probability distribution: for instance, they cannot represent any response in the variance to external forcings.
 This limitation restricts their usefulness for studying forced changes in distributions beyond the mean.
 III.2.2 Stationary statistics
-We integrate the two data-driven models inferred from data and specified in Eq.
-(7) and (8) for $T=10^{7}$, yielding time series of the same length as those obtained from the original triad system.
+We integrate the two data-driven models in Eq.
+(7) and Eq.
+(8) for $T=10^{7}$, yielding time series of the same length as those obtained from the original triad system.
 We then compare the modeled and original probability density functions (PDFs) and autocorrelation functions (ACFs), as shown in Figure 1.
-The neural model described in Equation (8) successfully captures the stationary distributions of $x_{1}(t),x_{2}(t)$ and $x_{3}(t)$.
+The neural model described in Eq.
+(8) successfully captures the stationary distribution of the triad system.
 In particular, it outperforms the linear model by accurately reproducing the non-Gaussian features of $x_{1}$.
 Both models also reproduce the autocorrelation functions well, as seen in the second row of Figure 1.
-In summary, the neural emulator defined in Equation (8) is capable of capturing both the PDF and autocorrelations of the full state vector $\mathbf{x}(t)$.
+In summary, the neural emulator is capable of capturing both the PDF and autocorrelations of the full state vector $\mathbf{x}(t)$.
 At the same time, the linear model serves as a strong baseline for evaluating more sophisticated data-driven approaches.
@@
 Impulse response functions.
-We now examine the impulse response of the original triad model (Equation (6)), the linear model (Equation (7)), and the nonlinear neural model (Equation (8)).
+We now examine the impulse response of the original triad model (Eq.
+(6)), the linear model (Eq.
+(7)), and the nonlinear neural model (Eq.
+(8)).
 Specifically, we analyze the time-dependent response of observables $A(x_{k}(t))$, with $k=1,2,3$, to an impulse perturbation imposed on $x_{1}(0)$ at time $t=0$.
-We choose as observables the mean and variance of the time-dependent perturbed distribution and refer to Appendix A for details on how to compute the response operator.
+We choose as observables the mean and variance of the distribution.
 The results, shown in Figure 2, reveal that the nonlinear neural model captures the mean response of the distribution quantitatively, while the linear model captures it only qualitatively.
-The most significant differences arise in the response of the variance: the linear model yields exactly zero variance response, clearly highlighting the limitations of linear dynamics [47].
+The most significant differences arise in the response of the variance: the linear model yields exactly zero variance response, as expected [Majda2010].
 In contrast, the neural model captures the variance response with relatively high fidelity.
@@
 Responses to a specific logarithmic forcing.
-we now consider a more targeted test involving a specific, time-dependent forcing.
-We introduce a logarithmic forcing term to the right-hand side of the original and data-driven triad model.
+we now consider a more targeted test and introduce a logarithmic forcing term to the right-hand side of both the original and data-driven triad model.
 The forcing is applied only to $x_{1}$ and takes the form $(0.01log(1+t),0,0)$ per time step.
@@
 Given a (practically) infinitely long trajectory of a stochastic dynamical system, and assuming the ideal condition where the full state vector is observed, it is possible to construct a neural stochastic model that is stable, reproduces the stationary statistics, and responds accurately to small but general external perturbations.
-The analysis does not rule out that other systems may struggle to capture responses even with full observations; however, these issues are here regarded as technical rather than fundamental.
-Such technical issues can be addressed by considering more complex architectures or by adding constraints withing a chosen architecture.
-While these results are encouraging, they remain far from any real-world scenario.
+The analysis does not rule out that other systems may struggle to capture responses even with full observations.
+These issues, while important, are considered here as technical, rather than fundamental, and point to the need for more sophisticated architectures, improved algorithms, or the inclusion of additional physical constraints within existing frameworks.
+While the results shown in this Section are encouraging, they remain far from any real-world scenario.
 In practice, we typically have access only to partial observations of a complex, interconnected system.
@@
 In this setting the variable $x_{1}$ clearly is the slow mode, forcing the much faster modes $x_{2}$ and $x_{3}$.
-We again integrate the system in (6) with an Euler–Maruyama scheme for a period of length $T=10^{7}$ with time step $dt=0.01$ (i.e.
+We again integrate the system in Eq.
+(6) with an Euler–Maruyama scheme for a period of length $T=10^{7}$ with time step $dt=0.01$ (i.e.
 $10^{5}$ Model Time Units).
@@
   2. ii.
-    It must correctly capture its response to perturbations as in the original triad system.
-Criterion (ii) is particularly demanding because, in the triad model defined by Eq.
-(6), perturbations can propagate through the system, inducing cascades of feedbacks.
-These time-dependent feedbacks must be effectively parametrized within the scalar model, despite the absence of direct observations of some components.
-While clearly nontrivial, this problem is representative of real-world data-driven modeling scenarios, where models are often built from partial observations of high-dimensional systems.
-Proposing a scalar data-driven model from partial observations that satisfies the two criteria above raises two central challenges:
+    It must correctly capture the mean and variance response to perturbations as in the original triad system.
+Both criteria are nontrivial, with the second posing the greater difficulty.
+Together they mirror core challenges in data-driven modeling built from partial observations.
+Satisfying these criteria leads to two central challenges:
   • Choice of variable to model.
@@
 The only viable option is to model the slow mode $x_{1}$, since the cumulative effect of the much faster $x_{2}$ and $x_{3}$ can be represented stochastically.
-Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model under partial observations [48, 43, 85, 86].
+Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model under partial observations [MajdaBook, FutureFromPast, Lacorata, BaldovinPlos].
 In realistic spatiotemporal climate dynamics, this choice translates into selecting appropriate climate fields together with a narrow range of spatial and temporal scales as we will show in Section IV.
@@
 Having identified the proper variable $x_{1}$ to model, the next challenge is to appropriately parametrize the cumulative effects of the unresolved components $x_{2}$ or $x_{3}$ onto the slow mode $x_{1}$.
-This leads to the classic problem of stochastic parametrization [87, 88, 2, 1, 67, 89].
-As we will show, the success of the resulting model, particularly in capturing forced responses, hinges critically on the quality of this parametrization.
+This leads to the classic problem of stochastic parametrization [ArnoldHM, Dawson, LucariniChekroun, GhilLucarini, MajdaPerspective, StochBAMS].
+As we will show, the success of the resulting model, particularly in capturing forced responses in the variance, hinges critically on the quality of this parametrization.
 We note that for the specific case of the triad model in Eq.
-(6), Majda and collaborators showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ [46, 47]; see discussion in Appendix B.
+(6), Majda and collaborators showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ [NormalForms, Majda2010]; see discussion in Appendix B.
 Our aim here is not to test such analytical strategies, which require full knowledge of the governing equations and are often limited in real-world settings; see footnote  [^1] Even moderately more complex systems already expose the practical limitations of rigorous stochastic mode reduction procedures; see Strounine et al.
-(2008) in the context of quasi-geostrophic modeling [8]..
+(2008) in the context of quasi-geostrophic modeling [STROUNINE2010145]..
 Rather, we use the triad model as a simple conceptual tool to highlight key challenges of data-driven modeling of multiscale systems.
-In realistic settings, both the governing equations and the full set of relevant variables are typically unknown [91].
-Even before specifying a model structure, identifying the slow variables is nontrivial and highly problem-dependent.
+In realistic settings, both the governing equations and the full set of relevant variables are typically unknown [Cecconi]: even before specifying a model structure, identifying the slow variables is nontrivial and highly problem-dependent.
 In what follows, we assume that the proper slow variable, $x_{1}$, has been correctly identified.
@@
     $L$ is a linear operator (here just a scalar), $n(x)$ is a nonlinear operator and $\sigma^{2}$ is the noise variance. $W$ is the standard Wiener process.
-  • Neural model with multiplicative noise. We consider a neural model as before, but now defined with multiplicative rather than additive noise:
+  • Neural model with multiplicative noise. We consider a neural model with multiplicative:
     $dx=(Lx+n(x))dt+\sigma(x)dW.$ (10)
+    We use the same drift $(Lx+n(x))$ fitted for the additive model in Eq.
+(9), so that differences can only arise from the stochastic parametrization.
 The multiplicative (state dependent) noise $\sigma(x)$ is fitted with an additional neural network, see Appendix C.
@@
 In contrast, incorporating multiplicative noise (Eq.
-(10)) leads to a substantially improved representation of the stationary distribution, see Figure 4(e).
-The mean response remains largely unchanged, decaying too slowly, but the model now captures a qualitatively correct signal in the ensemble variance, albeit with a similarly slow decay.
+(10)) leads to a substantially improved representation of the stationary distribution (Figure 4(d)) and the model now captures a qualitatively correct signal in the ensemble variance, albeit with a similarly slow decay.
 These results indicate that the inclusion of multiplicative noise improves the reduced-order model’s ability to represent both stationary statistics and variance responses to perturbations.
-In the full triad system, perturbing $x_{1}$ at time $t=0$ leads to changes in $x_{1}$, $x_{2}$ and $x_{3}$ at the next time step $dt$.
-Consequently, $x_{2}$ and $x_{3}$ feedback on $x_{1}$ at the consequent time step.
-Such time-dependent feedbacks require the inclusion of state dependent noise $\sigma(x)$ to effectively parametrize such interactions.
-Moreover, as discussed in Appendix B, multiplicative noise is not an ad-hoc addition but is theoretically justified and explicitly predicted by rigorous mode reduction strategies [60].
-However, even when multiplicative noise is considered, the responses to impulse perturbations are still only qualitatively right.
+Moreover, as discussed in Appendix B, the use of multiplicative noise in reduced-order modeling is not an ad-hoc addition but is theoretically justified and explicitly predicted by rigorous mode reduction strategies [MTV1].
 We emphasize two key points:
@@
 The variability of the unobserved variables $x_{2}$ and $x_{3}$ manifest as memory effects in the reduced order scalar model for $x_{1}$.
-One strategy for addressing such memory effects is the multilevel regression approach first introduced by Kratsov, Kondrashov and Ghil (2005) [6] and later generalized by Kondrashov, Checkroun and Ghil (2015) [7].
-The multilevel regression strategy augments the reduced-order model with a hierarchy of linear equations solving for the residuals from the deterministic dynamics and it can be thought as a practical implementation of the Mori–Zwanzig formalism in statistical mechanics [7, 2].
+One strategy for addressing such memory effects is the multilevel regression approach first introduced by Kratsov, Kondrashov and Ghil (2005) [KRAVTSOV] and later generalized by Kondrashov, Checkroun and Ghil (2015) [KONDRASHOV201533].
+The multilevel regression strategy augments the reduced-order model with a hierarchy of linear equations solving for the residuals from the deterministic dynamics and it can be thought as a practical implementation of the Mori–Zwanzig formalism in statistical mechanics [KONDRASHOV201533, LucariniChekroun].
 An alternative, pragmatic approach for handling non-Markovian effects is to coarse-grain the data in the temporal dimension, for instance, by averaging over a time window.
@@
 Due to its simplicity and practicality, this method is commonly used in the literature (see e.g.
-[4]) and is also adopted in the present study.
-We therefore analyze the time series $x_{1}(t)$ from the long simulation of the triad system in Eq.
+[Penland95]) and is also adopted in the present study.
+We consider again the time series $x_{1}(t)$ from the long simulation of the triad system in Eq.
 (6).
@@
 Results are shown in Figure 5.
-As before, the stochastic model with additive noise fails to capture the correct stationary and perturbed statistics.
-By contrast, the model with multiplicative noise not only reproduces the stationary distribution and improves the mean response, but also closely matches the variance response, capturing its peak magnitude and deviating only slightly in its temporal decay.
-One could, in principle, tune the averaging window to match the autocorrelation of the emulated $x(t)$ with that of the original $x_{1}(t)$, thereby ensuring closer agreement in the decay of responses (Figure 5(e,f)).
-We view this as a methodological refinement for future work, and leave it outside the scope of this study.
+Now, the model with multiplicative noise not only reproduces the stationary distribution but also closely matches the variance response, capturing its peak magnitude and deviating only slightly in its temporal decay.
+One could, in principle, tune the averaging window to match the autocorrelation of the emulated $x(t)$ with that of the original $x_{1}(t)$, thereby ensuring closer agreement in the decay of responses.
+Furthermore, training on multiple time steps should also result in a more accurate model.
+We view these as methodological refinements for future work, and leave it outside the scope of this study.
 Figure 5: Same as in Figure 4 however here the scalar emulators in Eq.
@@
 Even with access to very long time series, modeling real-world dynamical systems from data alone remains fundamentally constrained by observing only a subset of the full state vector.
-This creates two major challenges: (i) the model’s skill depends critically on the variables chosen, and (ii) once these variables are fixed, the influence of unobserved degrees of freedom must be adequately parametrized.
-While such issues are illustrated with low-dimensional models, they become far more acute in realistic spatiotemporal climate systems, where time-scale separation is weak and variable selection involves nontrivial choices of fields, dimensionality reduction methods, and spatiotemporal scales.
-There is no rigorous mathematical framework to guide these choices from data alone.
-Even in cases where the governing equations are known, such as in meteorology, and where large computational resources are available, emulators still operate on finite-resolution data, leading to subgrid processes.
-The cumulative effect of such unobserved processes still require parametrizations.
-This highlights that the problem of parameterizations, traditionally associated with numerical modeling [92, 93], is equally central to data-driven approaches with neural emulators.
-These considerations highlight intrinsic limitations of general-purpose emulators in capturing responses to external perturbations and in supporting Green’s function–like experiments [38].
+The most severe limitation arises from the choice of variables: (i) the model’s skill depends critically on which variables are observed and included.
+Once these variables are fixed, (ii) the influence of unobserved degrees of freedom must be adequately parametrized to capture responses in probability distributions.
+Although these issues can be illustrated with low-dimensional examples, they become much more severe in realistic spatiotemporal climate systems, where time-scale separation is weak and variable selection involves nontrivial decisions about which fields to include, how to perform dimensionality reduction, and which spatial and temporal scales to retain — that is, how to construct an appropriate coarse-grained representation of the system.
+There is no rigorous mathematical framework to guide these choices from data alone in high-dimensional dynamical systems.
+These considerations highlight intrinsic limitations of general-purpose emulators in capturing responses in mean and variance of the distribution to external perturbations.
 In contrast, reduced-order neural models are often tailored to specific problems, allowing one to leverage prior literature, insights, and results to guide variables and coarse-graining choices.
-In the next section, we present a concrete example where a simple reduced-order neural model succeeds in a task where a state-of-the-art neural emulator showed limitations [37].
+Such prior knowledge can be essential for constructing data-driven models and for clearly delineating their domains of validity.
 IV A real-world example: radiative responses to temperature perturbations in a climate model
-Building on the insights discussed above, we present a stochastic reduced-order model tailored to a specific real-world test-case.
-We first introduce the scientific problem at hand and the form of the proposed emulator to tackle this specific question.
+Building on the insights discussed above, we present a stochastic reduced-order model tailored to a specific real-world, high-dimensional test-case.
+We introduce the scientific problem at hand and the proposed emulator to tackle this specific question.
 We then highlight the coarse-graining choices defining the proper variables to model.
-Finally, we showcase the skill of this emulator to study causal mechanisms in large-scale climate dynamics.
+Finally, we showcase the skill of this emulator in studying causal mechanisms in large-scale climate dynamics, as well as its limitations and future work.
 IV.1 Scientific problem at hand
@@
 Radiative forcing (e.g., from changes in $\text{CO}_{2}$) drives surface warming, which alters sea surface temperatures (SSTs) and associated radiative feedbacks.
-Recent studies show that even for the same global-mean warming, different SST warming patterns can yield markedly different feedbacks.
-For example, eastern Pacific warming tends to amplify warming through positive feedbacks, while western Pacific (warm pool) warming has a stabilizing effect [38, 39].
-This sensitivity of Earth’s energy imbalance to spatial SST patterns is known as the “pattern effect.” The pattern effect describes the fast response of top-of-atmosphere (TOA) radiative fluxes to spatial temperature variations and can be formalized in energy balance models via a Taylor expansion around the steady-state temperature [94].
-A common approach to study the pattern effect is to relate SST patterns to TOA radiative fluxes using Green’s function experiments [95], where the radiative response to localized SST perturbations is quantified with atmosphere-only models [96, 97, 98, 39, 99].
+Recent studies show that even for the same global-mean warming, different SST warming patterns can yield different feedbacks.
+For example, eastern Pacific warming tends to amplify warming through positive feedbacks, while western Pacific warming has a stabilizing effect [blochjohnsonetal2024, Dong2019AttributingPacific].
+This sensitivity of Earth’s energy imbalance to spatial SST patterns is known as the “pattern effect.” The pattern effect describes the fast response of top-of-atmosphere (TOA) radiative fluxes to spatial temperature variations and can be formalized in energy balance models via a Taylor expansion around the steady-state temperature [Meyssignac].
+A common approach to study the pattern effect is to relate SST patterns to TOA radiative fluxes using Green’s function experiments [barsugli2002global], where the radiative response to localized SST perturbations is quantified with atmosphere-only models [zhouetal2017, Zhang2023UsingCM4, Dong2020IntermodelModels, Dong2019AttributingPacific, alessi2023surface].
 These experiments are valuable for testing model sensitivities but are computationally costly and typically require large SST perturbations (1–4 K) to obtain a measurable response within short simulations.
-Such large perturbations can trigger nonlinear behavior, undermining the linearity assumptions of the Green’s function framework [100].
+Such large perturbations can trigger nonlinear behavior, undermining the linearity assumptions of the Green’s function framework [williams2023circus].
 This motivates the need for fast, interpretable data-driven models that can simulate responses of large ensembles at low cost and with small perturbation amplitudes.
@@
 However, as shown by Van Loon et al.
-(2025) [37] in the context of the ACE2 emulator and the pattern effect, this strategy may not always succeed.
-A second, more targeted approach for large-scale climate dynamics is to design an emulator tailored to a given problem, e.g.
-[5].
-While this sacrifices generality, it offers key advantages: clarifying the choice of variables and scales to model, grounding the design in physical intuition and prior literature, and isolating only the processes most relevant to the problem.
-This focus also enhances interpretability and helps bridging the gap between simulation and understanding [101].
-IV.2 A reduced order model of surface temperature and radiative fluxes
-We aim to build a skillful emulator of large-scale sea surface temperature (SST) and radiative fluxes capable of capturing both stationary and perturbed statistics.
-Building on the previous sections, we adopt the following functional form for the neural model:
+(2025) [Senne] in the context of the ACE2 emulator and the pattern effect, this strategy may not always succeed.
+A second approach for large-scale climate dynamics is to design an emulator tailored to a given problem, e.g.
+[PENLAND1996534].
+While this sacrifices generality, it can also offer key advantages by clarifying the choice of variables and scales to model, grounding the design on prior literature, and isolating only the processes most relevant to the problem.
+IV.2 A reduced-order stochastic model of surface temperature and radiative fluxes
+Goals and model formulation.
+Our goal is to develop a skillful emulator of large-scale sea surface temperature (SST) and radiative fluxes that captures both stationary variability and responses to perturbations, enabling the study of their causal interactions.
+In this work, we focus on ensemble-mean responses, which allows for direct comparison with previous studies.
+We adopt the following functional form for the neural model:
 $d\mathbf{x}=(\mathbf{L}\mathbf{x}+\mathbf{n}(\mathbf{x}))dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W},$ (11)
-Where $\mathbf{W}$ is the standard Wiener process. As before, to improve readability we refer to Appendix C for technical details.
-As outlined throughout the paper, a central challenge lies in selecting preprocessing steps to define the coarse-grained dynamics $\mathbf{x}$ to be modeled.
-These choices are critical for identifying the proper variables, since the climate system involves not only SST and radiative fluxes but also many other interacting variables.
+where $\mathbf{W}$ is the standard Wiener process.
+In Appendix C, we propose a practical algorithm to fit the multivariate multiplicative noise $\mathbf{\Sigma}(\mathbf{x})$.
+Differences with the case of additive noise mainly arise in the reconstruction of the stationary density as further discussed in Appendix D.
+As outlined throughout the paper, a central challenge in data-driven modeling lies in selecting preprocessing steps to define the coarse-grained dynamics $\mathbf{x}$ to be modeled.
+These choices, guided by the study’s objectives, are critical for identifying the proper variables, since the climate system involves not only SST and radiative fluxes but also many other interacting variables.
 This is thus a practical instance of a partially observed state vector.
-Below, we describe the dataset and preprocessing procedure, which largely follows [102], with further details provided in that work.
+Below, we describe the dataset and preprocessing procedure, which largely follows [FDTPatternEffect], with further details provided in that work.
 IV.2.1 Data, preprocessing and the choice of proper variables
 Data.
-We consider outputs from the state-of-the-art coupled climate model GFDL-CM4 [103], which provides a 600-year pre-industrial control run with constant $\text{CO}_{2}$.
-The ocean component is MOM6 [104] with 0.25 horizontal resolution and 75 vertical layers, and the atmosphere is AM4 [105, 106] with  100 km resolution and 33 vertical layers.
+We consider outputs from the state-of-the-art coupled climate model GFDL-CM4 [cm4], which provides a 600-year pre-industrial control run with constant $\text{CO}_{2}$.
+The ocean component is MOM6 [OM4] with 0.25 horizontal resolution and 75 vertical layers, and the atmosphere is AM4 [AM4a, AM4b] with  100 km resolution and 33 vertical layers.
 The model in Eq.
@@
 We consider two climate fields: sea surface temperature (SST) and the global mean net radiative flux at the top of the atmosphere (TOA).
-The SST field is restricted to latitudes $45^{\circ},45^{\circ}]$, following earlier studies that highlight the importance of radiative feedbacks in the tropics and subtropics [39, 97].
+The SST field is restricted to latitudes $45^{\circ},45^{\circ}]$, following earlier studies that highlight the importance of radiative feedbacks in the tropics and subtropics [Dong2019AttributingPacific, Zhang2023UsingCM4].
 Choosing the proper variables.
-Climate dynamics involves many interacting fields beyond SST and radiative fluxes.
-Our goal is a reduced-order model that captures the coarse-grained dynamics while representing unresolved processes as noise.
-We coarse-grain the GFDL-CM4 control run in space and time, averaging daily data to monthly means and removing the seasonal cycle to focus on anomalies rather than periodic signals driven by solar forcing.
-Considering 1 month averages in TOA fluxes is not arbitrary: the immediate response of TOA fluxes to perturbations in SST is fast (e.g., $\sim 20$ days in tropical large-scale convective overturning motion in the GFDL-CM4 model) [102].
-This averaging filters out higher-frequency processes that mediate surface–TOA coupling, yielding a system that is more Markovian.
+We coarse-grain the GFDL-CM4 control run in space and time, averaging daily data to monthly means and removing the seasonal cycle to focus on anomalies rather than periodic signals.
+Considering 1 month averages in TOA fluxes is not arbitrary: the immediate response of TOA fluxes to perturbations in SST is fast (e.g., $\sim 20$ days in tropical large-scale convective overturning motion in the GFDL-CM4 model) [FDTPatternEffect].
+This averaging filters out higher-frequency processes that mediate surface–TOA coupling, leading to a “more Markovian” system .
 Given so, direct causal links between TOA fluxes at SST patterns can be detected as one time step responses; see Appendix A.
-We then high-pass filter with a cut-off of $10^{-1}$ years, removing (i) low-frequency (multidecadal) oscillations, which are poorly sampled even in 600-year runs, and (ii) slow drifts from incomplete ocean equilibration.
-This choice sacrifices the ability to capture multidecadal processes but ensures robust sampling of subdecadal statistics, yielding a larger effective sample size.
+We then high-pass filter the data with a cut-off of $10^{-1}$ years, removing (i) low-frequency (multidecadal) oscillations, which are poorly sampled even in 600-year runs, and (ii) slow drifts from incomplete ocean equilibration.
+This choice sacrifices the ability to capture multidecadal variability but ensures robust sampling of subdecadal statistics, yielding a larger effective sample size.
 Implicitly, the choice assumes that the response of the net radiative flux to SST perturbations occurs within a 10-year window.
-Next, we decompose the SST field $\mathbf{X}(t)\in\mathbb{R}^{N}$ into empirical orthogonal functions (EOFs):
+Next, we project the SST field $\mathbf{X}(t)\in\mathbb{R}^{N}$ onto an empirical orthogonal functions (EOFs) basis:
 $\mathbf{X}(t)=\sum_{i=1}^{n}x_{i}(t)\mathbf{e}_{i}+\sum_{j=n+1}^{N}y_{j}(t)\mathbf{e}_{j},$ (12)
-and retain the first $n=20$ modes $x_{i}(t)$.
-The combined steps of monthly averaging, high-pass filtering and projection over EOF bases define a coarse-grained system for the slow SST variability.
-This choice follows earlier work showing the skill of linear Markov models built on similar preprocessing [4, 5].
-For radiative fluxes, we focus on the global mean TOA net flux, as the TOA flux exhibits little spatial structure and we are mainly interest in its average value.
+and retain the first $n=20$ modes.
+The combined steps of monthly averaging, high-pass filtering and projection over EOF bases define a coarse-grained representation of the SST variability.
+This choice follows earlier work showing the skill of linear Markov models for SST built on similar preprocessing [Penland95, PENLAND1996534].
+For radiative fluxes, we focus on the global mean TOA net flux, as the TOA flux exhibits little spatial structure and we are mainly interested in its average value.
 The state vector is then defined as $\mathbf{x}(t)\in\mathbb{R}^{n+1}$, consisting of the $n$ SST principal components concatenated with the global mean TOA flux.
+Crucially, these choices yield an imperfect model that is skillful in some respects—such as capturing probability distributions and ensemble-mean responses to temperature perturbations—but inadequate in others, including autocorrelation structure and responses in ensemble variance.
+We discuss these successes and limitations in the following sections.
 Forced simulations
-We test the model’s ability to reconstruct changes in global-mean TOA radiative flux given prescribed SST trajectories from two 150-years long forced runs: 11pctCO2 and $4\timesCO$2.
+We will test the model’s ability to reconstruct changes in global-mean TOA radiative flux given prescribed SST trajectories from two 150-years long forced runs: 11pctCO2 and $4\timesCO$2.
 The 11pctCO2 experiment imposes a 1% $\text{CO}_{2}$ increase per year starting from pre-industrial conditions, while the $4\timesCO$2 experiment abruptly quadruples $\text{CO}_{2}$ at $t=0$.
@@
 To address this, we remove the portion of change of the net TOA radiative flux that is attributable solely to rising $\text{CO}_{2}$.
-This allow us to isolate and study the component of radiative feedback that is driven by changes in SST.
-This correction exploits the logarithmic dependence of global-mean forcing on $\text{CO}_{2}$ concentration [107, 108].
-As is standard, we subtract $8\text{Wm}^{-2}$ in the $4\timesCO$2 run [109], and apply a time-dependent correction $\alpha\log(C_{t}/C_{0})$ in the 11pctCO2 run, where $C_{t}$ and $C_{0}$ denote $\text{CO}_{2}$ concentrations in the forced and control simulations.
+This allow us to isolate and study the component of radiative feedback that is driven by changes in SST alone.
+This correction exploits the logarithmic dependence of global-mean forcing on $\text{CO}_{2}$ concentration [Pierrehumbert, Romps].
+As is standard, we subtract $8\text{Wm}^{-2}$ in the $4\timesCO$2 run [Zhao22investigation], and apply a time-dependent correction $\alpha\log(C_{t}/C_{0})$ in the 11pctCO2 run, where $C_{t}$ and $C_{0}$ denote $\text{CO}_{2}$ concentrations in the forced and control simulations.
 IV.2.2 Stationary statistics and forced responses
@@
 600 years at monthly time step).
-The stationary probability distributions of the leading SST mode and the net radiative TOA flux are well reproduced (Fig.
-6a,b).
-The main discrepancies arise in the autocorrelation functions, shown for the first SST mode (closely linked to ENSO [110]) and the TOA flux.
+The stationary probability distributions of the most energetic SST mode, corresponding to the El Niño Southern Oscillation (ENSO) [ENSOcomplexity], and the net radiative TOA flux are well reproduced Figures 6(a,b).
+Statistics and comparisons with the case of additive noise are reported in Appendix D.
+The main discrepancies arise in the autocorrelation functions, shown for the first SST mode and the TOA flux in Figure 6(c,d).
+This limitation is further discussed in Section “Limitations” below.
 Forced responses.
@@
 Before using our emulator for Green’s function–like studies to investigate the pattern-effect, we first assess its skill as a reconstruction tool.
-Specifically, we reconstruct global-mean TOA flux from prescribed SST trajectories in the two forced experiments, i.e.
-11pctCO22 and $4\timesCO$22.
+Specifically, we reconstruct the global-mean TOA flux from two prescribed, forced SST trajectories, i.e.
+from the 11pctCO22 and $4\timesCO$22 simulations.
 Forced SSTs are projected onto the first $n=20$ EOFs from the control run, yielding coarse-grained trajectories $\mathbf{x}_{\text{SST}}^{f}(t)$, $f$ denotes “forced.” Given $\mathbf{x}_{\text{SST}}^{f}(t)$, the emulator returns the TOA variable $x_{\text{TOA}}(t+1)$ at the next time step, initialized with $x_{\text{TOA}}(0)=0$.
-Results (Fig.
-6e,f) show that the emulator largely reproduces TOA flux changes in both scenarios.
-This reconstruction is possible because we focus on one-month averages, which reduce the time-dependent TOA response, driven by an atmospheric reorganization at sub-monthly scales, into a one-step prediction.
+Results (Figure 6(e,f)) show that the emulator largely reproduces TOA flux changes in both scenarios.
+This reconstruction is possible because of the focus on one-month averages, reducing the time-dependent TOA response, driven by an atmospheric reorganization at sub-monthly scales, into a one-step prediction.
 For the 1pctCO2 experiment, the linear trend in net TOA flux is $0.043Wm^{-2}yr^{-1}$ in GFDL-CM4 and $0.034Wm^{-2}yr^{-1}$ in the emulator.
-We then compute correlations between reconstructed and simulated radiative flux at the TOA, after removing a linear trend in 1pctCO2 and an initial 50-year transient in 4xCO2.
+We then compute correlations between the reconstructed and simulated radiative flux at the TOA, after removing a linear trend in 1pctCO2 and an initial 50-year transient in 4xCO2.
 At monthly resolution, correlations are $r=0.37$ (1pctCO2) and $r=0.43$ (4xCO2).
 If yearly averages are considered (as in the current literature, see e.g.
-[38]) these increase to $r=0.70$ (1pctCO2) and $r=0.74$ (4xCO2).
+[blochjohnsonetal2024]) these increase to $r=0.70$ (1pctCO2) and $r=0.74$ (4xCO2).
 All correlations are significant at the $5$ sigma level under a null hypothesis of zero correlation, tested via bootstrap resampling with replacement.
@@
 Note: the net flux at the TOA variable is negative upward (i.e., leaving the planet).
-IV.2.3 Interpretability and understanding: ensemble mean and variance responses
+IV.2.3 Interpretability and causal understanding through ensemble mean responses
 In the previous section, we showed that the reduced-order neural emulator can reconstruct the net radiative flux at the top of the atmosphere (TOA) from prescribed changes in the sea surface temperature (SST) field.
-This amounts to the mapping $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $\mathbf{x}_{\text{SST}}(t)\in\mathbb{R}^{n}$, $x_{\text{TOA}}(t)\in\mathbb{R}$, and $x_{\text{TOA}}(0)=0$.
+This amounts to the mapping $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $\mathbf{x}^{\text{f}}_{\text{SST}}(t)\in\mathbb{R}^{n}$, $x_{\text{TOA}}(t)\in\mathbb{R}$, and $x_{\text{TOA}}(0)=0$.
 We now ask the following question: What are the SST patterns that contribute the most to this skillfull prediction?
-Addressing this question requires moving beyond prediction to identify the causal influence of individual SST components on radiative flux.
+Addressing this question requires moving beyond prediction tasks and identify the causal influence of individual SST components on radiative fluxes.
 We frame the problem in terms of linear response theory, focusing on the direct link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $x_{j,\text{SST}}(t)$ denotes the $j$-th SST component.
-Response in ensemble mean.
-To quantify this link, we perturb the emulator in Eq.
-(11) following the procedure in Appendix A.
-Briefly, at $t=0$, we apply a small impulse $\Delta_{j}=10^{-2}\sigma_{j}$ to the $j$-th SST component, with $\sigma_{j}$ its stationary standard deviation.
-We then compute the ensemble-mean response of the net TOA flux at $t=1$, denoted $\delta\langle x_{\text{TOA}}(1)\rangle$.
-This quantity, referred to as $R^{\mu}_{\text{TOA},j}(1)$, represents the ensemble-mean response ($\mu$ indicates the mean) and is estimated over $10^{6}$ initial conditions to ensure robust statistics
-Response in ensemble variance.
-While the ensemble-mean response is informative, it captures only part of the adjustment.
-Large-scale climate responses can involve shifts of probability distributions.
-To partially account for this, we quantify the response in the variance of the net TOA radiative flux to surface warming perturbations.
-As above, the estimate is based on an ensemble of $10^{6}$ trajectories.
-For consistency in physical units, we express the response in terms of standard deviation rather than variance, denoting it by $R^{\sigma}_{\text{TOA},j}(1)$ ($\sigma$ indicates the standard deviation).
-This quantity measures the change in ensemble spread following a perturbation of the $j$-th SST component.
-Conceptually, this procedure provides a computationally efficient analogue of atmospheric-only Green’s function experiments, while shifting the focus from individual trajectories to full probability distributions.
-Sensitivity maps.
-The vectors of one-time-step responses in mean and standard deviation, $\mathbf{R}^{\mu}$ and $\mathbf{R}^{\sigma}$, are projected back to the full $N$-dimensional SST field using the EOFs, $\mathbf{e}\in\mathbb{R}^{n,N}$:
-$\begin{split}\mathbf{S}^{\mu,\sigma}=\mathbf{R^{\mu,\sigma}_{\text{TOA}}}(1)\cdot\mathbf{e}=\sum_{j}^{n}R^{\mu,\sigma}_{\text{TOA},j}(1)~\mathbf{e}_{j},\end{split}$ (13)
-where $\mathbf{S}^{\mu},\mathbf{S}^{\sigma}\in\mathbb{R}^{N}$ represent sensitivity maps in mean and standard deviation, respectively.
-At each ocean grid point $i$, these maps approximate the sensitivities $\frac{\partial x^{\mu,\sigma}_{\text{TOA}}}{\partial x_{\text{SST},i}}$, where $x^{\mu}_{\text{TOA}}$ and $x^{\sigma}_{\text{TOA}}$ respectively represent the ensemble mean and standard deviation of the mean, net radiative flux and $x_{\text{SST},i}$ is the SST at location $i$.
-The sensitivity maps in mean ($\mathbf{S}^{\mu}$) and standard deviation ($\mathbf{S}^{\sigma}$) are shown in Figure 7.
-The obtained sensitivity map $\mathbf{S}^{\mu}$ is in qualitative agreement with previous results in response theory (Figure 2(a) in [102]), with other data-driven approaches (see Figure 7(b) in [111] and Figure 2 in [112]), but, most importantly, with the sensitivity map obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [38], see Figure 5(a) in [39]).
+We refer the reader to Appendix A for details on the causal interpretation of the response operator.
+We apply an impulse to the $j$-th SST component at $t=0$ and compute the ensemble-mean response of the net TOA flux at $t=1$, denoted $\delta\langle x_{\text{TOA}}(1)\rangle$ (see Appendix A for details).
+This quantity, referred to as $R_{\text{TOA},j}(1)$, represents the one-step ensemble-mean response and is estimated using $10^{6}$ initial conditions to ensure robust statistics.
+By focusing on the one-step response, we isolate fast atmospheric adjustments while neglecting the slower oceanic feedbacks (see Section 7 in [FDTPatternEffect] for details).
+This setup is thus conceptually closer to atmospheric-only Green’s function experiments, where the ocean acts as a prescribed boundary condition rather than a dynamically coupled component of the system.
+The vector $\mathbf{R}_{\text{TOA}}(1)\in\mathbb{R}^{n}$ is projected back to the full $N$-dimensional SST field using the EOFs, $\mathbf{e}\in\mathbb{R}^{n,N}$:
+$\begin{split}\mathbf{S}=\mathbf{R_{\text{TOA}}}(1)\cdot\mathbf{e}=\sum_{j}^{n}R_{\text{TOA},j}(1)~\mathbf{e}_{j}.\end{split}$ (13)
+At each ocean grid point $i$, the sensitivity map $\mathbf{S}\in\mathbb{R}^{N}$ approximates the sensitivity $\frac{\partial x_{\text{TOA}}}{\partial x_{\text{SST},i}}$, where $x_{\text{TOA}}$ is the global mean radiative flux at the TOA and $x_{\text{SST},i}$ is the SST at location $i$.
+The sensitivity map $\mathbf{S}$ is shown in Figure 7.
+The obtained sensitivity map $\mathbf{S}$ is in qualitative agreement with previous results in response theory (Figure 2(a) in [FDTPatternEffect]), with other data-driven approaches (see Figure 7(b) in [kang2023recent] and Figure 2 in [senneNN]), and with the one obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [blochjohnsonetal2024], see Figure 5(a) in [Dong2019AttributingPacific]).
 In the latter case, the maps exhibit fewer spatial features due to the coarse resolution of the perturbation patches (spanning tens of degrees) in numerical models experiments.
+For additional statistical reconstruction approaches to the pattern effect that yield structures similar to those in Figure 7, see the recent analysis by Fredericks et al.
+(2025) [Leif].
 Figure 7: Sensitivity maps obtained using the neural emulator in Eq.
 (11).
-At each SST grid point $i$, we show the one–time-step response in the ensemble mean ($\mathbf{S}^{\mu}$) and ensemble standard deviation ($\mathbf{S}^{\sigma}$) of the global-mean net radiative flux at the TOA following an impulse perturbation in SST at that location.
+At each SST grid point $i$, we show the one–time-step response in the ensemble mean ($\mathbf{S}^{\mu}$) of the global-mean net radiative flux at the TOA following an impulse perturbation in SST at that location.
 The computation is carried out in a 21-dimensional reduced space spanned by empirical orthogonal functions, and the maps represent the projection of these low-dimensional results back to the full grid.
-Responses are estimated from an ensemble of $10^{6}$ initial conditions.
-The dominant feature of $\mathbf{S}^{\mu}$ is a dipole in the tropical Pacific: negative sensitivity in the west and positive sensitivity in the east [39].
+The dominant feature of $\mathbf{S}$ is a dipole in the tropical Pacific: negative sensitivity in the west and positive sensitivity in the east [Dong2019AttributingPacific].
 This implies that eastern Pacific warming produces a positive radiative feedback, while western Pacific warming produces the opposite.
-The mechanisms are well established [113, 100]: in the east, high climatological low-level cloud cover decreases under warming, reducing reflected shortwave radiation and yielding positive sensitivity.
+The mechanisms are well established [Pascale, williams2023circus]: in the east, high climatological low-level cloud cover decreases under warming, reducing reflected shortwave radiation and yielding positive sensitivity.
 In the west, deep convection transports surface warming vertically and then horizontally via gravity waves, strengthening the inversion over the eastern Pacific and other stratocumulus regions, increasing low-level cloud fraction and reflected shortwave radiation, thus producing a negative feedback.
-This analysis confirms that the accurate reconstructions in the forced regime (Fig.
-6e,f) arise from physically consistent mechanisms, and that coarse-grained neural models can recover such causal linkages.
-Finally, the variance response $\mathbf{S}^{\sigma}$ (Fig.
-7b) highlights regions where perturbations alter the spread of radiative responses across ensemble members.
-Perturbations in stratocumulus-dominated regions (northeast Pacific, off the coast of South America) yield negative variance responses: the ensemble distribution of net radiative flux narrows relative to the control, highlighting small uncertainties in the mean response studied above.
-By contrast, surface warming perturbations in the central and western Pacific lead to positive variance responses, reflecting a broader spread of the net radiative flux distribution after SST perturbations.
-To our knowledge, such variance responses have not been quantified before.
-Unlike the response in the mean, which has been examined extensively and shows broad agreement across atmospheric models [38], the response in ensemble variance has no established benchmark.
-Because it reflects changes in variability, it may differ significantly across atmospheric models.
-Moreover, estimating such responses requires very large ensembles, which are prohibitively expensive in conventional climate models.
-For this reason, the results presented here should be viewed as an exploratory step: they highlight potentially important regional structures, but a full climate interpretation will require systematic intercomparison and validation against other modeling frameworks, such as novel data-driven tools to estimate response through the fluctuation-dissipation theorem [114].
-IV.2.4 Main takeaways
-In this Section, we developed a stochastic neural emulator that captures the joint evolution of the sea surface temperature (SST) field and the global mean net radiative flux at the top of the atmosphere (TOA) in the GFDL-CM4 climate model.
+This analysis confirms that the accurate reconstructions in the forced regime (Figure 6(e,f)) arise from physically consistent mechanisms.
+IV.2.4 Limitations and a few ways forward
+Our coarse-grained representation enabled the development of a data-driven model to study the direct causal interaction $x_{\text{SST}}\rightarrow x_{\text{TOA}}$ through perturbation experiments, while also capturing the stationary distribution well.
+However, these coarse-graining choices also introduce some drawbacks.
+We discuss these limitations below, further delineating the model’s domain of validity and outlining possible ways forward:
+  • The model does not fully capture the decay of autocorrelations.
+For example, the first SST mode in Figure 6(a) suggests residual non-Markovian behavior.
+This is not an issue for our focus on direct causal links (responses at the shortest time step), but it would matter for responses over longer time horizons.
+Remedies include: (i) expanding the state vector by including slower oceanic variables and more EOF components, (ii) incorporating memory terms in the stochastic parameterization [KONDRASHOV201533], or (iii) further coarse-graining the temporal resolution (e.g., 3-month averaging).
+Such choices will depend on the intended scientific application.
+  • We found that responses in ensemble variance are sensitive to retraining the multiplicative noise term, so we do not present these results.
+The fitting algorithm in Appendix C requires far more data than available here (only 7200 samples for training and validation), to achieve high accuracy.
+In this data-scarce setting, exploring simple analytical forms for the multiplicative noise may be beneficial and it is left for future work.
+Importantly, data-driven approaches, such as the Fluctuation–Dissipation Theorem (FDT) [MajdaBook, Marconi], can provide an approximate benchmark for responses.
+This allows calibration of models with different stochastic terms, particularly to test variance responses where direct benchmarks are often unavailable.
+This strategy aligns with recent proposals such as [majdaSIAM] and will be explored in future work.
+  • While our emulators are empirically stable, stability is not guaranteed, which may matter if tested against large perturbations.
+Future work could impose additional constraints on the neural architectures, similar to previous approaches for physically constrained regression models as in [majdaConstraints, KONDRASHOV201533].
+IV.2.5 Main takeaways
+In this Section, we developed a stochastic neural emulator that captures the joint evolution of the sea surface temperature field and the global mean net radiative flux at the top of the atmosphere in the GFDL-CM4 climate model.
 Addressing a specific, well-posed problem, proved crucial in guiding the construction of the model: it enabled us to build on previous literature, restrict attention to just two variables and a limited range of spatial and temporal scales.
@@
 Such experiments focused on very large ensembles ($10^{6}$ trajectories), currently impossible to perform with climate models.
-These results demonstrate that reduced-order neural emulators can extend beyond reproducing stationary climate dynamics to enable efficient, ensemble-based perturbation experiments.
-V Discussion and Conclusion
-In summary, the first part of this work discussed the limitations of purely data-driven models of complex climate systems.
+V Discussion
+In summary, the first part of this work discussed important limitations of purely data-driven models of complex climate systems.
 We highlighted the potential of combining traditional reduced-order modeling approaches, including state-dependent (multiplicative) stochastic parameterizations, with recent advances in neural networks to study large-scale, low-frequency climate dynamics.
@@
     Unknown equations, partially observed state vector.
-This more realistic case presents fundamental difficulties, even with access to high-quality data and powerful neural models architectures.
-The model skill now depends critically on two factors: (a) the choice of variables to model and (b) how unobserved degrees of freedom are parameterized.
-These results highlight intrinsic limitations of purely data-driven-only modeling in high-dimensional climate systems, where the full set of relevant variables is typically unknown and no rigorous framework exists for selecting the proper variables to model.
-By variable selection we mean the choice of climate fields as well as their spatial and temporal scales (also including dimensionality reduction methodologies).
-Such choices are guided by the physical process of interest, and in the case of low-frequency climate dynamics, they motivate the development of reduced-order neural emulators tailored to specific phenomena as efficient alternative to general purpose emulators.
-A central challenge is to parameterize the cumulative effect of unresolved degrees of freedom.
-Model performance (especially in capturing responses to external forcings) depends critically on this choice and we showed that multiplicative (state-dependent) noise is preferable to additive noise.
-We emphasize that the analysis presented here applies to data-driven models of turbulent geophysical systems that must capture both stationary statistics and responses to external perturbations, as in (causal) Green’s function experiments [38].
-For Earth system models, these issues mainly concern autoregressive climate emulators, particularly coupled atmosphere–ocean systems, and should not be confused with short-term weather forecasting using neural networks (e.g., [115, 23]), where long-term stability and forced responses are less relevant.
-Our study also does not address statistics-based emulators for climate change projections, which have proven effective for investigating responses of key climate variables to CO2 forcing [116, 117]; see [116] for an example where simple approaches, such as linear pattern scaling, can outperform more complex models.
+This more realistic case presents fundamental difficulties, even with access to high-quality data and powerful neural model architectures.
+The most critical limitation is (a) identifying an appropriate coarse-grained representation; once this is determined, (b) the cumulative effect of unobserved degrees of freedom needs to be accurately parametrized.
+Learning such parametrizations from data remains a central challenge, and we showed that multiplicative noise can be advantageous over additive noise, though it requires a larger dataset for reliable training.
+These results highlight intrinsic limitations of purely data-driven modeling of high-dimensional, multiscale climate systems, where the full set of relevant variables is unobserved or even unknown  [^2] A brief remark: in partially observed settings, one might be tempted to appeal to Takens’ embedding theorem [Takens].
+This is not a practical option: the theorem does not extend to stochastic systems, and it becomes infeasible in high-dimensional settings [Cecconi, Baldovin, Lucente]..
+In this context, the “correct” choice of variables, scales, and stochastic parameterizations depends ultimately on the specific scientific goals [Smith2001].
+This perspective helps explain the continued success of Linear Inverse Models (LIMs) proposed by Penland, C.
+(1989) [Penland89], which rely on careful selection of climate fields and preprocessing choices to capture targeted aspects of climate variability; see, for example, [Penland95, PENLAND1996534, Penland2025] among others.
+We emphasize that the analysis presented here applies to data-driven models of turbulent geophysical systems that must capture both stationary statistics and responses to external perturbations, as in (causal) Green’s function experiments [blochjohnsonetal2024].
+For Earth system models, these issues mainly concern autoregressive climate emulators, particularly coupled atmosphere–ocean systems, and should not be confused with short-term weather forecasting using neural networks (e.g., [AIFS, fourcastnet3]), where long-term stability and forced responses are less relevant.
+Our study also does not address statistics-based emulators for climate change projections, which have proven effective for investigating responses of key climate variables to CO2 forcing [Bjorn, Gosha, AndreEmulation].
 Future work.
@@
 First, a stochastic formulation of neural autoregressive climate emulators is preferable to a purely deterministic one.
-Probabilistic formulations align more closely with the inherently stochastic and uncertain nature of climate dynamics [118].
-Recent advances in weather emulation already point in this direction such as FourecastNet3 [23], GenCast [23] and the novel work of Alet et al.
-(2025) [119].
-Second, the ability of climate emulators in capturing forced responses in the full probability distribution, akin to Green’s function experiments, remains largely unexplored.
-A promising way forward is to evaluate novel architectures on their ability to respond to external perturbations on a hierarchy of simplified models relevant to geophysical fluid dynamics, such as the Lorenz–96 system [120].
+Probabilistic formulations align more closely with the inherently stochastic and uncertain nature of climate dynamics [PalmerQJRMS].
+Recent advances in weather emulation already point in this direction such as FourecastNet3 [fourcastnet3], GenCast [fourcastnet3] and the novel work of Alet et al.
+(2025) [Alet].
+New interesting directions also include probabilistic emulation in the framework of stochastic interpolants [ProbabilisticSI, ProbabilisticSIPedram, Nikolaj].
+Second, the ability of climate emulators in capturing forced responses in the full probability distribution, akin to Green’s function experiments, remains largely unexplored, especially for responses in higher-order moments.
+A promising way forward is to evaluate novel architectures on their ability to capture the mean and variance response to external perturbations in a hierarchy of simplified models relevant for large scale climate variability, such as the 2-scale Lorenz–96 system [Lorenz96].
 These controlled testbeds enable rigorous quantification of forced responses and could inform the design of future architectures for data-driven climate modeling.
@@
 However, for models constructed from observational data, such direct perturbations are not possible.
-In this context, one could, in principle, leverage the Fluctuation-Dissipation Theorem (FDT) to predict the response operator from data alone [81, 17].
-Traditionally, the quasi-Gaussian approximation have yield reliable estimates of impulse responses in the ensemble mean [121, 47].
-Recent advances in score-based generative models have shown the potential to extend these predictions to responses in higher-order moments [114].
+In this context, one can leverage the Fluctuation-Dissipation Theorem (FDT) to predict the response operator from data alone [MajdaStructuralStability, Baldovin].
+Traditionally, the quasi-Gaussian approximation have yield reliable estimates of impulse responses in the ensemble mean [GRITSUN, Majda2010].
+Recent advances in score-based generative models have shown the potential to extend these predictions to responses in higher-order moments [LudovicoFabriAndre].
 Future work.
-In the absence of direct perturbation experiments, FDT provides a rigorous bridge between stationary and forced dynamics [49, 48].
-With sufficiently long integrations of current climate emulators, one can readily focus on representative large-scale modes and estimate their impulse responses from the stationary record, thereby extending evaluation beyond stationary statistics with minimal computational cost.
-A recent example illustrating this approach in the climate context can be found in [18].
+In the absence of direct perturbation experiments, FDT provides a rigorous bridge between stationary and forced dynamics [Marconi, MajdaBook].
+With sufficiently long integrations of current climate emulators, one can readily focus on representative large-scale modes and estimate their impulse responses from the stationary record, thereby extending evaluation beyond stationary statistics at minimal computational cost.
+A recent example illustrating this approach in the climate context can be found in [FabCausal].
+Importantly, the analysis reveals the need to test emulators in their ability to respond to higher-order moments.
+[LudovicoFabriAndre] have shown that coupling the FDT with generative models provide a flexible framework to estimate responses in the full probability distribution.
 Moreover, the response operator at the shortest time step encodes direct causal links, i.e.
 the input–output relationships across variables in a reduced-order emulator (see Appendix A).
-Future work will leverage this information to suppress spurious, non-causal input-output dependencies during training, similarly to the recent work of Chen and Liu (2024) [59] on causally constrained, sparse reduced-order models.
+Future work will leverage this information to suppress spurious, non-causal input-output dependencies during training, similarly to the recent work of Chen and Liu (2024) [nanChen], leading to causally constrained, sparse reduced-order models.
 V.3 A realistic application: radiative responses to temperature perturbations
@@
 The model was trained on a long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies.
-The emulator focuses on the evolution of 20 principal components of the SST field and on the scalar global mean, net radiative flux; unobserved degrees of freedom are parametrized through multiplicative noise through an additional neural network.
-The emulator largely reproduces the stationary statistics of the system and, more importantly, allows us to reconstruct TOA radiative fluxes in two forced simulations.
-We inferred the direct causal links between temperature modes and TOA fluxes through a perturbation approach, akin to Green’s function-like experiments with climate models [38], targeting the so-called “pattern effect” [39].
-Crucially, the analyses targeted both responses in the ensemble mean and variance.
+The emulator focuses on the evolution of 20 principal components of the SST field and on the scalar global mean, net radiative flux; unobserved degrees of freedom are parametrized as multiplicative noise through an additional neural network.
+The emulator largely reproduces the stationary statistics of the system and allows us to reconstruct TOA radiative fluxes in two forced simulations.
+We inferred the direct causal links between temperature modes and TOA fluxes through a perturbation approach, akin to Green’s function-like experiments with climate models [blochjohnsonetal2024], targeting the so-called “pattern effect” [Dong2019AttributingPacific].
 Such responses are then projected back onto the EOFs to obtain maps of sensitivity in mean and variance change of the radiative flux at the TOA to SST warming perturbations.
-We argued that the sensitivity map for the ensemble mean is consistent with recent studies, indicating that the emulator’s skill in reconstructing TOA fluxes under forcing arises from physically meaningful mechanisms.
-In contrast, the sensitivity map of the ensemble variance is a novel result, not previously examined in the literature.
-It highlights the potential for uncertainty quantification in perturbation experiments but it requires further investigation to better understand the underlying physical mechanisms.
-These findings highlight the potential of reduced-order neural emulators to perform thousands of perturbation experiments, shifting the focus of climate feedback analysis from mean responses to the responses of probability distributions.
+We discussed similarities of the sensitivity map for the ensemble mean with recent studies, indicating that the emulator’s skill in reconstructing TOA fluxes under forcing arises from physically meaningful mechanisms.We also identified limitations.
+In particular, the variance response is sensitive to the training of the multiplicative-noise component, restricting the emulator’s reliability in perturbation experiments beyond those targeting the ensemble mean.
+We outline possible paths forward to resolve this issue.
+These findings highlight the potential of reduced-order neural emulators to perform thousands of perturbation experiments, eventually shifting the focus of climate feedback analysis from mean responses to the responses of probability distributions.
+VI Conclusion
 Modeling the climate system from data alone is constrained by observing only a subset of variables of a multiscale, high-dimensional dynamical system.
-This raises the fundamental challenge of identifying the appropriate variables (both fields and spatiotemporal scales) a fundamental problem in physics, emphasized in classical work [122] and revisited in recent studies [123].
-However, large-scale climate dynamics is often well characterized by a few recurrent spatiotemporal patterns, or modes of variability.
-In this case, the multiscale structure of the climate system can be an advantage allowing us to derive data-driven models of the coarse-grained dynamics while treating unresolved, fast processes stochastically.
+This raises the fundamental challenge of identifying the proper variables (both fields and spatiotemporal scales), a fundamental problem in physics emphasized in classical work [Onsager] and revisited in recent studies [Lucente].
+Such choices are not unique and are largely determined by the study’s scientific goal rather than by the pursuit of an unattainable perfect model [Smith2001].
+In contrast to numerical weather prediction, which demands high spatial and temporal resolution, low-frequency climate variability is dominated by a few recurrent spatiotemporal patterns, or modes of variability, and their interactions.
+Thus, at least for certain tasks, the multiscale structure of the climate system can be an advantage, allowing us to focus on the coarse-grained dynamics while treating unresolved, fast processes stochastically.
 Traditional reduced-order models stand out from general-purpose emulators by targeting specific climate problems from the outset, guided by prior literature, physical insight, and a clear understanding of the relevant variables and scales at play.
-Such data-driven modeling approach has a long history in climate science [2], though most efforts have emphasized reproducing stationary statistics or short-term forecasting [3, 4, 75] rather than forced responses.
-A promising perspective is to build on such established reduced-order approaches, focusing on powerful neural networks to fit deterministic and stochastic components and on tools from response theory, such as the FDT, to guide model development.
-Together, these strategies can enable efficient algorithms to study how full probability distributions respond to external perturbations, opening new avenues for causal inference and attribution in climate science.
+Such data-driven modeling approach has a long history in climate science [LucariniChekroun], though most efforts have emphasized reproducing stationary statistics or short-term forecasting [Penland89, Penland95, KONDRAENSO] rather than forced responses.
+A promising perspective is to build on such established reduced-order approaches, including those developed by Majda and Qi [majdaSIAM, QiMajdaChaos], while leveraging flexible neural networks architectures to discover low-dimensional latent representations [latentPierre, PierreLatent] and fit deterministic and stochastic components.
+Furthermore, tools from response theory such as the FDT [FabCausal, LudovicoFabriAndre, ValerioData2] can be used to guide model development and eventually lead to reduced-order models that efficiently capture responses in the mean and variance to external perturbations, advancing causal inference and attribution analyses in climate science.
 Appendix A Response operator: practical estimation and causal interpretation
@@
 A.2 Causal interpretation of the response operator
-Causality in natural systems is inferred through external interventions: an agent perturbs the system and the response is measured [77, 78].
-Climate modeling follows the same principle through Green’s function experiments, where one variable is perturbed and the response of another is evaluated [38, 124].
-Linear response theory formalizes this principle, providing a rigorous framework for causal analysis both in models and directly from data [17, 18, 114].
+Causality in natural systems is inferred through external interventions: an agent perturbs the system and the response is measured [IsmaelNew, Aurell].
+Climate modeling follows the same principle through Green’s function experiments, where one variable is perturbed and the response of another is evaluated [blochjohnsonetal2024, LucariniPRL].
+Linear response theory formalizes this principle, providing a rigorous framework for causal analysis both in models and directly from data [Baldovin, FabCausal, LudovicoFabriAndre].
 Because responses are inherently time-dependent, causal links through responses can be distinguished in three forms, all encoded in the time-dependent response operator: (i) direct links, (ii) indirect links, and (iii) cumulative links (or causal sensitivity).
@@
 Finally, we summarize recent approaches for estimating responses directly from data and refer to Baldovin et al.
-(2020) [17] for details on the theoretical foundations of this approach.
+(2020) [Baldovin] for details on the theoretical foundations of this approach.
 Consider a $N$ dimensional dynamical system of the kind:
@@
 Note from panel (b) that causality in dynamical systems can involve cycles, in contrast with the acyclic structures often assumed in applications outside physics, see e.g.
-[125].
+[Kusner].
 Causal relations can then be probed by adding independent impulse perturbations to each degree of freedom and record the system’s response.
@@
 Since we are interested in causal relations, it suffices to consider the identity observable $A(x_{k}(t))=x_{k}(t)$, i.e.
-the response of the ensemble mean [17].
+the response of the ensemble mean [Baldovin].
 We now clarify the three different types of causal links mentioned at the start of this Section.
@@
 The direct causal links are encoded in the response operator $\mathbf{R}(dt)$ evaluated at the shortest time step $t=dt$.
-The intuition follows directly from the graph representation in Fig.
-8(b) and is rather straightforward.
+The intuition follows directly from the graph representation in Figure 8(b) and is rather straightforward.
   • Any perturbation applied to $x_{2}$ will affect $x_{2}$ itself and $x_{1}$ after one time step, but not $x_{3}$. Hence, we infer $\partial f_{3}/\partial x_{2}=0$.
@@
 Causal relations in physical systems extend beyond direct interactions.
-A variable $x_{j}$ may affect another variable $x_{k}$ indirectly through an intermediate variable $x_{i}$, forming a pathway $x_{j}\rightarrow x_{i}\rightarrow x_{k}$.
-Such indirect links are central in climate dynamics [126] and in complex systems more broadly [127].
-Within the response framework, these pathways are revealed by the time-dependent operator $R_{k,j}(t)$: the first non-zero response identifies the shortest causal path in the graph connecting $x_{j}$ to $x_{k}$.
+A variable $x_{j}$ may affect another variable $x_{k}$ indirectly through an intermediate variable $x_{i}$, forming a path $x_{j}\rightarrow x_{i}\rightarrow x_{k}$.
+Such indirect links are central in climate dynamics [RungeCausalEffects] and in complex systems more broadly [Ay].
+Within the response framework, these paths are revealed by the time-dependent operator $R_{k,j}(t)$: the first non-zero response identifies the shortest causal path in the graph connecting $x_{j}$ to $x_{k}$.
 Crucially, if $R_{k,j}(t)=0$ for all time $t$, then no external intervention on $x_{j}$ can ever influence the behavior of $x_{k}$.
@@
 Finally, Baldovin et al.
-[17] introduced a cumulative (in time) measure of causation between two variables, defined as $D_{k,j}=\int_{0}^{\tau_{max}}R_{k,j}(t)dt$.
+[Baldovin] introduced a cumulative (in time) measure of causation between two variables, defined as $D_{k,j}=\int_{0}^{\tau_{max}}R_{k,j}(t)dt$.
 Such cumulative link, can be also referred to as “causal sensitivity”.
@@
 A concrete example using climate models, is provided by the approach of Bloch-Johnson et al.
-(2024) [38], where causal links are inferred between patterns of surface warming and the net radiative flux at the top of the atmosphere.
+(2024) [blochjohnsonetal2024], where causal links are inferred between patterns of surface warming and the net radiative flux at the top of the atmosphere.
 Their method reconstructs responses to perturbations by: (i) applying surface warming perturbations, (ii) integrating an atmospheric-only model with and without the perturbation, and (iii) computing the average (in time) difference between the two runs after removing an initial transient.
@@
 quantities averaged over the invariant density, ensures the robustness of the theory and largely removes the dependence on chaos.
-For a detailed discussion of this point, we refer the reader to [128].
+For a detailed discussion of this point, we refer the reader to [FALCIONI1995481].
 Future efforts to estimate responses in climate science, which often rely on perturbing single trajectories, would therefore benefit from focusing on ensemble-averaged observables.
 Response operator from data alone.
-Given the assumption of Markovianity it is possible to infer the system’s response operator from a long stationary trajectory using the Fluctuation-Dissipation Theorem (FDT) [17].
-The FDT formula provides a link between stationary and perturbed dynamics by predicting forced responses from stationary statistics alone [129].
+Given the assumption of Markovianity it is possible to infer the system’s response operator from a long stationary trajectory using the Fluctuation-Dissipation Theorem (FDT) [Baldovin].
+The FDT formula provides a link between stationary and perturbed dynamics by predicting forced responses from stationary statistics alone [Falcioni].
 Specifically:
@@
 $\frac{\partial\ln\rho(\mathbf{x})}{\partial x_{j}}\Big|_{\mathbf{x}(0)}$ represents the so-called “score function” evaluated at the initial state, i.e.
-on the attractor [15].
+on the attractor.
 Therefore, in principle the response operator can be computed from data if the score $\nabla\ln\rho(\mathbf{x})$ is available.
-In practice, the score is often inaccessible, and most applications rely on the quasi-Gaussian (qG) approximation, by replacing $\rho(\mathbf{x})$ with a Gaussian distribution $\rho^{G}(\mathbf{x})$ with same mean and covariances [130, 48].
-For responses in the ensemble mean—which are sufficient to infer causal links—the qG approximation achieves high skill even for nonlinear systems [121, 17, 114].
-Recent work [114] has demonstrated that generative modeling techniques can be used to efficiently estimate the score directly from data, leading to high skill in computation of higher-order responses (up to the 5th moment) in strongly nonlinear regimes.
-Other fruitful avenues lie in frameworks at the intersection of Markov State Modeling and Koopman operator theory such as [131].
+In practice, the score is often inaccessible, and most applications rely on the quasi-Gaussian (qG) approximation, by replacing $\rho(\mathbf{x})$ with a Gaussian distribution $\rho^{G}(\mathbf{x})$ with same mean and covariances [Leith, MajdaBook].
+For responses in the ensemble mean the qG approximation achieves high skill even for nonlinear systems [GRITSUN, Baldovin, LudovicoFabriAndre].
+Recent work [LudovicoFabriAndre] has demonstrated that generative modeling techniques can be used to efficiently estimate the score directly from data, leading to high skill in computation of higher-order responses in strongly nonlinear regimes.
+Other interesting new proposals have been recently formulated in the context of Koopman theory and can be used for future exploration of forced responses from data alone [ValerioData1, ValerioData2, FDTKoopman].
 It is worth emphasizing that the central limitation discussed in the main text, i.e.
-the need to choose the proper variables, remains relevant for the inference of response operators via FDT as well (and for any causal inference methodology); see the discussion in [17].
+the need to choose the proper variables for the system to be Markovian, remains relevant for the inference of response operators via FDT as well (and for any causal inference methodology); see the discussion in [Baldovin].
 Appendix B Relevance of the the triad system for geophysical turbulence
 Here, we motivate the choice of the triad model analyzed in Section III as a building block for geophysical turbulent flows.
-We further present and test the associated analytical stochastic model for the slow variable $x_{1}$ (as discussed in Section III.3.1), illustrating that Markovian reduced-order models can sometimes be derived from partial observations.
-For detailed mathematical derivations and additional examples, we refer to the original work by Majda and collaborators [46, 60, 61, 48, 16].
+We further present and test the associated analytical stochastic model for the slow variable $x_{1}$ (as discussed in Section III.3.1).
+For detailed mathematical derivations and additional examples, we refer to the original work by Majda and collaborators [NormalForms, MTV1, MTV2, MajdaBook, majdaSIAM].
 B.1 Link to geophysical turbulence
-Projecting many geophysical turbulent flows, such as quasigeostrophic models, onto a complete set of orthogonal basis functions (spectral or empirical modes such as EOFs), give rise to the following generic form for the mode amplitudes:
+Projecting many geophysical turbulent flows, such as quasigeostrophic models, onto a complete set of orthogonal basis functions (spectral or empirical modes such as EOFs), give rise to the following generic form for the modes amplitudes:
 $d\mathbf{u}/dt=\mathbf{F}+\mathbf{L}\mathbf{u}+\mathbf{B}(\mathbf{u},\mathbf{u}),$ (16)
@@
 The linear operator $\mathbf{L}=\mathbf{L}^{\prime}+\mathbf{D}$, is formed by a skew-symmetric operator $\mathbf{L}^{\prime}$ and by a diagonal dissipation matrix $\mathbf{D}$ with negative real entries.
-The quadratic nonlinearity should be energy-conserving, satisfying $\mathbf{u}\cdot\mathbf{B}(\mathbf{u},\mathbf{u})=0$ [132, 60, 48].
+The quadratic nonlinearity should be energy-conserving, satisfying $\mathbf{u}\cdot\mathbf{B}(\mathbf{u},\mathbf{u})=0$ [Lorenz, MTV1, MajdaBook].
 Eq.
-(16) is the typical abstract form of dry cores of atmospheric models [67].
-Galerkin truncations of turbulent systems often reduce to this structure, as in the Lorenz ’63 and Charney–DeVore models [132, 133, 68].
+(16) is the typical abstract form of dry cores of atmospheric models [MajdaPerspective].
+Galerkin truncations of turbulent systems often reduce to this structure, as in the Lorenz ’63 and Charney–DeVore models [Lorenz, Charney, CROMMELIN].
 $\mathbf{u}\in\mathbb{R}^{N}$ represents the full turbulent geophysical system, with $N$ infinite in principle and finite but very large in applications ($N\gg 1$).
 The state $\mathbf{u}$ is then decomposed into slow and fast components, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, i.e.
-$\mathbf{x}$ will represent the first $n$ components of $\mathbf{u}$ and $\mathbf{y}$ the next NnN-n.
+$\mathbf{x}$ represents the first $n$ components of $\mathbf{u}$ and $\mathbf{y}$ the next NnN-n.
 We consider the simple case of $N=3$: one large-scale resolved mode $x_{1}$ and two fast modes $\mathbf{y}=(x_{2},x_{3})$.
-Majda and collaborators [46] considered an exact projection of the slow mode $x_{1}$ and the two fast modes $(x_{2},x_{3})$ onto the dynamics in Eq.
+We then project the three modes onto the dynamics in Eq.
 (16).
@@
 (6)) to this study is that it provides a minimal multiscale setting with linear and quadratic nonlinear interaction of the kind that naturally arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs.
-For additional details see also Section 4 of [16].
+Furthermore, the model is characterized by fast decays of correlations, and generates PDFs that are regular and quasi-Gaussian, as often seen in the analysis of large-scale modes of climate variability [MajdaStructuralStability].
+For additional details see [NormalForms] and Section 4 of [majdaSIAM].
 B.2 Analytical stochastic model for
@@
 Starting from the triad model in Eq.
-(6), Majda and collaborators [60] showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ using standard projection techniques [48, 46, 47].
-The mode reduction predicts the functional form of the dynamics governing the resolved, slow variable [60, 61, 62, 63].
-The resulting stochastic equation is
+(6), Majda and collaborators [MTV1] showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ using standard projection techniques [MajdaBook, NormalForms, Majda2010].
+The mode reduction predicts the functional form of the dynamics governing the resolved, slow variable [MTV1, MTV2, MTV3, MTV4].
+The resulting stochastic equation is:
 $dx=(\tilde{F}+ax+bx^{2}-cx^{3})\,dt+(A-Bx)\,dW_{2}+\sigma\,dW_{3},$ (17)
@@
 $W_{2}$ and $W_{3}$ are independent Wiener processes.
-Interestingly, the equation for the coarse-grained variables $x_{1}$ display a structure richer than that of the full triad system in Eq.
+Interestingly, the equation for the coarse-grained variable $x_{1}$ displays a structure richer than that of the full triad system in Eq.
 (6).
 In addition to the original linear and quadratic interactions, the reduced equation includes new terms: an effective forcing $\tilde{F}$, a cubic nonlinearity, and both additive and multiplicative noise.
-The additive noise results from the linear interactions across fast modes, while the multiplicative noise reflects the advection of the slow mode by the fast modes [67].
+The additive noise results from the linear interactions across fast modes, while the multiplicative noise reflects the advection of the slow mode by the fast modes [MajdaPerspective].
 Thus, Eq.
 (17) demonstrates that multiplicative noise arises from theoretical considerations, rather than being introduced ad hoc, providing a firm basis for its use in reduced-order stochastic climate models.
-The scalar model above has served as a prototype in several contexts: it has been used to test the Fluctuation–Dissipation Theorem in climate dynamics [81, 114] as well as for inverse modeling of the leading principal component of a quasi-geostrophic model and the observed variability associated with the North Atlantic Oscillation (NAO), demonstrating its relevance for theoretical studies in climate dynamics.
+The scalar model above has been used to test the Fluctuation–Dissipation Theorem in climate dynamics [MajdaStructuralStability, LudovicoFabriAndre] as well as for inverse modeling of the leading principal component of a quasi-geostrophic model and the observed variability associated with the North Atlantic Oscillation (NAO), demonstrating its relevance for theoretical studies in climate dynamics.
 Figure 9: In each panel: $x_{1}$ Triad” represent statistics of $x_{1}$ in the original triad model in Eq.
@@
 The results are reported in Figure 9.
-The analysis show that, given $\epsilon\ll 1$, it is possible to build a Markovian model of the slow variable $x_{1}$ reproducing both stationary and perturbed statistics as in the original system.
-Finally, as shown in this Section, rigorous model reduction strategies such as those developed by Majda and collaborators [61] and by Roberts [134, 135] (among others [2]) offer invaluable theoretical guidelines for multiscale climate modeling.
-While such strategies where traiditionally developed in the case when the equation of motions are known, they can possibly serve as inspiration and leading to new directions in future work involving data-driven emulators.
-Finally, as shown in this Section, rigorous model reduction frameworks, such as those developed by Majda and collaborators [61] and by Roberts [134, 135], among others [2], provide invaluable theoretical guidance for multiscale climate modeling.
+The analysis shows that, given $\epsilon\ll 1$, it is indeed possible to build a Markovian model of the slow variable $x_{1}$ reproducing both stationary and perturbed statistics as in the original system.
+Finally, as shown in this Section, rigorous model reduction frameworks, such as those developed by Majda and collaborators [MTV2] and by Roberts [Roberts2006, ROBERTS200812], among others [LucariniChekroun], provide invaluable theoretical guidance for multiscale climate modeling.
 While these strategies were traditionally developed for systems with known equations of motion, they may also inspire future work aimed at incorporating such principles into data-driven-only emulators.
@@
 where $\bm{\xi}(t)$ is a Gaussian white noise with zero mean, unit variance, and $\sqrt{\Delta t}\bm{\xi}(t)$ is the increment of the Wiener process $\Delta\mathbf{W}=\mathbf{W}(t+\Delta t)-\mathbf{W}(t)$.
-Without loss of generality, we rescale time units by the time sampling interval such that a non-dimensional time increment is $\Delta t=1$ [6, 136, 7, 8].
+Without loss of generality, we rescale time units by the time sampling interval such that a non-dimensional time increment is $\Delta t=1$ [KRAVTSOV, KravtsovReview, KONDRASHOV201533, STROUNINE2010145].
 The strategy is then to fit a rescaled version of Eq.
@@
 In practice, this is solved as a linear regression problem by minimizing the mean squared error.
-We use the linear regression module from the scikit-learn library [137].
+We use the linear regression module from the scikit-learn library [scikit-learn].
     • Nonlinear drift.
 We consider the following form: $\Delta\mathbf{x}\approx\mathbf{f}(\mathbf{x}(t))=\mathbf{L}\mathbf{x}(t)+\mathbf{n}(\mathbf{x}(t))$, where $\mathbf{L}\in\mathbb{R}^{N,N}$ is the linear operator and $\mathbf{n}(\mathbf{x}(t))$ is the nonlinear operator parameterized by a multilayer perceptron (MLP).
-Both operators are fitted jointly by minimizing the mean squared error using a version of the gradient descent algorithm (Adam, [138]) with weight decay.
+Both operators are fitted jointly by minimizing the mean squared error using a version of the gradient descent algorithm (Adam, [kingma2014adam]) with weight decay.
 The MLP considered here always consisted of one hidden layer.
@@
       In practice, the loss function minimizes the mean squared error (MSE) between $\mathbf{C}(t)$ and $\mathbf{\Sigma}(\mathbf{x}(t))\mathbf{\Sigma}(\mathbf{x}(t))^{\text{T}}$ at each batch.
-A minimizer of the MSE loss approximates the conditional mathematical expectation $\mathbb{E}[\mathbf{C}(t)|\mathbf{x}(t)$], that is the covariance matrix averaged over all residuals corresponding to the same state vector, as a function of this state vector, and can be seen as a generalization of a method presented in [139] to multiple dimensions.
-The neural network architecture used here is identical to the one used for fitting the drift.
+A minimizer of the MSE loss approximates the conditional mathematical expectation $\mathbb{E}[\mathbf{C}(t)|\mathbf{x}(t)$], that is the covariance matrix averaged over all residuals corresponding to the same state vector, as a function of this state vector, and can be seen as a generalization of a method presented in [adler2018deep] to multiple dimensions.
 The positivity of the diagonal terms of the lower triangular matrix $\mathbf{\Sigma}(\mathbf{x}(t))$ is enforced through a softplus activation function.
@@
 $r(t)^{2}$.
-We then train a neural network to have $g(x(t))\approx r(t)^{2}$ by minimizing the mean squared error [139].
+We then train a neural network to have $g(x(t))\approx r(t)^{2}$ by minimizing the mean squared error [adler2018deep].
 The neural network is again a MLP with one hidden layer with 30 neurons.
@@
     $\begin{split}x(t+1)=x(t)+f(x(t))+\sqrt{g(x(t))}\xi(t).\end{split}$ (24)
+Appendix D Emulator of the SST field and net radiative flux: additive vs multiplicative noise
+In the main text, we considered the model defined in Eq. (11) to emulate the joint variability of SST and net radiative flux at the TOA. Here we examine a simpler variant with additive noise:
+$d\mathbf{x}=(\mathbf{L}\mathbf{x}+\mathbf{n}(\mathbf{x}))dt+\mathbf{\Sigma}d\mathbf{W}.$ (25)
+The drift term $\mathbf{L}\mathbf{x}+\mathbf{n}(\mathbf{x})$ is identical to the one used in the multiplicative-noise model in the main text.
+We performed a 600-year-long emulation following the same protocol as in the main text.
+The resulting autocorrelation functions, the reconstruction of net radiative flux, and the ensemble-mean response are all nearly indistinguishable from those obtained with multiplicative noise (Eq.
+(11)), as shown in Figures 6(c,d); 6(e,f); and 7.
+This demonstrates that the deterministic component (shared by both emulators) is responsible for setting these features.
+The main difference arises in the reconstruction of probability distributions.
+Figure 10 shows the PDFs for all 20 SST modes, as well as for the global-mean net radiative flux at the TOA.
+The PDF of the leading mode, associated with ENSO variability, is accurately reproduced by the emulator with multiplicative noise.
+Table 1 reports the first four moments of the distributions of this first mode for the climate model and for the two emulators.
+The most pronounced discrepancy appears in the standard deviation, where the multiplicative-noise emulator shows substantially closer agreement.
+The multiplicative formulation also succeeds in capturing the PDF of the second mode.
+Interestingly, for higher modes, each explaining a much smaller fraction of the variance, the emulator with additive noise performs slightly better.
+These differences, while visible, are modest if comparing the scale with the values of the first mode.
+We suspect that the very limited dataset (only 7200 samples in total for training and validation) is the main factor preventing the network from learning the multiplicative noise structure with higher accuracy.
+With a larger dataset, we expect the performance of the multiplicative-noise emulator to achieve more accurate reconstructions.
+Table 1: Mean, standard dev., skewness and kurtosis of SST Mode 1 (ENSO variability) for 600 years long runs.
+Mode 1 (ENSO) Mean St.Dev. Skewness Kurtosis
+CM4 Climate model 0. 28.27 -0.10 3.16
+Eq. (11) (Multiplicative noise) 2.83 29.07 -0.08 2.88
+Eq. (25) (Additive noise) -0.65 42.31 -0.13 3.33
+Figure 10: Panel (a-t): stationary distributions obtained the first 20 principal components of the sea surface temperature (SST) field as in the original GFDL-CM4 climate model and emulated by the neural emulator with multiplicative (Eq.
+(11)) and additive (Eq.
+(25)) noise.
+Units: nondimensional.
+Panel (u): same as panels (a-t) but for the global mean radiative flux at the top of the atmosphere (TOA).
+Units: $W/m^{2}$.
+The stationary distributions have been obtained after simulating the neural emulators for 600 years.
+Note that the system was trained on only 7200 time steps, which limits the accuracy of the learned multiplicative noise term.
 Acknowledgements.
@@
 I am especially grateful to Pasha Perezhogin, Rory Basinski, Chris Pedersen, and Andre Souza for their thoughtful feedback and for many enriching discussions on data-driven modeling of the Climate system.
-I acknowledge an interesting email exchange with Prof.
+I acknowledge an insightful email exchange with Prof.
 Tony Roberts, who drew my attention to recent advances on stochastic model reduction.
+I thank two anonymous reviewers for their very useful and constructive comments.
 This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.